### v1

# Scalable multiphoton generation from cavity-synchronized single-photon sources

Abstract
[…]
I Introduction and main results
Efficient sources of single and indistinguishable photons Eisaman et al. (2011); Senellart et al. (2017); Slussarenko and Pryde (2019) are a fundamental requirement to perform all kinds of quantum information tasks with photons: photonic quantum computation Knill et al. (2001); Kok et al. (2007); Tiecke et al. (2014); Tiarks et al. (2016); Takeda and Furusawa (2019), networking Lodahl (2017), simulation Aspuru-Guzik and Walther (2012); Hartmann (2016), communication Briegel et al. (1998); Sangouard et al. (2011); Llewellyn et al. (2020), cryptography Gisin et al. (2002), or boson sampling Brod et al. (2019); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a). Scaling up these protocols requires the generation of a large product-state of $N\gg 1$ indistinguishable photons,
[…]
propagating along $N$ or more channels. The generation of this state with high fidelity and efficiency demands the use of nearly deterministic, nearly identical and perfectly synchronized single-photon sources (SPSs), each of them producing just one photon.
The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al. (2016), which can be deterministically excited He and Barkai (2006a, b) and decay spontaneously, producing individual photons that are collected into the desired channels. The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs. A long list of setups includes single atoms McKeever et al. (2004); Hijlkema et al. (2007), single molecules Brunel et al. (1999); Ahtee et al. (2009); Rezai et al. (2018), trapped ions Almendros et al. (2009); Stute et al. (2012); Higginbottom et al. (2016); Sosnova et al. (2018), and atomic ensembles Chou et al. (2004); Matsukevich et al. (2006); Farrera et al. (2016); Ripka et al. (2018), as well as solid-state systems Aharonovich et al. (2016) such as quantum dots Santori et al. (2002); Nawrath et al. (2019); Wang et al. (2019b); Schnauber et al. (2019); Thoma et al. (2016); Dusanowski et al. (2019); Kirsanske et al. (2017); Liu et al. (2018); Uppu et al. (2020) or color centers in diamond Rodiek et al. (2017); Sipahigil et al. (2014); Alléaume et al. (2004); Gaebel et al. (2004); Wu et al. (2007); Babinec et al. (2010); Tchernij et al. (2017); Neu et al. (2011); Schröder et al. (2011); Wang et al. (2019c). In the microwave regime, superconducting quantum circuits have been used to create nearly deterministic SPSs Houck et al. (2007); Pechal et al. (2014); Peng et al. (2016); Forn-Díaz et al. (2017); Gasparinetti et al. (2017); Pfaff et al. (2017); Eder et al. (2018); Zhou et al. (2020), which have the advantage of being externally tuneable Peng et al. (2016); Forn-Díaz et al. (2017); Pfaff et al. (2017), fast Pechal et al. (2014); Zhou et al. (2020), and readily integrated on-chip with very low losses Zhou et al. (2020). Currently, quantum dots in micro-pillar cavities are proving the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al. (2019c); Uppu et al. (2020), but it is hard to manufacture many of them identically, and tuning them also remains elusive for more than two emitters Patel et al. (2010); Flagg et al. (2010); Moczała-Dusanowska et al. (2019); Ellis et al. (2018); Kambs and Becher (2018).
[…]
$P_{N}=({\cal P}_{1})^{N}S_{N},$ (2)
of the probability ${\cal P}_{1}$ of generating an individual photon on each of the $N$ channels independently, times the correlation error $S_{N}$ introduced by the circuit. An ideal scheme should scale exponentially $(S_{N}=1)$, creating $N$ synchronous and completely uncorrelated photons. However, typical active demultiplexing schemes introduce errors that scale as $S_{N}=\frac{1}{N}\left[\eta^{N}+(N-1)\left(\frac{1-\eta}{N-1}\right)^{N}\right]$ Lenzini et al. (2017), where $\eta$ characterizes the switching efficiency. Imperfect switching is a critical issue in any large photonic circuit Li et al. (2015), but even in the limit of lossless switches ($\eta=1$), the demultiplexing scheme introduces a detrimental factor $S_{N}=1/N$ to the scaling in Eq. (2). This can significantly limit the achievable $N$-photon efficiencies in practical applications such as boson sampling.
Figure 1: Cavity-QED scheme for deterministically generating a large number of synchronized and independent single-photons. (a) Idealized implementation with NN two-level emitters coupled to a common driven cavity mode. Each emitter additionally couples to independent output channels, which collect the $N$ emitted single-photons on parallel for later usage. (b) Efficient on-chip implementation using circuit-QED. Here, transmon qubits realize the NN two-level emitters, each of them capacitively coupled to an independent transmission-line antenna and to a common transmission-line resonator. Controlling the shape $\Omega(t)$ of the resonator drive, we can trigger all SPSs simultaneously and on-demand. The NN emitted microwave single-photons are collected by the on-chip antennas with high efficiency. These antennas are also used to send a DC current to each qubit and calibrate their frequencies into resonance with the resonator drive.
[…]
$S_{N}=1+\epsilon N(N-1).$ (3)
The correction factor $\epsilon$ stems from residual cavity-induced correlations. For realistic experimental conditions at the optimal working point, we find that $\epsilon\sim 10^{-8}-10^{-6}$, making the multiphoton generation scheme scalable up to $N\sim 100-1000\ll\epsilon^{-1/2}$, depending on the implementation and the noise sources.
To show the favorable scaling of the scheme in a realistic setup, we study a circuit-QED implementation with flux-tunable transmon emitters and microwave transmission lines DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014), considering dephasing noise, internal loss, and disorder. As shown in Fig. 1(b), this implementation is fully integrated on-chip, which allows the output antennas to have collection and transmission efficiencies above $99\%$ Zhou et al. (2020). For a strongly driven cavity, we show it is feasible to build microwave SPSs with single-photon purity and indistinguishability above $99\%$, as measured via standard Hanbury Brown and Twiss (HBT), and Hong-Ou-Mandel (HOM) experiments. Most importantly, we predict an overall single-photon efficiency or brightness of ${\cal P}_{1}\gtrsim 0.99$, a synchronized $30$-photon generation efficiency of $P_{30}\sim({\cal P}_{1})^{30}\sim 0.72$, and even $P_{100}\sim 0.33$ for $N=100$ SPSs. This means that for pulsed microwave SPSs with a repetition rate of $R\sim 0.4$ MHz, we can achieve $30$-photon generation rates of $C_{30}\sim 300$ kHz, and $100$-photon rates of $C_{100}\sim 100$ kHz. Remarkably, this is more than seven orders of magnitude more efficient than state-of-the-art boson sampling experiments with up to $N=14$ photons Wang et al. (2019a). The performance of our synchronized multiphoton generation scheme benefits from the high efficiency of the superconducting circuit implementation, as well as from the absence of switches and other optical elements. We also show that our predictions are robust to small inhomogeneities and disorder in the SPS parameters.
The paper is organized as follows. In Sec. II we introduce the setup, model, main approximations, and identify the optimal parameter regime of the scheme. In Sec. III, we introduce realistic parameter sets for a circuit-QED implementation and quantify its performance via the single- and multiphoton generation efficiencies, single-photon purity, and indistinguishability. The most important results of the paper are presented in Sec. IV, where we characterize the scalability of the synchronized $N$-photon generation. In Sec. V we analyze the effect of inhomogeneities and disorder in the SPS parameters, and finally in Sec. VI we summarize our conclusions. We complement the discussion by analytical and numerical methods to quantify the photon counting and correlations, which are briefly presented in Appendices A, B, and C.
II Setup and synchronization
In this section, we introduce a general cavity-QED model and possible nanophotonic implementations of the multiphoton emitter [cf. Sec. II.1]. We then discuss how to achieve cavity-mediated synchronization [cf. Sec. II.2], and finally we explain the basic working principle and the parameter conditions to achieve an efficient multiphoton generation [cf. Sec. II.3].
II.1 General cavity-QED model
Let us consider $N$ two-level systems or “qubits” coupled to an optical or microwave cavity as shown in Figs. 1(a) and (b). The cavity mode of frequency $\omega_{c}$ is coherently driven by a time-dependent field with frequency $\omega_{d}$, amplitude $\Omega(t)$ and detuning $\Delta=\omega_{d}-{\color[rgb]{0,0,0}\omega_{c}}.$ In a frame rotating with frequency $\omega_{d}$, the Hamiltonian of this system reads
$H(t)={}$ $-\Delta a^{{\dagger}}a+\Omega(t)(a+a^{{\dagger}})$
$-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}a).$ (4)
Here, $^{{\dagger}}, aa$ are the cavity creation and annihilation operators, whereas $\sigma_{j}^{+}=|e\rangle_{j}\langle g|$, $\sigma_{j}^{-}=|g\rangle_{j}\langle e|$, and $\sigma_{j}^{z}=|e\rangle_{j}\langle e|-|g\rangle_{j}\langle g|$ are standard Pauli operators for qubits $j=1,\dots,N$, with ground and excited states denoted by $|g\rangle_{j}$, and $|e\rangle_{j}$, respectively. Each qubit has a possibly different frequency $\omega_{q}^{j}$ and detuning $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ with respect to the cavity drive. The interaction strength $g_{j}$ between the $j$-th qubit and the cavity satisfies the rotating-wave approximation (RWA) $g_{j}\ll\omega_{q}^{j}\sim\omega_{d},$ and coupling terms can be written in the standard Jaynes-Cummings form. We control the emitters by modulating the envelope of the cavity drive
$\Omega(t)=\Omega_{0}f(t),$ (5)
using a smooth square pulse $f(t)$ of duration $T$ and maximum amplitude $\Omega_{0}.$
Additionally, each qubit $j$ is coupled to an independent decay channel or “antenna”, which collects the emitted photons [cf. Fig.1(a) and (b)]. We describe this qubit-antenna interaction in the the Born-Markov approximation, introducing the $\gamma_{j}$ at which a qubit deposits photons into its own antenna. We also introduce a photon loss rate $\gamma_{\rm loss}^{j}$ characterizing the emission of photons into any other unwanted channel. Finally, we consider the cavity decay rate $\kappa$ and introduce white noise dephasing rates $\gamma_{\phi}^{j}$ on each of the qubits. The complete dynamics of this open system is described with a master equation,
$\dot{\rho}(t)={}$ $-i[H(t),\rho]+\kappa{\cal{D}}[a]\rho+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho$
$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\rho+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\rho,$ (6)
modelling the mixed quantum state of the qubits and the cavity mode $\rho(t)$, with the system Hamiltonian $H(t)$ from Eq. (4), and the Lindblad terms ${\cal{D}}[x]\rho=x\rho x^{\dagger}-(x^{\dagger}x\rho+\rho x^{\dagger}x)/2$.
The idealized setup in Fig. 1(a) admits various implementations. The two-level systems $\sigma_{j}^{-}$ could be neutral atoms Welte et al. (2018) or ions Casabone et al. (2015) trapped inside an optical cavity field $a$ that is localized between macroscopic mirrors. In this prototypical cavity-QED implementation the individual decay channels of each qubit would require the use of high-aperture lenses Araneda et al. (2018) or tapered nano-fibers Reitz et al. (2013) to collect the photons independently, which is very challenging to integrate and to scale-up with high collection efficiencies. Nanophotonic structures such as photonic crystals Lodahl et al. (2015) or integrated photonic circuits Antón et al. (2019); Mehta et al. (2020) are other promising platforms to realize our setup. The common mode and the independent output channels can be integrated and scaled up. In this scenario, the main limitation arise from the creation of nearly identical or tuneable emitters such as quantum dots Ellis et al. (2018), or alternatively the trapping of many atoms Goban et al. (2015) or ions Mehta et al. (2020) near these surfaces.
However, in this work we will focus on circuit-QED to discuss an efficient and scalable integration of multiple two-level emitters with individual decay channels. As sketched in Fig. 1(b), one may use superconducting transmon qubits as quantum emitters DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014). These qubits may be capacitively coupled to both a common transmission-line resonator, as well as to individual transmission-line waveguides that efficiently collect the microwave single-photons Zhou et al. (2020). In Secs. III and IV we analyze in detail the performance and scalability of this superconducting circuit platform.
II.2 Cavity-mediated synchronization and residual correlations
The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode, using a strong drive $|\Omega_{0}|\gg\kappa,|\Delta|$. We model this state using a displacement of the Fock operator
$a=\alpha+\delta a,$ (7)
where the amplitude $\alpha(t)$ is given by the classical harmonic oscillator equation
$\dot{\alpha}(t)={}$ $-\left[\kappa/2-i\Delta\right]\alpha(t)-i\Omega_{0}f(t).$ (8)
Once the pulse is switched on ($f(t)=1$), $|\alpha(t)|$ grows and reaches the steady state value
$|\alpha_{\rm ss}|=\frac{\Omega_{0}}{\sqrt{(\kappa/2)^{2}+\Delta^{2}}}\gg 1.$ (9)
Similarly, once we switch the drive off $(f(t\geq T)=0),$ the cavity displacement vanishes in a time scale $\sim 1/\kappa.$
As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, described by the effective Hamiltonian, $\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j}g_{j}|\alpha(t)|(\sigma_{j}^{-}+\sigma_{j}^{+}).$ In the asymptotic limit, the large and steady cavity amplitude allows for a strong cavity-mediated driving strength on the qubits
$\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (10)
which we use to excite all SPSs simultaneously.
In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed. To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
$g_{j},|\delta_{j}|\ll\kappa,|\Delta|.$ (11)
These conditions [cf. App. A.2] ensure that the cavity reaches a steady state (9) where quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated. The dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\},$ is then modeled by the master equation
$\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\cal{D}}\left[\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}\right]\tilde{\rho}$
$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$, (12)
with the effective Hamiltonian $\tilde{H}(t)$
$\tilde{H}(t)={}$ $-\frac{1}{2}\sum_{j=1}^{N}(\delta_{j}-\delta^{j}_{\rm cm})\sigma_{j}^{z}+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}f(t)(\sigma_{j}^{-}+\sigma_{j}^{+})$
$+\sum_{j>l}J^{jl}_{\rm cm}\left(\sigma_{j}^{+}\sigma_{l}^{-}+\sigma_{l}^{+}\sigma_{j}^{-}\right).$ (13)
In addition to the cavity-mediated driving $\Omega_{\rm cm}^{j}$ in Eq. (10), the qubits experience a Lamb-shift
$\delta^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\Delta}{(\kappa/2)^{2}+\Delta^{2}},$ (14)
and undergo long-range cavity-mediated interactions, with couplings
$J^{jl}_{\rm cm}={}$ $\frac{g_{j}g_{l}\Delta}{(\kappa/2)^{2}+\Delta^{2}}.$ (15)
The cavity also induces a collective decay or “superradiance” on the qubits, described by a collective jump operator $W_{\rm cm}=\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma^{-}_{j}.$ The cavity-mediated decay rates
$\gamma^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\kappa}{(\kappa/2)^{2}+\Delta^{2}},$ (16)
combined with the local decay and dephasing rates $\gamma_{j}$, $\gamma_{j}^{\rm loss}$, and $\gamma_{\phi},$ describe the incoherent dynamics of the qubit, as shown in Eq. (12). We will see that, for an optimal performance of the synchronized $N$-photon emitter, the residual cavity-mediated interactions $J_{\rm cm}^{jl}$ and decay $\gamma_{\rm cm}^{j}$ must be much smaller than the cavity-mediated driving $\Omega_{\rm cm}^{j}.$
II.3 Synchronization dynamics and parameter regime
Our goal is to realize the synchronized excitation and emission of each qubit, so that they act as nearly independent SPSs. The first stage of operation involves exciting each two-level systems from the ground state $|g\rangle_{j}$ to the excited state $|e\rangle_{j}$ via a fast cavity-mediated $\pi$-pulse of duration $T\sim\pi/(2\Omega_{\rm cm}^{j})$. We then expect that all emitters will produce synchronized photons, one on each of the antennas, on a timescale $t\gg 1/\gamma_{j}$. To ensure that this procedure efficiently generates $N$ nearly indistinguishable and independent single-photons, the system parameters must satisfy the conditions,
$\Omega_{\rm cm}\gg{}$ $\gamma\gg J_{\rm cm},\gamma_{\rm cm},\gamma_{\rm loss},\gamma_{\phi},$ (17)
and $\delta=\delta_{\rm cm}$, as explained in the following.
First, to achieve synchronization and indistinguishability of the photon emissions, the system parameters should be as homogeneous as possible, and specially the qubit frequencies $\omega_{q}\approx\omega_{q}^{j}$, couplings $g\approx g_{j}$, and antenna decays $\gamma\approx\gamma_{j}$. In the remainder of the paper we will thus assume that all system parameters are homogeneous, except in Sec. V, where we analyze inhomogeneities and disorder.
Second, to achieve high efficiency and purity of single-photon emission we need to drive the emitters on resonance ($\delta=\delta_{\rm cm}$). We must also excite the qubits very fast ($\Omega_{\rm cm}\gg\gamma$) to suppress events where the qubit emits two photons—it creates a photon during the excitation pulse, gets excited by the remaining of the pulse, and emits a second photon—. Note that we can mitigate this effect using three-level emitters as SPSs Fischer et al. (2016), but this is not the focus of the present work.
Third, to achieve nearly independent photons in a product state (1), the photons must be created faster than the speed of the cavity mediated correlations and interactions ($\gamma\gg J_{\rm cm},\gamma_{\rm cm}$). This is the only intrinsic limitation of the proposed multiphoton scheme. In Sec. III-IV we show that state-of-the-art circuit-QED setups satisfy this requirement and allow the synchronization of $N\sim 100-1000$ nearly independent SPSs.
The fourth condition, $\gamma\gg\gamma_{\rm loss}$, ensures an efficient collection of photons by the antennas, with minimal losses, while the last condition, $\gamma\gg\gamma_{\phi}$, ensures that those photons are phase-coherent and indistinguishable as detected by HOM interference [cf. Sec. III.5]. The degree to which we satisfy these conditions is limited by the technology of the SPSs—atoms, dots, superconductors, etc—, but we will see that they are well met by state-of-the-art circuit-QED setups [cf. Sec. III].
Figure 2: Synchronization of two deterministic SPSs via a cavity. (a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity [cf. red dash-dotted line]. This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf. blue solid line]. After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$. (b) Long time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its own antenna [cf.  green dashed line]. Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. orange dotted line]. The parameters correspond to a state-of-the-art circuit-QED implementation, and are presented in row C of Table 1.
$\frac{\gamma}{2\pi}[{\rm MHz}]$ $\frac{\gamma_{\rm loss}}{2\pi}[{\rm kHz}]$ $\frac{\gamma_{\phi}}{2\pi}[{\rm kHz}]$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\phi}/\gamma$ $\gamma_{\rm loss}/\gamma$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
A 0.4 0 0 $1.0\cdot 10^{-2}$ 0 0 $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.997 0.994 $4.4\cdot 10^{-3}$ $2.0\cdot 10^{-5}$
B 0.4 1 1 $1.0\cdot 10^{-2}$ $2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.994 0.989 $4.4\cdot 10^{-3}$ $2.5\cdot 10^{-3}$
C 1.0 5 10 $2.6\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $5.0\cdot 10^{-3}$ $3.8\cdot 10^{-4}$ $4.7\cdot 10^{-5}$ 0.989 0.979 $1.0\cdot 10^{-2}$ $9.7\cdot 10^{-3}$
D 3.0 50 100 $7.7\cdot 10^{-2}$ $3.3\cdot 10^{-2}$ $1.7\cdot 10^{-2}$ $1.3\cdot 10^{-4}$ $1.6\cdot 10^{-5}$ 0.968 0.936 $3.0\cdot 10^{-2}$ $3.1\cdot 10^{-2}$
Table 1: Parameters and figures of merit for a circuit-QED implementation of two synchronized SPSs. We study four parameter sets: A, B, C, and D, which show increasing antenna coupling $\gamma$ and decoherence rates $\gamma_{\rm loss}$, $\gamma_{\phi}$. The rest of the parameters are common to all sets and read $g/2\pi=0.2{\rm MHz}$, $\kappa/2\pi=400{\rm MHz}$, $\Delta/2\pi=50{\rm MHz}$, and $\Omega_{0}/2\pi=40{\rm GHz}$, which leads to $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$ photons in the cavity on average. The detuning of the qubits is chosen to compensate for the cavity-induced Lamb-shift $\delta_{q}=\delta_{\rm cm}$. Depending on the decoherence rates of sets A-D, the inequalities in Eq. (17) are satisfied differently, as shown in columns 5th to 9th, and this influences the performance of the four figures of merit: single-photon generation probability ${\cal P}_{1}$ [cf. Sec. III.2], synchronized two-photon generation efficiency or probability $P_{2}\approx({\cal P}_{1})^{2}$ [cf. Sec. III.3], multi-photon contamination $g^{(2)}_{\rm HBT}[0]$ [cf. Sec. III.4], and photon distinguishability $g^{(2)}_{\rm HOM}[0]$ [cf. Sec. III.5]. All these quantities are calculated using the full master equation (6), the methods in Appendices B.1 and C, and evolving up to a final time $t=15/\gamma$.
Fig. 2 illustrates the two processes in the cavity-mediated synchronization method, using two sychronized SPSs with parameters that satisfy Eq. (17). Fig. 2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|\sim 10^{4}$ [cf. red dash-dotted line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf. blue solid line]. Fig. 2(b) shows the long time dynamics, in which the photons are generated. Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf. blue solid line]. To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs. III.2-III.3. For non-optimized parameters satisfying Eq. (17), we reach ${\cal P}_{1}\approx 0.99$ [cf. green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. orange dotted line]. This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq. (2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$). We remark that this example only illustrates the working principle of the setup. As we show in Secs. III and IV, the parameters can be further optimized to improve the efficiency and scalability of the multi-photon emitter.
III Performance of a circuit-QED implementation
In this section, we describe typical state-of-the-art parameters to realize the synchronized multiphoton emitter in a circuit-QED implementation [cf. Sec. III.1]. Subsequently, we quantify the performance of the scheme using four figures of merit: (i) efficiency of single-photon generation ${\cal P}_{1}$ [cf. Sec. III.2], (ii) efficiency of synchronized $N$-photon generation $P_{N}$ [cf. Sec. III.3], (iii) single-photon purity as quantified by HBT correlations, $p=1-g^{(2)}_{\rm HBT}[0]$ [cf. Sec. III.4], and (iv) photon indistinguishability as quantified by HOM interference, $I=1-g^{(2)}_{\rm HOM}[0]$ [cf. Sec. III.5].
III.1 State-of-the-art parameters
In Fig. 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits. For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency on the order of a few GHz. These superconducting resonators support strong drives and can hold a large number of photons in steady state Pietikäinen et al. (2017); Jenkins et al. (2014); Macri et al. (2020). In our design we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al. (2017); Jenkins et al. (2014).
As two-level emitters, we consider flux-tunable transmon qubits DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Wang et al. (2020) weakly coupled to the resonator mode with strength $g/2\pi=0.2{\rm MHz}$. The cavity mediated driving amplitude is $\Omega_{\rm cm}/2\pi=|\alpha_{\rm ss}|g/2\pi\approx 39$MHz. This is strong but still smaller than typical transmon anharmonicities $U/2\pi\sim 300-400{\rm MHz}$ Pechal et al. (2014); McKay et al. (2019); Zhou et al. (2020), thus preventing leakage to higher excitation states. The frequency of the qubits—which will be on the few GHz range—can be fine-tuned by sending a DC current through the antenna that is coupled to each qubit [cf. Fig. 1(b)]. This is required to bring the qubits in resonance with the external drive, and to compensate the cavity-induced Lamb-shifts $\delta_{q}=\delta_{\rm cm}$ [See Sec. V for small deviations of this condition].
The antenna of each qubit is not only used for frequency calibration, but its main purpose is to collect the photons emitted by each qubit [cf. Fig. 1(b)]. The coupling rate between the qubit and antenna $\gamma$ can be chosen depending on the estimated decoherence parameters $\gamma_{\rm loss}$ and $\gamma_{\phi}$, so that the photons are efficiently ($\gamma\gg\gamma_{\rm loss}$) and coherently ($\gamma\gg\gamma_{\phi}$) collected. In our analysis, we consider four different combinations of the parameter sets ($\gamma,\gamma_{\rm loss}$$\gamma_{\phi}$), shown in Table 1. Parameter set D corresponds to the typical relaxation $\gamma_{\rm loss}/2\pi\sim 50$kHz and dephasing rates $\gamma_{\phi}/2\pi\sim 100$kHz found in microwave quantum photonics experiments Pechal et al. (2014); Zhou et al. (2020); Wang et al. (2020). For this case we choose a relatively large decay rate of $\gamma/2\pi=3$MHz so that the emission time is faster than the decoherence timescales, but at the same time it is still slower than the effective driving $\Omega_{\rm cm}/2\pi\sim 39$MHz [cf. column D in Table 1]. In addition, parameter set C corresponds to decoherence rates reported in quantum computing experiments Barends et al. (2013); Caldwell et al. (2018); McKay et al. (2019); Andersen et al. (2019), which have reached relaxation and dephasing rates as low as $\gamma_{\rm loss}/2\pi\sim 5$kHz and $\gamma_{\phi}/2\pi\sim 10$kHz, respectively. For these parameters we can choose a slightly lower coupling to the antenna of $\gamma/2\pi\sim 1$MHz, and thus improve the ratio $\gamma/\Omega_{\rm cm}\sim 2.6\cdot 10^{-2}$ [cf. column C in Table 1]. Finally, set B corresponds to near-future parameters and set A to the ideal case of no decoherence, which we use as a reference to quantify the intrinsic limitations of the synchronization method. In the remainder of the paper, we use these four parameter sets A-D to characterize the performance and scaling of the synchronized SPSs for this circuit-QED implementation.
Figure 3: Few-photon statistics ${\cal P}_{n}$ and two-photon dependence $D_{2}$ as a function of various system parameters such as (a) cavity occupation $|\alpha_{\rm ss}|^{2}$, (b) qubit-cavity coupling $g/\kappa$, (c) qubit-antenna decay $\gamma/\Omega_{\rm cm}$, (d) qubit-drive detuning $(\delta-\delta_{\rm cm})/\Omega_{\rm cm}$, (e) dephasing $\gamma_{\phi}/\gamma$, and (f) relaxation loss $\gamma_{\rm loss}/\gamma$. The probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) are plotted using the left vertical axis, whereas the two-photon dependence $D_{2}$ (red), corresponds the right vertical axis. Solid lines are computed using the effective master equation in Eq. (12) and the discrete symbols using the full master equation in Eq. (6), which show excellent agreement. Parameters are given by set C of Table 1, except for the quantity that is changed on each subplot.
III.2 Efficiency of single-photon generation
Without loss of generality, in our discussion about photon generation efficiencies or probabilities, we assume perfect photo-detectors  [^1] Detector inefficiencies $\eta_{D}<1$ give rise to an extra exponential factor $\sim(\eta_{D})^{N}$ Lenzini et al. (2017) in the $N$-photon probability (2). See also Refs. Eisaman et al. (2011); Slussarenko and Pryde (2019) for more details.. In this scenario, the efficiency of a single photon source can be extracted from the photon emission statistics. For a single emitter, ${\cal P}_{1}$ is the probability of generating exactly an isolated photon in a Fock state Senellart et al. (2017). Similarly, we define ${\cal P}_{n}^{j}$ as the probability for the synchronized SPS to emit a Fock state of $n$ photons in the $j$-th antenna. Formally,
${\cal P}_{n}^{j}={\rm Tr}\{\Lambda_{n}^{j}\rho_{\rm ext}(t)\},$ (18)
where $\rho_{\rm ext}(t)$ is the state of the extended system including the photons emitted in all antennas. The operator $\Lambda_{n}^{j}=|n\rangle_{j}\langle n|$ projects $\rho_{\rm ext}(t)$ on the Fock state $|n\rangle_{j}$ of $n$ photons propagating in the independent antenna $j$. Computing ${\cal P}_{1}$ or ${\cal P}_{n}^{j}$ seems to require a large Hilbert space, but there are shortcuts, such as the quantum jump formalism or our new photon counting approach based on master equations [cf. App. B].
A nearly deterministic single-photon source should have ${\cal P}_{1}\approx 1$, other few-photon probabilities ${\cal P}_{0}$ and ${\cal P}_{2}$ should be strongly suppressed, and ${\cal P}_{n\geq 3}\approx 0$ negligible. This also applies to each of the probabilities ${\cal P}_{n}^{j}$ in our synchronized setup, if we want it to operate as a collection of $N$ independent emitters.
To verify this, we have computed the few-photon probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) for a setup with $N=2$ superconducting qubits, varying different parameters around the set C from Table 1. The results in Figs. 3(a)-(f), have been computed using the full model (6) and the effective equation (12), and are shown in lines and markers, respectively. The agreement between both methods confirms that the adiabatic elimination of the cavity is an excellent approximation over a broad range of parameters.
In Fig. 3(a) we plot the effect of cavity occupation $|\alpha_{\rm ss}|^{2}$ (controlled via the cavity drive $\Omega_{0}$) on the single-photon efficiency of the SPS. All errors reduce with the cavity population $|\alpha_{\rm ss}|^{2},$ because a larger cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$ increases the speed at which the qubit is excited and reduces the probability of emitting two consecutive photons. However, the driving strength is limited by our need to have an off-resonant cavity and by the anharmonicity of the qubits [cf. Sec. III.1].
Fig. 3(b) illustrates the need to find an optimum value of the coupling $g$, within inequalities in Eq. (17). Initially, increasing $g$ increases the driving $\Omega_{\rm cm}$ and reduces errors. However, a large coupling strength enhances the cavity-mediated dipole-dipole interactions $J_{\rm cm}\sim g^{2}$ and the collective decay $\gamma_{\rm cm}\sim g^{2}$. Both result in the transfer of excitations between qubits, increasing the probability of no emission ${\cal P}_{0}$ and of two photon emission ${\cal P}_{2}.$ For the parameters in Table 1 we find an optimal operation point around $g/\kappa=4\cdot 10^{-4}$.
In Fig. 3(c) we analyze the effect of the coupling to the antenna coupling $\gamma.$ Once more, there is an optimal point that satisfies (17), balancing the imperfections due to unwanted re-excitations ($\Omega_{\rm cm}\gg\gamma$) and cavity-mediated effects ($\gamma\ll J_{\rm cm},\gamma_{\rm cm}$). The optimal lays around $\gamma/\Omega_{\rm cm}\sim 2\cdot 10^{-2}$ for the parameter set C in Table 1.
Fig. 3(d) clearly shows an optimal operation of the emitters when they are on resonance with the drive $\delta=\delta_{\rm cm}$. At this point, a $\pi$ pulse excites the qubits with high fidelity, and the Lamb-shift is compensated.
Finally, in Fig. 3(e) and Fig. 3(f), we observe that the dephasing $\gamma_{\phi}$ and the losses into uncontrolled channels $\gamma_{\rm loss}$ have a negligible influence on the few-photon statistics ${\cal P}_{n}$ as long as they satisfy $\gamma_{\phi},\gamma_{\rm loss}\ll\gamma$ [cf. Eq. (17)]. Nevertheless, when $\gamma_{\phi},\gamma_{\rm loss}\sim\gamma$ we see a rapid increase of events where no photon is detected ${\cal P}_{0}.$ This is due to emitters becoming effectively off-resonant from the drive when $\gamma_{\phi}\sim\gamma$, or due to a decrease in the the collection efficiency in the antenna, when $\gamma_{\rm loss}\sim\gamma.$
Figure 4: Multi-photon contamination of the SPS measured via the second-order correlation function. (a) Schematic of the Hanbury Brown and Twiss (HBT) experiment using a beam-splitter (BS) and two-photon coincidence measurements. (b) HBT correlation function $G^{(2)}_{\rm HBT}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set C of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HBT}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HBT}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. Other parameters correspond to set C (blue) and set B (red). Decoherence has a small detrimental effect on the observed multi-photon contamitation via $g^{(2)}_{\rm HBT}[0]$.
III.3 Efficiency of synchronized N-photon generation
The quantity ${\cal P}_{1}$ is the probability of generating one photon on a given antenna irrespective of the photons emitted in the rest of the channels. To study SPS synchronization, however, we are interested in the probability $P_{N}$ of emitting exactly one photon on each of the $N$ available antennas or channels. This can be expressed using projectors onto single-photon Fock states $\Lambda_{1}^{j}=|1\rangle_{j}\langle 1|$ on each the channels
$P_{N}={\rm Tr}\{\Lambda_{1}^{1}\cdots\Lambda_{1}^{N}\rho_{\rm ext}(t)\},$ (19)
An efficient generation of $N$ synchronized and independent single-photons should satisfy $P_{N}\approx({\cal P}_{1})^{N}\leq{\cal P}_{1}$, meaning that the $N$-photon product state (1) is generated with high fidelity. This happens for all parameter sets in Table 1. The two-photon generation probability satisfies $P_{2}\approx({\cal P}_{1})^{2}\lesssim 1,$ and deviates from unity because of errors in the single-photon efficiency ${\cal P}_{1}\lesssim 1$.
To quantify more precisely the independence of the $N$ emitted photons, we define the $N$-photon dependence or demultiplexing error as,
$D_{N}=\frac{P_{N}}{({\cal P}_{1})^{N}}-1,$ (20)
which describes the deviation from the ideal limit of $N$ perfectly independent and synchronized SPSs. This quantity is also related to the demultiplexing inefficiency in Eq. (2) as $S_{N}=1+D_{N}$.
We have computed $P_{N}$ and $D_{N}$ in the simplest case of two SPSs, using the same photon counting methods introduced in Appendix B. The results of these simulations are shown as red curves of Figs. 3(a)-(f) [cf. red curves and right vertical axis]. We observe that $D_{2}$ is nearly insensitive to changes in the cavity occupation $|\alpha_{\rm ss}|^{2}$, the dephasing $\gamma_{\phi}/\gamma$, and out-coupling efficiency $\gamma_{\rm loss}/\gamma$, with typical values on order $D_{2}\sim 10^{-7}$ for parameters around set C of Table 1. Nevertheless, the variables $g$, $\gamma$, and $\delta$ can drastically change $D_{2}$ in various orders of magnitude. In particular, for larger cavity coupling $g$ or smaller antenna decay $\gamma$, $D_{2}$ increases because the cavity-mediated correlations $\sim g^{2}$ become more important in the timescale of the photon emission $\sim 15/\gamma$ [cf. right inequality in Eq. (17)].
Finally, we note that the exact pulse shape $f(t)$ is not relevant for the synchronization dynamics as long as it approximately induces a $\pi$-pulse on the qubits. To study this, we have performed the calculations in Figs. 3(a)-(f) using an ideal square pulse and a smooth step pulse $f(t)=\frac{1}{2}\tanh(\frac{t}{\tau_{r}})\{2-\tanh(\frac{T}{\tau_{r}})-\tanh(\frac{t-T}{\tau_{r}})\}$ with duration $T$ and ramp time $\tau_{r}$. When optimizing over $T$ and $\tau_{r}$ of the smooth pulse, we obtain a deviation smaller than $1\%$ with respect to the result of the simpler square $\pi$-pulse with $T=\pi/(2\Omega_{\rm cm})$. Therefore, in the remainder of the paper we safely consider the ideal square pulse, but we keep in mind that a realistic experimental pulse shape gives similar results when properly optimized.
III.4 Single-photon purity
A standard figure of merit to experimentally quantify the amount of multi-photon contamination of a SPS is the second-order correlation function $G_{\rm HBT}^{(2)}$ Kiraz et al. (2004); Fischer et al. (2016). As sketched in Fig. 4(a), this is measured in a Hanbury Brown and Twiss (HBT) setup, where the output of a given SPS is beam-splitted and measured via coincidences in two intensity detectors. For pulsed emission, the correlation function $G_{\rm HBT}^{(2)}(\tau)$ is defined as Kiraz et al. (2004); Fischer et al. (2016)
$G_{\rm HBT}^{(2)}(\tau)\!=\!$ $\int\!\mathrm{d}t\,\langle b^{j{\dagger}}_{\rm out}(t)b^{j{\dagger}}_{\rm out}(t+\tau)b^{j}_{\rm out}(t+\tau)b^{j}_{\rm out}(t)\rangle,$ (21)
where $\tau$ is the time delay between the two photon detections, and $b^{j}_{\rm out}(t)$ is the annihilation operator for an outgoing photon on the $j$-th antenna Gardiner and Zoller (2004). Appendix C contains details on the input-output theory and the calculation of these correlations. For simplicity, our analysis assumes a homogeneous setup and thus we omit the $j$ index in $G_{\rm HBT}^{(2)}(\tau)$.
Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference. (a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements. (b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$. Parameters correspond to set C of Table 1. (c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$. (d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively. Other parameters correspond to set C (blue) and set B (red). The indistinguishability dramatically reduces with dephasing, a behaviour in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al. (2003).
In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for train of 4 excitation pulses with a repetition rate
$R=\gamma/15.$ (22)
We observe clear peaks at integer multiples of the repetition time $\tau=n/R$, corresponding to the detection of two photons coming from different pulses. Relevant to characterize the few-photon statistics of the SPS is the small peak that appears near zero time delay, $\tau\sim 0.$ A small area in this peak signals a small probability of two- and multi-photon emission Kiraz et al. (2004). In Fig. 4(c) we enlarge the region $\tau\sim 0$ and show that the area of the zeroth peak increases with $g/\kappa,$ consistently with the behavior of ${\cal P}_{2}$ in Fig. 3(b) for the same parameters.
To quantify more precisely the amount of multi-photon contamination of the SPS, and in a way that is independent of the input power, it is convenient to define the normalized correlation $g_{\rm HBT}^{(2)}[0]$ at zero delay as Fischer et al. (2016)
$g_{\rm HBT}^{(2)}[0]=\frac{\int_{0}^{1/(2R)}d\tau G_{\rm HBT}^{(2)}(\tau)}{\left(\int_{0}^{1/(2R)}dt\langle b_{\rm out}^{j{\dagger}}(t)b_{\rm out}^{j}(t)\rangle\right)^{2}}.$ (23)
Here, the numerator corresponds to the area of the zeroth peak at $\tau\sim 0$ and the normalization is the area of one of the high peaks at $\tau\sim n/R$, and thus $g_{\rm HBT}^{(2)}[0]\sim{\cal P}_{2}/({\cal P}_{1})^{2}$. In Figs. 4(d)-(e), we plot the normalized correlation $g_{\rm HBT}^{(2)}[0]$ as a function of $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma$, respectively. We observe that $g_{\rm HBT}^{(2)}[0]$ is very insensitive to both types of decoherence, which is also consistent to the behaviour of ${\cal P}_{2}$ in Fig. 3(e)-(f) for the same parameters. In addition, we see that parameter set B gives a two-photon contamination of $g_{\rm HBT}^{(2)}[0]\sim 2.5\cdot 10^{-3}$, slightly lower than parameter set C [cf. Fig. 4(e)-(f) and Table 1]. Finally, we note that the single-photon purity $p$ is sometimes defined as $p=1-g_{\rm HBT}^{(2)}[0]$, which for parameter sets B and C of our circuit-QED implementation is higher than $p=0.99$.
III.5 Indistinguishability
Another important aspect of a high-performance multiphoton demultiplexing scheme is the degree of indistiguishability of the photons emitted by two different sources Senellart et al. (2017). Experimentally, the distinguishability of two sources is typically quantified using Hong-Ou-Mandel (HOM) interference as shown in Fig. 5(a). Photons coming from two sources interfere on a beam-splitter (BS), and a pair of detectors count coincidences. The result of these measurements are the HOM correlations $G_{\rm HOM}^{(2)}(\tau)$, which for pulsed emission read Fischer et al. (2016)
$G_{\rm HOM}^{(2)}(\tau)\!=\!\int dt\langle b_{\text{BS}}^{1{\dagger}}(t)b_{BS}^{2{\dagger}}(t+\tau)b_{BS}^{2}(t+\tau)b_{BS}^{1}(t)\rangle,$ (24)
where $\tau$ is the delay time between the two photon detections. In addition, $b_{BS}^{1}(t)=[b^{1}_{\rm out}(t)+b^{2}_{\rm out}(t)]/\sqrt{2}$, and $b_{BS}^{2}(t)=[b^{1}_{\rm out}(t)-b^{2}_{\rm out}(t)]/\sqrt{2}$, describe the output fields of photons after interfering at the BS [cf. Appendix C].
Figure 6: Scalability performance of the synchronized multiphoton generation scheme. (a) NN-photon generation efficiency $P_{N}$ as a function of $N$, for parameter sets A (green), B (blue), C (red), and D (grey), as given in Table 1. Data points up to $N=12$ correspond to QT calculations averaged over $M=6000$ trajectories, which show excellent agreement (within the statistical error $\Delta P_{N}\sim 10^{-3}$) with the prediction for perfectly independent and synchronized SPSs, i.e. $P_{N}\approx({\cal P}_{1})^{N}$. [cf. black lines and inset]. (b) NN-photon generation rate $C_{N}=RP_{N}$ for pulsed excitation with repetition rate $R=\gamma/15$. Despite parameter set D shows the lowest $P_{N}$ efficiency, its faster repetition rate $R\sim 1.2$MHz allows for appreciably higher $C_{N}$. (c) Demultiplexing error $D_{N}$ up to $N=5$, and quadratic fit $D_{N}=\epsilon N(N-1)$ (cf. solid lines and inset). The fitted scaling factor depends on the system parameters [cf. Eq. (27)] and read $\epsilon_{A}=6.7\times 10^{-7}$, $\epsilon_{B}=6.7\times 10^{-7},\epsilon_{C}=1.0\times 10^{-7}$, and $\epsilon_{D}=1.3\times 10^{-8}$ for parameter sets A, B, C, and D, respectively. The extrapolation of all these calculations are valid as long as $D_{N}\lesssim 1$.
Due to the bosonic nature of the emitted photons, two perfectly indistinguishable photons will bunch on either output port of the BS, resulting in a vanishing HOM correlation $G_{\rm HOM}^{(2)}(\tau)$ at zero time delay $\tau\sim 0$. This is perfect HOM interference, and any deviation from it (assuming an ideal BS) can be used to quantify the distinguishability of the generated single-photons.
A standard figure of merit of indistinguishability that accounts for losses and other imperfections is the normalized HOM correlation function $g_{\rm HOM}^{(2)}[0]$ Fischer et al. (2016)
$g_{\rm HOM}^{(2)}[0]=\frac{\int_{0}^{1/(2R)}d\tau G_{\rm HOM}^{(2)}(\tau)}{\prod_{k=1}^{2}\left(\int_{0}^{1/(2R)}dt\langle b_{BS}^{k{\dagger}}(t)b_{BS}^{k}(t)\rangle\right)}.$ (25)
In analogy to Eq. (23), the numerator of Eq. (25) corresponds to the area below the zeroth peak and the normalization to the area of one of the high peaks at the repetition times $\tau=n/R$. The indistinguishability of two SPSs can be simply defined as $I=1-g^{(2)}_{\rm HOM}[0].$
We have computed $G_{\rm HOM}^{(2)}(\tau)$ as a function of $\tau$, for a setup with two emitters and a train of 4 excitation pulses [cf. Fig. 5(b)]. Here, we observe clear peaks at the repetition times $\tau=n/R$ due to the detection of two nearly independent photons coming from two different pulses. The correlations $G_{\rm HOM}^{(2)}(\tau)$ are strongly suppressed at the origin, $\tau\sim 0$, showing only a minor zeroth peak which is enlarged in Fig. 5(c). The small area of this zeroth peak manifests the high indistinguishability of the single-photons prepared in our setup with parameter set C. We also see that this behaviour is nearly insensitive to the coupling $g/\kappa$, but may strongly depend on the decoherence parameters.
Figures 5(d)-(e) display $g^{(2)}_{\rm HOM}[0]$ as function of the decoherence rates $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma.$ Note that in our implementation with parameters B and C we reach values above $I=0.99$, as inferred from Fig. 5(e) and Table 1. We also find that the distinguishability $g^{(2)}_{\rm HOM}[0]$ is quite insensitive to $\gamma_{\rm loss}/\gamma$, but it dramatically increases with dephasing, following the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al. (2003). This contrasts with the behaviour of $D_{2}$ in Fig. 3(e), illustrating that the dependence error $D_{2}$ of two SPSs is qualitatively different from the distinguishability $g^{(2)}_{\rm HOM}[0]$ measured by HOM interference. Consequently, both figures of merit have to be considered when designing high-performance multiplexed and synchronized SPSs.
IV Scalability of synchronized multiphoton generation
We are now in position to discuss the most important results of the present work: the scalability performance of the multiphoton emitter. To do so, we analyze the $N$-photon generation probability $P_{N}$ and the demultiplexing error $D_{N}$ as a function of $N$ and the various system parameters. From these results we demonstrate the feasibility to synchronize $N\sim 100-1000$ emitters with high $N$-photon generation rates $C_{N}=RP_{N}$.
We calculate $P_{N}$ in Eq. (19) by solving the master equation (12), which includes all qubit-qubit correlations induced by the cavity ($\sim J_{\rm cm},\gamma_{\rm cm}$), as well as the decoherence effects ($\sim\gamma_{\phi},\gamma_{\rm loss}$). The Hilbert space of the system grows exponentially as $2^{N}$, but using quantum trajectories (QT) [cf. Appendix B.2] we can estimate the photon statistics of the multiphoton emission and thereby $P_{N}$ up to moderately large numbers of SPSs. Figure. 6(a) displays $P_{N}$ as a function of $N$ computed from an average over $M=6000$ trajectories and considering the four parameter sets of Table 1. In the four cases and up to $N=12$, we numerically confirm that the SPSs are nearly perfectly synchronized and independent, satisfying $P_{N}\approx({\cal P}_{1})^{N}$ within the the statistical error $\Delta P_{N}\sim 10^{-3}$ [cf. black lines in Fig. (6)]. The reduction of $N$-photon generation probability with $N$ is thus mainly limited by the imperfections in the single-photon efficiency ${\cal P}_{1}$ and not by the synchronizing and demultiplexing scheme. For the decoherence-free parameters A, we predict a $12$-photon efficiency as high as $P_{12}\approx 0.97$, and when including realistic decoherence sources as in parameter set C [cf. inset in Fig. 6(a)], this reduces only to $P_{12}\approx 0.88$. For a pulsed excitation of the SPSs with a repetition rate $R=\gamma/15$ (as in Secs. III.4-III.5), these high efficiencies imply $12$-photon generation rates $C_{12}=RP_{12}$ of $\sim 0.1-1$MHz [cf. Fig. 6(b) and inset]. Since parameter set D has the largest antenna decay $\gamma/2\pi\sim 3$MHz, it shows the largest $N$-photon generation rates $C_{N}$ despite having the lowest efficiencies $P_{N}$ and the largest decoherence rates.
To quantify more precisely the deviations from the ideal scaling $P_{N}\approx({\cal P}_{1})^{N}$ in Figs. 6(a)-(b), we compute the $N$-photon dependence or demultiplexing error $D_{N}$ as defined in Eq. (20). For $N=2$ we know this error is as small as $D_{2}\sim 10^{-7}$ [cf. Figs. 3] and therefore the QT calculations with an uncertainty $\Delta P_{N}\sim 10^{-3}$ do not provide enough precision. To address this problem, we developed a photon counting approach based on the master equation [cf. Appendix B.1], which does not suffer from any statistical uncertainty. In this alternative method we simulate the photon counters at each antenna by an additional two-level system. This increases the Hilbert space dimension to $2^{2N}$ and thus limits the numerical computations to a maximum of $N\sim 5$ emitters.
In the inset of Fig. 6(c) we show the results of $D_{N}$ as a function $N$ for the four parameter sets of Table 1, and up to $N=5$. We confirm that $D_{N}<10^{-5}$ is well below the QT precision, and most importantly, we observe that the data is very well approximated by the quadratic fit
$D_{N}=\epsilon N(N-1),$ (26)
where the scaling factor $\epsilon\ll 1$ depends on the system parameters [cf. inset of Fig. 6(c)]. To analyze how the cavity-induced correlations influence $\epsilon$ and thus the scalability of the multiphoton scheme, we vary the system parameters around parameter set B and extract $\epsilon$ from fits to numerical calculations of $D_{N}$ up to $N=5$. The results of this analysis are represented by the data points in Fig. 7(a), which correspond to independent sweeps over each of the system parameters. Remarkably, the apparently complicated behaviour of the system is very well approximated by the analytical formula [cf. solid lines in Fig. 7(a)],
$\epsilon\approx(1.42)\frac{g^{4}}{\gamma^{2}}\frac{\kappa}{\left[(\kappa/2)^{2}+\Delta^{2}\right]^{3/2}},$ (27)
which manifests the dominant dependence on $g$, $\gamma$, $\kappa$, and $\Delta$, and omits the marginal effects of the effective driving $\Omega_{\rm cm}$, and the decoherence rates $\gamma_{\rm loss}$ and $\gamma_{\phi}$. This confirms our intuition that the synchronization and demultiplexing performance of our scheme is limited exclusively by residual cavity-induced correlations, provided $J_{\rm cm},\gamma_{\rm cm}\ll\gamma$ [cf. Eq. (17)].
Although Eq. (27) was extracted from a variation of parameters around set B of Table 1, this expression accurately predicts the scaling factors $\epsilon$ obtained from the fits on all parameters sets A-D in Fig. 6(c) with an error smaller than $8\%$. Interestingly, we can exploit Eq. (27) to predict new parameter sets whose demultiplexing error grows quickly to $D_{N}\gtrsim 1$ for moderate system sizes $N\lesssim 10$. Setups with these characteristics can then be used to study the validity limits of the approximated $D_{N}$ scaling in Eq. (26). In Fig. 7(b) we illustrate this analysis by calculating $D_{N}$ for two parameter sets similar to B, but: (i) with a $10$-fold coupling enhancement $g/2\pi=2$MHz (blue data), and (ii) with a $1/2$ reduction of decay $\gamma/2\pi=0.2$MHz in addition to the same $10$-fold increase in $g$ (red data). We perform $D_{N}=\epsilon N(N-1)$ fits to both data sets up to $N=5$, and we clearly observe the on-set of deviations from this scaling only when $D_{N}\gtrsim 1$ [cf. solid red line in Fig. 7]. For $D_{N}\ll 1$, both data sets show excellent agreement with the quadratic scaling (26), and this constitutes the range of validity of this approximation.
Figure 7: Limits on the scalability of the synchronized multiphoton generation. (a) Scaling factor $\epsilon$ as a function of various system parameters. The dimensionless variable $x=1$ corresponds to the parameters of set B in Table 1, which are denoted by $\tilde{g}$, $\tilde{\gamma}$, $\tilde{\Delta}$, $\tilde{\kappa}$, $\tilde{\Omega}_{\rm cm}$, $\tilde{\gamma}_{\rm loss}$, $\tilde{\gamma}_{\phi}$, and in this plot we change each of them from $25\%$ to $400\%$. We use these results to numerically extract Eq. (27). (b) Demultiplexing error $D_{N}$ calculated up to $N=10$ with QT averaged over $M=10^{4}$ trajectories. Both parameter sets are similar to B, except for $g/2\pi=2$MHz (blue data) and $g/2\pi,\gamma/2\pi]=[2,0.2]$MHz (red data). Solid lines correspond to fits of $D_{N}=\epsilon N(N-1)$ up to $N=5$, which result in the $\epsilon$ factors shown in the legend. For both data sets we observe that the quadratic scaling is a good approximation as long as $D_{N}\ll 1$, confirming the expected limit of validity.
[…]
V Inhomogeneity effects
So far, we have analyzed the performance of the scheme assuming a perfectly homogeneous setup. In large scale implementations, however, the SPS parameters will unavoidably present some degree of inhomogeneity. In this last section we analyze the impact of these imperfections.
We consider inhomogeneity deviations on all qubit parameters, which we generically denote by
$y_{j}=y+\delta y_{j},$ (28)
where $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$ are the average qubit parameters discussed in the previous sections [cf. Table 1] and $\delta y_{j}$ are random deviations over each of them. To statistically describe each of these deviations $\delta y_{j}$, we assume that they are distributed according to a Gaussian probability distribution,
$\Pi(\delta y_{j})=\frac{1}{\sqrt{2\pi(\Delta y)^{2}}}{\rm exp}\left(-\frac{1}{2}\left[\frac{\delta y_{j}}{\Delta y}\right]^{2}\right).$ (29)
Here, $\Delta y$ denotes the standard deviation associated to the disorder on each qubit parameter $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$.
To quantify the impact of each of the inhomogeneities $\delta y_{j}$ on the performance of the multiphoton emitter, we compute the average $N$-photon efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ as
$\langle\!\langle P_{N}\rangle\!\rangle_{y}=\frac{1}{M}\sum_{m=1}^{M}P_{N}[m]_{y},$ (30)
where the probabilities $P_{N}[m]_{y}$ are calculated with the master equation method of Appendix B.1 for each realization $m=1,\dots,M$ of the independent disorder $y$. The results of these computations for $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ in the case of disorder on the antenna couplings $\Delta\gamma$, as well as in the decoherence rates $\Delta\gamma_{\phi}$ and $\Delta\gamma_{\rm loss}$, are shown in Fig. (8). We perform calculations up to $N=4$ slightly inhomogeneous photon sources, for parameters B and C, and we do not observe any detrimental effect up to a disorder strength of $10\%$ of the average values [cf. Fig. 8(a)-(b)]. This is not surprising in the case of inhomogeneous decoherence rates $\gamma_{\phi}^{j}$ and $\gamma_{\rm loss}^{j}$ since their effects have nothing to do with the synchronization dynamics and are very marginally small anyways $\gamma_{\phi}^{j},\gamma_{\rm loss}^{j}\ll\gamma$. In the case of inhomogeneous antenna couplings $\gamma_{j}=\gamma+\delta\gamma_{j}$, they imply slightly different emission time scales for the emitted single-photons,
$t\gg\frac{1}{\gamma+\delta\gamma_{j}}=\frac{1}{\gamma}\left(1-\frac{\delta\gamma_{j}}{\gamma}+{\cal O}^{2}\right),$ (31)
but since we consider a long average waiting time $t=15/\gamma$, the effect of $10\%$ inhomogeneity in $\delta\gamma_{j}/\gamma$, is negligible on that timescale.
Figure 8: Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{y}$, as a function of disorder in antenna decay $y=\Delta\gamma/\gamma$, qubit dephasing $y=\Delta\gamma_{\phi}/\gamma_{\phi}$, and waveguide loss $y=\Delta\gamma_{\rm loss}/\gamma_{\rm loss}$. (a)-(b) Up to a disorder strength of $10\%$, we do not observe any impact of these three disorder sources on the synchronization and demultiplexing dynamics, as calculated up to $N=4$ for parameter sets B and C, respectively. All results are obtained from averaging $M=10^{3}$ realizations for each type of disorder.
On the other hand, inhomogeneities in the qubit frequencies $\delta\omega^{j}_{q}$ and in the couplings $\delta g_{j}$ are much more harmful for the multiphoton synchronization performance and therefore they need to be controlled more precisely in an experimental implementation of the device. Inhomogeneous couplings $g_{j}$, in particular, induce different cavity-mediated driving strengths $\Omega^{j}_{\rm cm}=\Omega_{\rm cm}+|\alpha_{\rm ss}|\delta g_{j}$, and therefore different times to realize an exact $\pi$-pulse on each qubit $\sim\pi/(2\Omega_{\rm cm}^{j})$. Since we control only the global duration of the cavity pulse, we optimally set it to the average $\pi$-pulse time $T=\pi/(2\Omega_{\rm cm})$, but this unavoidably leads to slightly different probabilities $P_{e}^{j}$ of preparing the excited states $|e\rangle_{j}$ on each qubit. Explicitly, we have
$P_{e}^{j}\approx\sin^{2}\left(\frac{\pi}{2}\frac{\Omega_{\rm cm}^{j}}{\Omega_{\rm cm}}\right)=1-\frac{\pi^{2}}{4}\left(\frac{\delta g_{j}}{g}\right)^{2}+{\cal O}^{4},$ (32)
and therefore a low coupling disorder $\delta g_{j}\ll g$ is required for a high performance of the synchronized multiphoton emitter. To precisely quantity the deviation from the ideal photon independence condition, we define the average dependence or demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{y}$ over disorder $y$ as
$\langle\!\langle D_{N}\rangle\!\rangle_{y}=\frac{\langle\!\langle P_{N}\rangle\!\rangle_{y}}{\left(\langle\!\langle{\cal P}_{1}\rangle\!\rangle_{y}\right)^{N}}-1.$ (33)
Here, the average single-photon efficiency reads,
$\langle\!\langle{\cal P}_{1}\rangle\!\rangle_{y}=\frac{1}{MN}\sum_{j=1}^{N}\sum_{m=1}^{M}{\cal P}_{1}^{j}[m]_{y},$ (34)
with ${\cal P}_{1}^{j}[m]_{y}$ calculated from $m=1,\dots,M$ realizations of the disorder $y$ and for each of the $j=1,\dots,N$ SPSs in the setup. In Figs. 9(a)-(b) we compute $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=4$ for parameter set B, and disorder strengths $\Delta g/g=0\%$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green). To achieve reasonable statistical errors $\Delta D_{N}\sim 10^{-6}$, these calculations require up to $M=3\cdot 10^{5}$ realizations. We observe that the $N$-photon efficiency and dependence, $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$, are minimally modified as long as $\delta g_{j}/g\lesssim 1\%$. Fortunately, current superconducting circuit technology allows for less than $1\%$ of geometrical errors in device fabrication, which is thus enough to suppress the disorder effects in the design of capacitive qubit-cavity couplings $g_{j}$.
Figure 9: Impact of inhomogeneities in qubit-cavity couplings $\delta g_{j}$ and qubit frequencies $\delta\omega_{j}$ on the performance of the synchronized multiphoton emitter. (a) Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ up to $N=4$ for coupling disorder strengths $\Delta g/g=0$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green). (b) Average demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=3$, for the same disorder values as in (a). (c)-(d) Analogous calculations as in (a)-(b), but considering $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ with disorder in the qubit frequencies $\delta\omega^{j}_{q}$. We consider disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green). All calculations are performed using the parameter set B of Table 1 and up to $M=3\cdot 10^{5}$ realizations of each type of disorder.
Finally, we analyze the impact of inhomogeneous qubit frequencies $\omega_{q}^{j}=\omega_{q}+\delta\omega^{j}_{q}$, which lead to slightly different qubit detunings with respect to the cavity driving frequency $\omega_{d}$, i.e. $\delta_{j}=\delta-\delta\omega_{q}^{j}$, where $\delta=\omega_{d}-\omega_{q}$ is the average detuning. We can set $\omega_{d}$ to compensate for the cavity-induced Lamb-shift $\delta=\delta_{\rm cm}$, but the remaining inhomogeneities will lead to slightly different probabilities of preparing the $|e\rangle_{j}$ states on each qubit, i.e.
$P_{e}^{j}\approx 1-\frac{1}{4}\left(\frac{\delta\omega^{j}_{q}}{\Omega_{\rm cm}}\right)^{2}+{\cal O}{\color[rgb]{0,0,0}\left(\frac{\delta\omega^{j}_{q}}{\Omega_{\rm cm}}\right)^{4}.}$ (35)
Therefore, frequency inhomogeneties should also be controlled $\delta\omega_{q}^{j}\ll\Omega_{\rm cm}$, in order to have a high quality synchronization and demultiplexing. We discussed in Sec. III.1 that the qubit frequencies can be fine-tuned by sending specific DC currents on each antenna channel. With this we can make all qubits nearly resonant up to a tuning imprecision in the range $\Delta\omega_{q}/2\pi\lesssim 1$MHz DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Arrangoiz-Arriola et al. (2019). This means that a frequency disorder of order $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ is achievable with state-of-the-art technology. In Figs. 9(c)-(d), we calculate $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ up to $N=4$ and for frequency disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 0\%$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ (green). These results confirm that under these realistic disorder conditions the $N$-photon generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ is minimally altered with respect to the homogeneous prediction [cf. blue data in Fig. 9(c)]. Moreover, we observe that the demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ reduces with higher frequency disorder $\Delta\omega_{q}/\Omega_{\rm cm}$ due to the larger independence of the photon emission processes. This occurs at the expense of reducing the generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ so it is not a good limit for our purpose.
The synchronized multiphoton generation scheme is thus resilient to moderate disorder in all system paremeters and we expect high $N$-photon generation efficiencies in realistic circuit-QED implementations of the device.
VI Conclusions and Outlook
In this work, we propose a scalable design for generating efficiently a large number of synchronized, independent and indistinguishable photons propagating over independent channels. The synchronization is provided by a strongly driven resonator in the bad cavity and weak-coupling limits. The resonator allows for a simultaneous and robust control of all emitters, which deposit photon into individual waveguides, at a high rate and with a high collection efficiency. Although our scheme can be implemented in cavity QED and nanophotonic platforms Welte et al. (2018); Casabone et al. (2015); Araneda et al. (2018); Reitz et al. (2013); Lodahl et al. (2015); Antón et al. (2019); Ellis et al. (2018); Goban et al. (2015), we have discussed a particularly efficient superconducting circuits implementation, with microwave resonators, superconducting qubits and microwave transmission lines Pechal et al. (2014); Zhou et al. (2020); Wang et al. (2020); Barends et al. (2013); Caldwell et al. (2018); McKay et al. (2019); Andersen et al. (2019); Dassonneville et al. (2020a).
The only intrinsic limitation for the scalability of the synchronized multiphoton device are the cavity-mediated interactions and collective decay, which can create correlations between SPSs and the emitted photons. Nevertheless, for state-of-the-art circuit QED parameters, we show that these correlations are strongly suppressed on the timescale of photon emission, and that they only induce a nearly negligible quadratic demultiplexing error $D_{N}\approx\epsilon N(N-1)$ with $\epsilon\sim 10^{-8}-10^{-6}$. Remarkably, this allows for the synchronization of up to hundreds of nearly independent single-photons, even in the presence of decoherence and disorder.
Given that, as we show, each SPS can achieve single-photon efficiency, purity, and indistinguishability above $99\%$, and the parallel operation in our device enables the efficient creation of large $N$-photon states. For instance, we predict a $30$-photon probability of $P_{30}\approx 0.72$ at a rate $C_{30}\approx 270$kHz and a $100$-photon probability of $P_{100}\approx 0.33$ at a rate $C_{100}\approx 130$kHz. This is seven orders of magnitude more efficient than the most sophisticated multiplexed SPSs up to $N=14$ Lenzini et al. (2017); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a); Antón et al. (2019); Hummel et al. (2019). All these promising figures of merit can be further improved when implementing a more refined model for the SPSs such as three-level emitters Fischer et al. (2016), but this is independent from the efficient synchronization scheme we propose.
The realization of scalable and deterministic sources of multi-photon states opens the door to large-scale quantum information processing with photons such as boson sampling Brod et al. (2019), but also quantum networking Llewellyn et al. (2020), simulation Sparrow et al. (2018), and computation Takeda and Furusawa (2019). Moreover, the setup and ideas introduced in this work can be further extended to generate correlated multi-photon states Meyer-Scott et al. (2019); Pfaff et al. (2017) with engineered entanglement patterns. Many-body methods such as Matrix-product-states Huang et al. (2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Sánchez-Muñoz et al. (2018); Dassonneville et al. (2020b); Lescanne et al. (2020); Jeannic et al. (2020) could be use to characterize them in the lab.
Acknowledgements
The authors thank S. Paraoanu, F. Luis, R. Martínez, R. Dassonneville, O. Buisson, Z. Wang, E. Torrontegui, M. Pino, A. González-Tudela, and D. Porras, for helpful discussions. Work in Madrid was supported by project PGC2018-094792-B-I00 (MCIU/AEI/FEDER, UE), CAM/FEDER project No. S2018/TCS-4342 (QUITEMAD-CM), and CSIC Quantum Technology Platform PT-001. T.R. further acknowledges funding from the EU Horizon 2020 program under the Marie Skłodowska-Curie grant agreement No. 798397. M.L. acknowledges support from the China Scholarship Council and National Natural Science Foundation of China under Grant No. 11475021.
Appendix A Derivation of effective model for cavity synchronization
In this appendix we outline the derivation of the effective master equation (12) by performing a displacement transformation on the cavity mode [cf. Sec. A.1], and then an adiabatic elimination of the cavity fluctuations [cf. Sec. A.2].
A.1 Coherent displacement of driven cavity mode
The quantum Langevin equations Gardiner and Zoller (2004) associated to the master equation (6) read
$\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-i\Omega-\sqrt{\kappa}a_{\rm in},$ (36)
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}a\sigma_{j}^{z}$ (37)
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})+2ig_{j}(a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}a)$ (38)
$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$
Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}(t)$ corresponds to the input noise field of the cavity mode, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$. A coherent driving on the cavity induces a coherent state component on this mode, and thus it is convenient to displace it using the transformation $a\rightarrow\alpha+\delta a$ as in Eq. (7). Here, $\delta a$ corresponds to the quantum fluctuation of the cavity mode around its classical value $\alpha$. By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation reads
$\delta\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]\delta a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-\sqrt{\kappa}a_{\rm in}(t),$ (39)
and for the qubits,
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}\delta a\sigma_{j}^{z}+ig_{j}|\alpha|\sigma_{j}^{z}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (40)
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})+2ig_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}\delta a)$
$-2ig_{j}|\alpha|(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}(t)+{\rm h.c.}\right].$ (41)
Here, $\alpha(t)$ is determined via the classical differential equation (8). Notice that, without loss of generality, we have absorbed the phase of $\alpha$ in the definition of all raising and lowering operators, so that only the magnitude $|\alpha|$ has a physical effect in the dynamics of the Langevin equations (39)-(41).
The master equation associated to the above Langevin equations (39)-(41) with the displaced cavity reads
$\dot{\rho}^{\prime}(t)={}$ $-i[H^{\prime}(t),\rho^{\prime}]+\kappa{\cal{D}}[\delta a]\rho^{\prime}+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}$
$+\sum_{j=1}^{N}\gamma^{j}_{\rm loss}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\rho^{\prime},$ (42)
where the system Hamiltonian after the displacement transformation $H^{\prime}(t)$ is given by,
$H^{\prime}(t)={}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
$+\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (43)
Notice that the master equation (42) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (43), which acts on the qubits rather than on the cavity, i.e.
$\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
Another important advantage of the displaced master equation (42) is its efficiency for numerical simulations in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$, where most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is very weakly populated. This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits.
A.2 Adiabatic elimination of cavity fluctuations
In the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state $\alpha_{\rm ss}=-i\Omega/(\kappa/2-i\Delta)$, with the quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$. In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only. To do so, we formally integrate Eq. (39) and apply the Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$. As a result, we get
$\delta a(t)\approx-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t).$ (45)
Replacing the above expressions in the original quantum Langevin equations (37)-(38), we obtain the effective dynamics of the qubits, given by
$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss}+\gamma_{\rm cm}^{j})/2-i(\delta_{j}-\delta^{j}_{\rm cm})\right]\sigma_{j}^{-}$ (46)
$+i\Omega_{\rm cm}^{j}f(t)\sigma_{j}^{z}+\sum_{l\neq j}\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{z}\sigma_{j}^{-}+\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{z}a_{\rm in}$
$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$
$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j}+\gamma_{\rm cm}^{j})(1+\sigma_{j}^{z})-2i\Omega_{\rm cm}^{j}f(t)(\sigma_{j}^{+}-\sigma_{j}^{-})$
$-2\sum_{l\neq j}\left[\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{+}\sigma_{l}^{-}+{\rm h.c.}\right]$
$-2\sqrt{\gamma^{j}_{\rm cm}}\left[\sigma_{j}^{+}a_{\rm in}+{\rm h.c.}\right]-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]$
$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right],$ (47)
where the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (10) and (14)-(16). Without loss of generality, we have reabsorbed the constant phase $ie^{i\Theta}$, with $\Theta={\rm Arctan}(2\Delta/\kappa)$ in all annihilation operators.
Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs. (46)-(47). This effective master equation after adiabatic elimination of the cavity mode is given in Eqs. (12)-(13).
Appendix B Photon counting and calculation of photon generation probabilities
Photon detection and counting lies at the heart of Quantum Optics Gardiner and Zoller (2004); Plenio and Knight (1998) and the study of light-matter interactions Meystre (2006); Obukhov et al. (2012), and thus plenty of methods have been developed over time. In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs. In Sec. B.1 we introduce an original photon counting method that we developed based on extending the master equation formalism. Then, in Sec. B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach. This is useful when the Hilbert space of the system becomes very large, but at the expense of losing precision in the computation of the averages.
B.1 Photon counting in master equation formalism
In the master equation dynamics the information about the emission of photons into the bath is omitted, and therefore one typically resorts to the input-output formalism (cf. Sec. C and Ref. Gardiner and Zoller (2004)) in order to relate measurable photonic quantities to multi-time system correlations. Nevertheless, obtaining the photon statistics from these system correlations involves the computation of multi-dimensional integrals over time, which can be very computationally costly and inefficient when scaling up the number of emitters or photons to probe.
To tackle the above problem, we developed a non-conventional photon counting method that extends the master equation formalism by reincorporating the information of the emitted photons. To do so, we simulate photon counters at each output channel $j=1,\dots,N$ of the system as quantum “boxes” that dynamically count the number of quantum jumps performed by each emitter and thereby the emitted photons. An adequate modeling of the photon counters is crucial to ensure that their presence does not alter the physical dynamics of the system and this is what we detail in the following. First, our method requires extending the Hilbert space of the system ${\cal H}_{\rm sys}$ as
${\cal H}_{\rm ext}={\cal H}_{\rm sys}\otimes{\cal H}^{\rm c}_{1}\otimes\dots\otimes{\cal H}^{\rm c}_{N},$ (48)
[…]
${\rm dim}\{{\cal H}_{\rm ext}\}={\rm dim}\{{\cal H}_{\rm sys}\}(1+N_{c})^{N},$ (49)
but as long as the total dimension fits in ${\rm dim}\{{\cal H}_{\rm ext}\}\lesssim 2^{10}$ our method provides a fast and efficient way to numerically calculate the few-photon statistics of multiple emitters within a purely master equation approach. For any Markovian master equation for a system state $\tilde{\rho}(t)$, one can find an extended master equation that incorporates the few-photon counting statistics in the dynamics. The general recipe is actually very simple and thus we explain it directly on the synchronized multiphoton device described by Eq. (12). In this case, the extended master equation reads
$\dot{\rho}_{\rm ext}={}$ $-i[\tilde{H}(t),\rho_{\rm ext}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}S_{j}^{\dagger}]\rho_{\rm ext}$
${}+$ $\sum_{j=1}^{N}\gamma^{j}_{\rm loss}{\cal{D}}[\sigma_{j}^{-}]\rho_{\rm ext}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\rho_{\rm ext}$
$+{\cal{D}}\left[\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}\right]\rho_{\rm ext},$ (50)
where $\rho_{\rm ext}$ is the density operator of the extended system including counters. Importantly, the extended master equation looks identical to the original in Eq. (12), except for the inclusion of the counting operators $S_{j}^{\dagger}$ in the Lindblad terms $\sim\gamma_{j}{\cal{D}}[\sigma_{j}^{-}S_{j}^{\dagger}]$ associated to the photon emissions we want to characterize. The counting operator is a cyclic and unitary operator defined as
$S_{j}^{\dagger}=\sum_{n=0}^{N_{c}-1}|n+1\rangle_{j}\langle n|+|0\rangle_{j}\langle N_{c}|,$ (51)
and therefore every time the emitter $j$ performs a quantum jump and decays into its antenna with rate $\gamma_{j}$, $S_{j}^{\dagger}$ adds a new photon to the counter box $j$ as
$S_{j}^{\dagger}|n\rangle_{j}={}$ $|n+1\rangle_{j},\quad n\neq N_{c}.$ (52)
When a counter reaches its maximum state $|N_{c}\rangle_{j}$, additional quantum jumps would reset the counter as,
$S_{j}^{\dagger}|N_{c}\rangle_{j}={}$ $|0\rangle_{j},$ (53)
and therefore it is very important to choose $N_{c}$ large enough to avoid reaching this limit and properly account for the physically emitted photons. Using the cyclic definition of $S_{j}^{\dagger}$ in Eq. (51), as well as their unitarity properties $S_{j}^{\dagger}S_{j}=S_{j}S_{j}^{\dagger}=1$, we can take partial trace on the extended master equation (50) and show that we exactly recover the system dynamics as
$\tilde{\rho}(t)={\rm Tr}_{\rm c}\{\rho_{\rm ext}(t)\},$ (54)
where $\tilde{\rho}(t)$ is the state of the system governed by the original master equation without photon counting (12).
In practical calculations we therefore solve for the extended state $\rho_{\rm ext}(t)$ in Eq. (50), and we then obtain the few-photon statistics of the emitted photons by taking simple expectation values on $\rho_{\rm ext}(t)$. In particular, the probability ${\cal P}_{n}^{j}$ to count $n$ photons in channel $j$ is calculated as
${\cal P}_{n}^{j}(t)={\rm Tr}\{\Lambda_{n}^{j}\rho_{\rm ext}(t)\},$ (55)
where the projection operators $\Lambda_{n}^{j}$ on the counter Fock states $|n\rangle_{j}$ are defined as
$\Lambda_{n}^{j}=|n\rangle_{j}\langle n|.$ (56)
More generally, the probability to detect $n_{1},\dots,n_{N}$ photons in output channels $j=1,\dots,N$, respectively, is obtained by products of the projectors as
$\Pi_{n_{1}\dots n_{N}}^{1\dots N}={\rm Tr}\{\Lambda_{n_{1}}^{1}\dots\Lambda_{n_{N}}^{N}\rho_{\rm ext}\}.$ (57)
Using equation (57) we can calculate the full few-photon statistics of the system as long as the evolution of the extended master equation (50) is numerically tractable. This method is particularly suited for characterizing the efficiency of SPSs since in that case we expect the photon statistics to be strongly peaked at $n_{j}\approx 1$ and therefore taking $N_{c}=1$ or $N_{c}=2$ on each counter may be enough. In most calculations shown in the main text, we take $N_{c}=1$ and quantify the probability of emitting one photon on each of the $N$ channels simultaneously by computing
$P_{N}=\Pi_{1\dots 1}^{1\dots N}={\rm Tr}\{\Lambda_{1}^{1}\dots\Lambda_{1}^{N}\rho_{\rm ext}\}.$ (58)
In the general case, however, it is important to ensure that the occupation of the last state of the counters is negligible, so that the photon statistics is not affected by the finite size of the counters.
B.2 Photon counting in quantum trajectories formalism
The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al. (1992); Molmer et al. (1993); Dum et al. (1992). Here, the physics of quantum jumps is explicitly simulated during the open system evolution and therefore it is very natural to count them and thereby infer the photon statistics.
The QT interpretation requires re-expressing the master equation (12) of our SPS synchronization and demultiplexing system as Plenio and Knight (1998); Daley (2014)
$\dot{\tilde{\rho}}(t)=-i(H_{\rm nh}\tilde{\rho}-\tilde{\rho}H_{\rm nh}^{\dagger})+\sum_{q=1}^{Q}c_{q}\tilde{\rho}c_{q}^{\dagger}.$ (59)
Here, $H_{\rm nh}$ is the non-Hermitian Hamiltonian of the system given by
$H_{\rm nh}=\tilde{H}-\frac{i}{2}\sum_{q=1}^{Q}c_{q}^{\dagger}c_{q},$ (60)
with $\tilde{H}$ the standard system Hamiltonian in Eq. (13) and $c_{q}$ denoting the $Q=3N+1$ jump operators appearing in the master equation (12). Using the index $j=1,\dots,N$, which describes each of the $N$ SPSs, the jump operators $c_{q}$, with $q=1,\dots,Q=3N+1$ can be decomposed as
$c_{j}={}$ $\sqrt{\gamma_{j}}\sigma_{j}^{-},$ (61)
$c_{N+j}={}$ $\sqrt{\gamma_{\rm loss}^{j}}\sigma_{j}^{-},$ (62)
$c_{2N+j}={}$ $\sqrt{2\gamma_{\phi}^{j}}\sigma_{j}^{+}\sigma_{j}^{-},$ (63)
$c_{3N+1}={}$ $\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}.$ (64)
[…]
and stochastic quantum jumps that project the quantum state at random times as
$|\Psi_{m}(t)\rangle\rightarrow\frac{c_{q}|\Psi_{m}(t)\rangle}{\langle\Psi_{m}(t)|c_{q}^{\dagger}c_{q}|\Psi_{m}(t)\rangle},$ (66)
where the specific jump operator $c_{q}$ is also randomly chosen from the $q=1,\dots,Q$ possibilities on each jump process. When solving for $M\gg 1$ realizations, one can obtain the density matrix $\tilde{\rho}(t)$ of the system from the ensemble average $\tilde{\rho}(t)=(1/M)\sum_{m=1}^{M}|\Psi_{m}(t)\rangle\langle\Psi_{m}(t)|$ or calculate the expectation value of any system operator $X$ as $\langle X(t)\rangle=(1/M)\sum_{m=1}^{M}\langle\Psi_{m}(t)|X(t)|\Psi_{m}(t)\rangle$.
Importantly, if we record the information of how many jumps of each type $c_{q}$ occurred on each of the $M$ realizations $|\Psi_{m}(t)\rangle$, we can directly access the photon statistics of the system from this data. For instance, the probability ${\cal P}_{n}^{j}$ to generate $n$ photons on antenna $j=1,\dots,N$ is calculated in the QT approach as
${\cal P}_{n}^{j}=\frac{{\cal N}(c_{j}|n)}{M},$ (67)
where ${\cal N}(c_{j}|n)$ denotes the number of trajectories that registered $n$ jumps with a given operator $c_{j}$ of Eq. (61). Similarly, the $N$-photon probability $P_{N}$ of generating one photon on each of the $N$ independent antennas can be statistically obtained as
$P_{N}=\frac{{\cal N}(c_{1},c_{2},\dots,c_{N}|1,1,\dots,1)}{M},$ (68)
where ${\cal N}(c_{1},c_{2},\dots,c_{N}|1,1,\dots,1)$ denotes the number of trajectories that registered exactly one jump of each $c_{j}$ operator in Eq. (61), for $j=1,\dots,N$.
When calculating many trajectories $M\gg 1$ of the system dynamics, we can gather enough quantum jump data and determine the photon emission probabilities (67) and (68) with low statistical error. Since this error decreases as $\sim M^{-1/2}$, we typically require on the order of $M\sim 10^{3}-10^{4}$ trajectories to obtain meaningful results with an error on order $\sim 10^{-2}-10^{-3}$. In practice, this makes the calculation of ${\cal P}^{j}_{n}$ and $P_{N}$ less precise than the extended master equation method in Sec. B.1. Nevertheless, the advantage of QT formalism is that we evolve pure states instead of density matrices and that it avoids extending the Hilbert space dimension to include counters as in the extended master equation method. This two key aspects allows us to dramatically reduce the Hilbert space for the simulations and thus to treat much larger systems composed of many more SPSs [cf. the scalability calculations in Fig. 6].
Appendix C Photon correlations and input-output formalism
In Secs. III.4-III.5, we discuss how to quantify multi-photon contamination and photon indistinguishability in the emission of SPSs via second-order photon correlation functions. These photon correlations are measured via coincidence counts either in the Hanbury Brown and Twiss (HBT) or the Hong-Ou-Mandel (HOM) configurations, and can be expressed as
$G_{\rm HBT}^{(2),j}(\tau)={}$ $\int_{0}^{\infty}\!\!\!\!dt\langle b^{j{\dagger}}_{\rm out}(t)b^{j{\dagger}}_{\rm out}(t+\tau)b^{j}_{\rm out}(t+\tau)b^{j}_{\rm out}(t)\rangle,$ (69)
$G_{\rm HOM}^{(2),jl}(\tau)\!={}$ $\!\!\!\int_{0}^{\infty}\!\!\!\!\!\!dt\langle b_{BS}^{1(jl){\dagger}}\!(t)b_{BS}^{2(jl){\dagger}}\!(t\!+\!\tau)b_{BS}^{2(jl)}(t\!+\!\tau)b_{BS}^{1(jl)}\!(t)\rangle.$ (70)
In the HBT correlations (69), the operators $b^{j}_{\rm out}(t)$ annihilates an output photon on the antenna channel $j$ at time $t$, and in HOM correlations (70), $b_{BS}^{1(jl)}(t)=[b^{j}_{\rm out}(t)+b^{l}_{\rm out}(t)]/\sqrt{2}$, and $b_{BS}^{2(jl)}(t)=[b^{j}_{\rm out}(t)-b^{l}_{\rm out}(t)]/\sqrt{2}$ correspond to the photonic output operators after passing through a beamsplitter that connects antennas two antennas $j$ and $k$.
As explained in Secs. III.4-III.5, it is convenient to define normalized second-order correlation functions at zero time delay, which in the case of pulsed emission read,
$g_{\rm HBT}^{(2),j}[0]{}$ $=\frac{\int_{0}^{\frac{1}{2R}}d\tau G_{\rm HBT}^{(2),j}(\tau)}{\left(\int_{0}^{\frac{1}{2R}}dt\langle b_{\rm out}^{j{\dagger}}(t)b_{\rm out}^{j}(t)\rangle\right)^{2}},$ (71)
$g_{\rm HOM}^{(2),jl}[0]{}$ $=\frac{\int_{0}^{\frac{1}{2R}}d\tau G_{\rm HOM}^{(2),jl}(\tau)}{\prod_{k=1}^{2}\left(\int_{0}^{\frac{1}{2R}}dt\langle b_{BS}^{k(jl){\dagger}}(t)b_{BS}^{k(jl)}(t)\rangle\right)}.$ (72)
To express the photon correlations in Eqs. (69)-(72) in terms of two-time system correlations, we can use the input-output relation, which read
$b^{j}_{\rm out}(t)=b^{j}_{\rm in}(t)+\sqrt{\gamma_{j}}\sigma_{j}^{-}(t),$ (73)
where $\sigma_{j}^{-}(t)$ is the Pauli operator of qubit $j=1,\dots,N$. In addition, the input field $b^{j}_{\rm in}(t)$ is the same operator that appears in the quantum Langevin equations (41) or (47), and it can be expressed as a Fourier transform over the annihilation operators $b_{j}(\omega)$ of photons of frequency $\omega$ propagating in antenna $j$, namely
$b_{\rm in}^{j}(t)={}$ $\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}d\omega e^{-i\omega t}b_{j}(\omega).$ (74)
If we use the input-output relation (73) into the correlations functions in Eqs. (71)-(72), and consider that all antennas are initially in vacuum state $|0\rangle$, we obtain
$g_{\rm HBT}^{(2),j}[0]{}$ $=\frac{\int_{0}^{\frac{1}{2R}}\!d\tau\!\int_{0}^{\infty}\!dt\langle\sigma^{+}_{j}(t)\sigma^{+}_{j}(t+\tau)\sigma^{-}_{j}(t+\tau)\sigma^{-}_{j}(t)\rangle}{\left(\int_{0}^{\frac{1}{2R}}dt\langle\sigma_{j}^{+}(t)\sigma_{j}^{-}(t)\rangle\right)^{2}},$ (75)
and
$g_{\rm HOM}^{(2),jl}[0]{}$ $=\frac{\int_{0}^{\frac{1}{2R}}\!d\tau\!\!\int_{0}^{\infty}\!\!dt\langle\xi_{1}^{(jl){\dagger}}\!(t)\xi_{2}^{(jl){\dagger}}\!(t\!+\!\tau)\xi_{2}^{(jl)}\!(t\!+\!\tau)\xi^{(jl)}_{1}\!(t)\rangle}{\prod_{k=1}^{2}\left(\int_{0}^{\frac{1}{2R}}dt\langle\xi_{k}^{(jl){\dagger}}(t)\xi_{k}^{(jl)}\!(t)\rangle\right)}.$ (76)
Here, the superposition system operators between qubits $j$ and $l$ read
$\xi_{1}^{(jl)}(t)=(\sqrt{\gamma_{j}}\sigma_{j}^{-}(t)+\sqrt{\gamma_{l}}\sigma_{l}^{-}(t))/\sqrt{2},$ (77)
$\xi_{2}^{(jl)}(t)=(\sqrt{\gamma_{j}}\sigma_{j}^{-}(t)-\sqrt{\gamma_{l}}\sigma_{l}^{-}(t))/\sqrt{2}.$ (78)
Finally, we can use the Eq. (12) together with the quantum fluctuation regression theorem Gardiner and Zoller (2004) to calculate the system expectation values $\langle\sigma_{j}^{+}(t)\sigma_{j}^{-}(t)\rangle$ and $\langle\xi_{k}^{(jl){\dagger}}(t)\xi_{k}^{(jl)}\!(t)\rangle$, as well as the two-time system correlation functions $\langle\sigma^{+}_{j}(t)\sigma^{+}_{j}(t+\tau)\sigma^{-}_{j}(t+\tau)\sigma^{-}_{j}(t)\rangle$ and $\langle\xi_{1}^{(jl){\dagger}}\!(t)\xi_{2}^{(jl){\dagger}}\!(t\!+\!\tau)\xi_{2}^{(jl)}\!(t\!+\!\tau)\xi^{(jl)}_{1}\!(t)\rangle$. After calculating these quantities, we replace them into Eqs. (75)-(76), and perform the corresponding integrals over time $t$ and time delay $\tau$ to obtain the results for the HBT and HOM second-order correlations functions shown in Secs. III.4-III.5.

### v2

@@
 (2019), cryptography Gisin et al.
-(2002), or boson sampling Brod et al.
+(2002), metrology Motes et al.
+(2015); Ge et al.
+(2018), boson sampling Brod et al.
 (2019); Wang et al.
@@
 (2017); Wang et al.
-(2019a).
+(2019a), or even quantum optical neural networks Steinbrecher et al.
+(2019).
 Scaling up these protocols requires the generation of a large product-state of $N\gg 1$ indistinguishable photons,
@@
 The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al.
-(2016), which can be deterministically excited He and Barkai (2006a, b) and decay spontaneously, producing individual photons that are collected into the desired channels.
+(2016), which can be deterministically excited He and Barkai (2006); He (2006) and decay spontaneously, producing individual photons that are collected into the desired channels.
 The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs.
@@
 (2016); Dusanowski et al.
-(2019); Kirsanske et al.
+(2019); Kiršanskė et al.
 (2017); Liu et al.
@@
 1(a) and (b).
-The cavity mode of frequency $\omega_{c}$ is coherently driven by a time-dependent field with frequency $\omega_{d}$, amplitude $\Omega(t)$ and detuning $\Delta=\omega_{d}-{\color[rgb]{0,0,0}\omega_{c}}.$ In a frame rotating with frequency $\omega_{d}$, the Hamiltonian of this system reads
-$H(t)={}$ $-\Delta a^{{\dagger}}a+\Omega(t)(a+a^{{\dagger}})$
-$-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}a).$ (4)
+This cavity mode of frequency $\omega_{c}$ is externally driven by a coherent field with time-dependent amplitude $\Omega(t)$ and frequency $\omega_{d}$.
+The total Hamiltonian of the system reads
+$H(t)={}$ $\omega_{c}a^{{\dagger}}a+2\Omega(t)\cos(\omega_{d}t)(a+a^{{\dagger}})$
+$+\frac{1}{2}\sum_{j=1}^{N}\omega_{j}^{q}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(a+a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+}).$ (4)
 Here, $^{{\dagger}}, aa$ are the cavity creation and annihilation operators, whereas $\sigma_{j}^{+}=|e\rangle_{j}\langle g|$, $\sigma_{j}^{-}=|g\rangle_{j}\langle e|$, and $\sigma_{j}^{z}=|e\rangle_{j}\langle e|-|g\rangle_{j}\langle g|$ are standard Pauli operators for qubits $j=1,\dots,N$, with ground and excited states denoted by $|g\rangle_{j}$, and $|e\rangle_{j}$, respectively.
-Each qubit has a possibly different frequency $\omega_{q}^{j}$ and detuning $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ with respect to the cavity drive.
-The interaction strength $g_{j}$ between the $j$-th qubit and the cavity satisfies the rotating-wave approximation (RWA) $g_{j}\ll\omega_{q}^{j}\sim\omega_{d},$ and coupling terms can be written in the standard Jaynes-Cummings form.
+Each qubit $j$ has a possibly different frequency $\omega_{q}^{j}$ and a coupling $g_{j}$ to the common cavity mode.
+Note that we do not assume rotating wave approximation (RWA) to allow for a large cavity occupation $\langle a^{\dagger}a\rangle\gg 1$.
 We control the emitters by modulating the envelope of the cavity drive
@@
 II.2 Cavity-mediated synchronization and residual correlations
-The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode, using a strong drive $|\Omega_{0}|\gg\kappa,|\Delta|$.
-We model this state using a displacement of the Fock operator
+The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode.
+This can be achieved via a very strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning.
+We model the resulting state using a displacement of the Fock operator
 $a=\alpha+\delta a,$ (7)
 where the amplitude $\alpha(t)$ is given by the classical harmonic oscillator equation
-$\dot{\alpha}(t)={}$ $-\left[\kappa/2-i\Delta\right]\alpha(t)-i\Omega_{0}f(t).$ (8)
-Once the pulse is switched on ($f(t)=1$), $|\alpha(t)|$ grows and reaches the steady state value
+$\dot{\alpha}(t)={}$ $-\left[\kappa/2+i\omega_{c}\right]\alpha(t)-2i\Omega_{0}f(t){\color[rgb]{0,0,0}\cos(\omega_{d}t)}.$ (8)
+Once the pulse is switched on ($f(t)=1$), $|\alpha(t)|$ grows and stabilizes around the steady state value
 $|\alpha_{\rm ss}|=\frac{\Omega_{0}}{\sqrt{(\kappa/2)^{2}+\Delta^{2}}}\gg 1.$ (9)
 Similarly, once we switch the drive off $(f(t\geq T)=0),$ the cavity displacement vanishes in a time scale $\sim 1/\kappa.$
-As shown in App.
-A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, described by the effective Hamiltonian, $\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j}g_{j}|\alpha(t)|(\sigma_{j}^{-}+\sigma_{j}^{+}).$ In the asymptotic limit, the large and steady cavity amplitude allows for a strong cavity-mediated driving strength on the qubits
-$\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (10)
-which we use to excite all SPSs simultaneously.
+As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
+[…]
+This is the mechanism that allows us to excite all SPSs simultaneously.
 In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed.
 To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
-$g_{j},|\delta_{j}|\ll\kappa,|\Delta|.$ (11)
-These conditions [cf.
-App.
-A.2] ensure that the cavity reaches a steady state (9) where quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated.
-The dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\},$ is then modeled by the master equation
+$g_{j},|\delta_{j}|\ll\kappa,|\Delta|{\color[rgb]{0,0,0}\ll\omega_{j}^{q},\omega_{c}},$ (11)
+where $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ is the detuning between qubit $j$ and cavity drive.
+These conditions ensure that the cavity reaches a steady state (9), where small quantum fluctuations $\delta a\ll|\alpha_{\rm ss}|$ can be adiabatically eliminated and fast oscillations neglected.
+As shown in App.
+A.2, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
 $\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\cal{D}}\left[\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}\right]\tilde{\rho}$
 $+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$, (12)
-with the effective Hamiltonian $\tilde{H}(t)$
+with the effective Hamiltonian $\tilde{H}(t)$ given by
 $\tilde{H}(t)={}$ $-\frac{1}{2}\sum_{j=1}^{N}(\delta_{j}-\delta^{j}_{\rm cm})\sigma_{j}^{z}+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}f(t)(\sigma_{j}^{-}+\sigma_{j}^{+})$
 $+\sum_{j>l}J^{jl}_{\rm cm}\left(\sigma_{j}^{+}\sigma_{l}^{-}+\sigma_{l}^{+}\sigma_{j}^{-}\right).$ (13)
-In addition to the cavity-mediated driving $\Omega_{\rm cm}^{j}$ in Eq. (10), the qubits experience a Lamb-shift
+Here, the cavity-mediated driving $\Omega_{\rm cm}^{j}$ reads
+$\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (14)
+with the steady state coherent amplitude given in Eq. (9). In addition, the qubits experience a Lamb-shift
 $\delta^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\Delta}{(\kappa/2)^{2}+\Delta^{2}},$ (15)
@@
 Third, to achieve nearly independent photons in a product state (1), the photons must be created faster than the speed of the cavity mediated correlations and interactions ($\gamma\gg J_{\rm cm},\gamma_{\rm cm}$).
-This is the only intrinsic limitation of the proposed multiphoton scheme.
+This is the only intrinsic limitation of the proposed multiphoton scheme, and its effect can be mitigated in the case of a bad cavity $g\ll\kappa$ with strong coherent drive $|\alpha_{\rm ss}|\gg 1$.
 In Sec.
@@
 Figure 2: Synchronization of two deterministic SPSs via a cavity.
-(a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity [cf.
-red dash-dotted line].
+(a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity with fast residual oscillations due to non-RWA terms [cf.
+red solid and black dashed lines].
 This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf.
-blue solid line].
+blue dash-dotted line].
 After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$.
@@
 Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf.
-orange dotted line].
-The parameters correspond to a state-of-the-art circuit-QED implementation, and are presented in row C of Table 1.
+yellow dotted line].
+(c)-(d) Deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|$ between the exact (6) and effective (12) models for the qubit occupations $x=\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$, and photon generation probabilities $x={\cal P}_{1},P_{2}$ as a function of time.
+We observe great agreement and no heating effects despite the large cavity driving strength $\Omega_{0}=4\omega_{d}\gg\kappa,|\Delta|$.
+These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row C of Table 1 and discussed in Sec.
+III.1.
 $\frac{\gamma}{2\pi}[{\rm MHz}]$ $\frac{\gamma_{\rm loss}}{2\pi}[{\rm kHz}]$ $\frac{\gamma_{\phi}}{2\pi}[{\rm kHz}]$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\phi}/\gamma$ $\gamma_{\rm loss}/\gamma$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
@@
 Fig.
-2 illustrates the two processes in the cavity-mediated synchronization method, using two sychronized SPSs with parameters that satisfy Eq.
+2(a)-(b) illustrates the two processes in the cavity-mediated synchronization method, using two synchronized SPSs with parameters that satisfy Eq.
 (18).
 Fig.
-2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|\sim 10^{4}$ [cf.
-red dash-dotted line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf.
-blue solid line].
+2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ [cf.
+red solid line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf.
+blue dash-dotted line].
 Fig.
@@
 Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf.
-blue solid line].
+blue dash-dotted line].
 To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs.
@@
 green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf.
-orange dotted line].
+yellow dotted line].
 This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq.
 (2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
-We remark that this example only illustrates the working principle of the setup.
-As we show in Secs.
-III and IV, the parameters can be further optimized to improve the efficiency and scalability of the multi-photon emitter.
+We performed the numerical simulations in Figs.
+2(a)-(b) using both the full model in Eqs.
+(4)-(6) and the effective model in Eqs.
+(12)-(17).
+They show an excellent agreement with deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf.
+Fig.
+2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim 10^{4}\gg 1$.
+This demonstrates that the strong cavity driving $\Omega_{0}\gg\kappa,|\Delta|$ does not introduce heating in the system, but only creates a large coherent state in the cavity with some fast oscillations around the mean value $|\alpha_{\rm ss}|^{2}\gg 1$ [cf.
+solid red and dashed black lines in Fig.
+2(a)].
+Since this fast oscillations have no detrimental impact in the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (12) which uses the mean steady value (9) for describing the effect of the cavity field (see App.
+A for more details on the approximations involved in the model).
 III Performance of a circuit-QED implementation
@@
 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits.
-For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency on the order of a few GHz.
-These superconducting resonators support strong drives and can hold a large number of photons in steady state Pietikäinen et al.
+For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz.
+These superconducting resonators support strong drives and can hold a large number of photons in steady state $|\alpha_{\rm ss}|^{2}\gg 1$ Pietikäinen et al.
 (2017); Jenkins et al.
-(2014); Macri et al.
-(2020).
-In our design we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al.
+(2014); Macrì et al.
+(2020), as it is required in our scheme.
+For this, we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al.
 (2017); Jenkins et al.
@@
 (18)].
+Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference.
+(a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements.
+(b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$.
+Parameters correspond to set C of Table 1.
+(c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$.
+(d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
+Other parameters correspond to set C (blue) and set B (red).
+The indistinguishability dramatically reduces with dephasing, a behaviour in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al.
+(2003).
 Finally, we note that the exact pulse shape $f(t)$ is not relevant for the synchronization dynamics as long as it approximately induces a $\pi$-pulse on the qubits.
@@
 For simplicity, our analysis assumes a homogeneous setup and thus we omit the $j$ index in $G_{\rm HBT}^{(2)}(\tau)$.
-Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference.
-(a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements.
-(b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$.
-Parameters correspond to set C of Table 1.
-(c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$.
-(d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
-Other parameters correspond to set C (blue) and set B (red).
-The indistinguishability dramatically reduces with dephasing, a behaviour in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al.
-(2003).
 In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for train of 4 excitation pulses with a repetition rate
@@
 Appendix C].
+Due to the bosonic nature of the emitted photons, two perfectly indistinguishable photons will bunch on either output port of the BS, resulting in a vanishing HOM correlation $G_{\rm HOM}^{(2)}(\tau)$ at zero time delay $\tau\sim 0$.
+This is perfect HOM interference, and any deviation from it (assuming an ideal BS) can be used to quantify the distinguishability of the generated single-photons.
 Figure 6: Scalability performance of the synchronized multiphoton generation scheme.
@@
 The extrapolation of all these calculations are valid as long as $D_{N}\lesssim 1$.
-Due to the bosonic nature of the emitted photons, two perfectly indistinguishable photons will bunch on either output port of the BS, resulting in a vanishing HOM correlation $G_{\rm HOM}^{(2)}(\tau)$ at zero time delay $\tau\sim 0$.
-This is perfect HOM interference, and any deviation from it (assuming an ideal BS) can be used to quantify the distinguishability of the generated single-photons.
 A standard figure of merit of indistinguishability that accounts for losses and other imperfections is the normalized HOM correlation function $g_{\rm HOM}^{(2)}[0]$ Fischer et al. (2016)
@@
 (2016), but this is independent from the efficient synchronization scheme we propose.
-The realization of scalable and deterministic sources of multi-photon states opens the door to large-scale quantum information processing with photons such as boson sampling Brod et al.
-(2019), but also quantum networking Llewellyn et al.
-(2020), simulation Sparrow et al.
-(2018), and computation Takeda and Furusawa (2019).
-Moreover, the setup and ideas introduced in this work can be further extended to generate correlated multi-photon states Meyer-Scott et al.
+Scalable and deterministic sources of multi-photon states will be a key resource for realizing large-scale quantum information processing such as quantum optical neural networks Steinbrecher et al.
+(2019) or fault-tolerant photonic quantum computation Takeda and Furusawa (2019).
+In the short-term, the implemenation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al.
+(2019) or quantum metrology Motes et al.
+(2015); Ge et al.
+(2018).
+Moreover, the setup and ideas introduced in this work can be further extended to scale up the generation of correlated multi-photon states Meyer-Scott et al.
 (2019); Pfaff et al.
-(2017) with engineered entanglement patterns.
+(2017) with engineered entanglement patterns Besse et al.
+(2020).
 Many-body methods such as Matrix-product-states Huang et al.
-(2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Sánchez-Muñoz et al.
+(2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Muñoz et al.
 (2018); Dassonneville et al.
 (2020b); Lescanne et al.
-(2020); Jeannic et al.
+(2020); Le Jeannic et al.
 (2021) could be use to characterize them in the lab.
@@
 The quantum Langevin equations Gardiner and Zoller (2004) associated to the master equation (6) read
-$\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-i\Omega-\sqrt{\kappa}a_{\rm in},$ (36)
-$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}a\sigma_{j}^{z}$ (37)
-$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$
-$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})+2ig_{j}(a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}a)$ (38)
-$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$
+$\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}$
+$-2i\Omega(t)\cos(\omega_{d}t),$ (37)
+$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(a+a)^{{\dagger}}\sigma_{j}^{z}$
+$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (38)
+$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})-2ig_{j}(a^{{\dagger}}+a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
+$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (39)
 Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}(t)$ corresponds to the input noise field of the cavity mode, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$.
@@
 By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation reads
-$\delta\dot{a}={}$ $-\left[\kappa/2-i\Delta\right]\delta a-i\sum_{j=1}^{N}g_{j}\sigma_{j}^{-}-\sqrt{\kappa}a_{\rm in}(t),$ (39)
+$\delta\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]\delta a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}(t),$ (40)
 and for the qubits,
-$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2-i\delta_{j}\right]\sigma_{j}^{-}+ig_{j}\delta a\sigma_{j}^{z}+ig_{j}|\alpha|\sigma_{j}^{z}$
-$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (40)
-$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})+2ig_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}-\sigma_{j}^{+}\delta a)$
-$-2ig_{j}|\alpha|(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
+$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(\delta a+\delta a^{{\dagger}})\sigma_{j}^{z}$
+$+2ig_{j}{\rm Re}\{\alpha\}\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}$
+$-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (41)
+$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-2ig_{j}(\delta a^{{\dagger}}+\delta a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
+$-4ig_{j}{\rm Re}\{\alpha\}(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
 $-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}(t)+{\rm h.c.}\right].$ (42)
 Here, $\alpha(t)$ is determined via the classical differential equation (8).
-Notice that, without loss of generality, we have absorbed the phase of $\alpha$ in the definition of all raising and lowering operators, so that only the magnitude $|\alpha|$ has a physical effect in the dynamics of the Langevin equations (39)-(41).
 The master equation associated to the above Langevin equations (40)-(42) with the displaced cavity reads
@@
 where the system Hamiltonian after the displacement transformation $H^{\prime}(t)$ is given by,
-$H^{\prime}(t)={}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
-$+\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (43)
-Notice that the master equation (42) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (43), which acts on the qubits rather than on the cavity, i.e.
-$\Omega(t)(a+a^{\dagger})\rightarrow\sum_{j=1}^{N}g_{j}|\alpha|(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
+$H^{\prime}(t)={}$ $\omega_{c}\delta a^{{\dagger}}\delta a+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
+$+\sum_{j=1}^{N}2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
+Notice that the master equation (43) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (44), which acts on the qubits rather than on the cavity as indicated in Eq.
+(10).
 This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
-Another important advantage of the displaced master equation (42) is its efficiency for numerical simulations in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$, where most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is very weakly populated.
+The displaced master equation (43) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$ because most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is only weakly populated.
 This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits.
+In addition, if we consider the cavity in the weak coupling and bad cavity limit given by Eq.
+(11), the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq.
+(44).
+On the one hand, since $g_{j}\ll\omega_{c}+\omega_{q}^{j}$ and the fluctuation is weakly populated, we can approximate the qubit-fluctuation coupling as
+$g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})\approx g_{j}(\delta a^{\dagger}\sigma_{-}+{\rm h.c.}).$ (45)
+On the other hand, the cavity-mediated driving on the qubits can also be simplified by applying RWA as
+$2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})\approx g_{j}|\alpha_{\rm ss}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi)}+{\rm h.c.}).$ (46)
+Here, $|\alpha_{\rm ss}|=\Omega_{0}/\sqrt{(\kappa/2)^{2}+\Delta^{2}}$ and $\phi={\rm Arctan}(2\Delta/\kappa)-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile.
+To demonstrate this last approximation (46), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha(0)=0$.
+Then, the solution of the classical equation (8) for $\alpha(t)$ reads
+$\alpha(t)={}$ $|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}(1-e^{-(\kappa/2-i\Delta)t})$
+$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa/2+2i\omega_{d})t}),$ (47)
+where we have also used the inequalities $|\Delta|,\kappa\ll\omega_{d}$ valid for our parameter conditions.
+The first term in (47) reaches the steady state value $\sim|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ in a timescale $\sim 1/\kappa$ and the second term causes the fast but low amplitude oscillation observed in Fig.
+2(a).
+Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss}|\gg 1$ as it is the case for our parameters.
+Since in bad cavity limit (11) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa$, we can use the steady state value $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits.
+Similarly, after the cavity reaches the steady state and it is switched off ($f(t)=0$), the dynamics in Eq.
+(8) predicts an exponential decay of $\alpha(t)$ with rate $\sim\kappa$.
+In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha(t)\approx 0$ for $f(t)=0$.
+In summary, under the above approximations we have $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}f(t)$ so that when replacing it in the last term of Eq.
+(44), and applying RWA provided $g_{j}|\alpha_{\rm ss}|\ll\omega_{d}$, we obtain Eq.
+(46).
+Replacing Eqs.
+(45)-(46) in (44), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (43) with the RWA Hamiltonian:
+$H^{\prime}(t)\approx{}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
+$+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (48)
+Here, we have defined $\Omega_{\rm cm}^{j}=g_{j}|\alpha_{\rm ss}|$ as in Eq.
+(14) and absorbed the constant phase $\phi$ in the definition of the qubit and fluctuation operators.
+The qubit detunings read $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ as in the main text.
 A.2 Adiabatic elimination of cavity fluctuations
-In the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state $\alpha_{\rm ss}=-i\Omega/(\kappa/2-i\Delta)$, with the quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$.
+As discussed in the previous subsection, in the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state, with its quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$.
 In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only.
 To do so, we formally integrate Eq.
-(39) and apply the Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$.
+(40) and apply Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$.
 As a result, we get
-$\delta a(t)\approx-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t).$ (45)
-Replacing the above expressions in the original quantum Langevin equations (37)-(38), we obtain the effective dynamics of the qubits, given by
+$\delta a(t)\approx{}$ $-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t)$
+${\color[rgb]{0,0,0}-i\sum_{j}\frac{g_{j}}{\kappa/2+i(2\omega_{d}-\Delta)}\sigma_{j}^{+}(t).}$ (49)
+Replacing the above expressions in the original quantum Langevin equations (38)-(39), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations
 $\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss}+\gamma_{\rm cm}^{j})/2-i(\delta_{j}-\delta^{j}_{\rm cm})\right]\sigma_{j}^{-}$ (50)
@@
 $-2\sqrt{\gamma^{j}_{\rm cm}}\left[\sigma_{j}^{+}a_{\rm in}+{\rm h.c.}\right]-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]$
-$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right],$ (47)
-where the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs.
-(10) and (14)-(16).
-Without loss of generality, we have reabsorbed the constant phase $ie^{i\Theta}$, with $\Theta={\rm Arctan}(2\Delta/\kappa)$ in all annihilation operators.
+$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (51)
+Here, the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (14)-(17).
 Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs.
 (50)-(51).
-This effective master equation after adiabatic elimination of the cavity mode is given in Eqs.
+This effective master equation, which incorporates the adiabatic elimination of the cavity mode, is given in Eqs.
 (12)-(13).
 Appendix B Photon counting and calculation of photon generation probabilities
-Photon detection and counting lies at the heart of Quantum Optics Gardiner and Zoller (2004); Plenio and Knight (1998) and the study of light-matter interactions Meystre (2006); Obukhov et al.
-(2012), and thus plenty of methods have been developed over time.
+Photon detection and counting lies at the heart of Quantum Optics and thus plenty of methods have been developed over time Gardiner and Zoller (2004); Plenio and Knight (1998).
 In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs.
@@
 The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al.
-(1992); Molmer et al.
+(1992); Mølmer et al.
 (1993); Dum et al.

### v3

@@
 propagating along $N$ or more channels.
-The generation of this state with high fidelity and efficiency demands the use of nearly deterministic, nearly identical and perfectly synchronized single-photon sources (SPSs), each of them producing just one photon.
+The generation of this state with high fidelity and efficiency demands the use of nearly deterministic, nearly identical, and perfectly synchronized single-photon sources (SPSs), each of them producing just one photon.
 The best on-demand SPSs for this task rely on few-level quantum systems Fischer et al.
-(2016), which can be deterministically excited He and Barkai (2006); He (2006) and decay spontaneously, producing individual photons that are collected into the desired channels.
+(2016), which can be deterministically excited He and Barkai (2006a, b) and decay spontaneously, producing individual photons that are collected into the desired channels.
 The great experimental progress in controlling single quantum systems with cavity-enhanced light-matter interactions has allowed many realizations of nearly deterministic SPSs.
@@
 (2016); Dusanowski et al.
-(2019); Kiršanskė et al.
+(2019); Kirsanske et al.
 (2017); Liu et al.
@@
 (2020).
-Currently, quantum dots in micro-pillar cavities are proving the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al.
+Currently, quantum dots in micro-pillar cavities provide the best overall numbers including generation efficiencies, distinguishability, and single-photon purity Reimer and Cher (2019); Wang et al.
 (2019c); Uppu et al.
@@
 […]
-$P_{N}=({\cal P}_{1})^{N}S_{N},$ (2)
-of the probability ${\cal P}_{1}$ of generating an individual photon on each of the $N$ channels independently, times the correlation error $S_{N}$ introduced by the circuit.
+${\color[rgb]{0,0,0}C_{N}=R}({\cal P}_{1})^{N}S_{N},$ (2)
+with ${\cal P}_{1}$ the probability of generating an individual photon pulse on each of the $N$ channels independently, and $S_{N}$ the correlation error introduced by the synchronizing circuit.
 An ideal scheme should scale exponentially $(S_{N}=1)$, creating $N$ synchronous and completely uncorrelated photons.
@@
 (3).
-This can significantly limit the achievable $N$-photon efficiencies in practical applications such as boson sampling.
+This can significantly limit the achievable $N$-photon generation rates in practical applications requiring $N\gg 1$ such as single-photon boson sampling.
 Figure 1: Cavity-QED scheme for deterministically generating a large number of synchronized and independent single-photons.
 (a) Idealized implementation with NN two-level emitters coupled to a common driven cavity mode.
-Each emitter additionally couples to independent output channels, which collect the $N$ emitted single-photons on parallel for later usage.
+Each emitter additionally couples to independent output channels, which collect the $N$ emitted single photons in parallel for later usage.
 (b) Efficient on-chip implementation using circuit-QED.
 Here, transmon qubits realize the NN two-level emitters, each of them capacitively coupled to an independent transmission-line antenna and to a common transmission-line resonator.
-Controlling the shape $\Omega(t)$ of the resonator drive, we can trigger all SPSs simultaneously and on-demand.
+Controlling the shape $\Omega(t)$ of the resonator drive, we can trigger all SPSs simultaneously and on demand.
 The NN emitted microwave single-photons are collected by the on-chip antennas with high efficiency.
@@
 […]
-$S_{N}=1+\epsilon N(N-1).$ (3)
-The correction factor $\epsilon$ stems from residual cavity-induced correlations.
-For realistic experimental conditions at the optimal working point, we find that $\epsilon\sim 10^{-8}-10^{-6}$, making the multiphoton generation scheme scalable up to $N\sim 100-1000\ll\epsilon^{-1/2}$, depending on the implementation and the noise sources.
+${\color[rgb]{0,0,0}P_{N}=({\cal P}_{1})^{N}(1+D_{N})},$ (3)
+except for a small error $D_{N}=[P_{N}-({\cal P}_{1})^{N}]/({\cal P}_{1})^{N}\ll 1$, characterizing the residual correlation or dependence between photons.
+This small deviation is shown to be quadratic in $N$:
+${\color[rgb]{0,0,0}D_{N}}=\epsilon N(N-1),$ (4)
+with the correction factor $\epsilon$ stemming from residual cavity-induced correlations.
+We work out the parameter conditions to reach $\epsilon\sim 10^{-4}-10^{-7}$, making the multiphoton generation scheme scalable up to $N\sim 100-1000\ll\epsilon^{-1/2}$, depending on the implementation and the noise sources.
+The $N$-photon generation rate is then obtained as $C_{N}=RP_{N}$ with $R$ the repetition rate.
 To show the favorable scaling of the scheme in a realistic setup, we study a circuit-QED implementation with flux-tunable transmon emitters and microwave transmission lines DiCarlo et al.
@@
 (2020).
-For a strongly driven cavity, we show it is feasible to build microwave SPSs with single-photon purity and indistinguishability above $99\%$, as measured via standard Hanbury Brown and Twiss (HBT), and Hong-Ou-Mandel (HOM) experiments.
-Most importantly, we predict an overall single-photon efficiency or brightness of ${\cal P}_{1}\gtrsim 0.99$, a synchronized $30$-photon generation efficiency of $P_{30}\sim({\cal P}_{1})^{30}\sim 0.72$, and even $P_{100}\sim 0.33$ for $N=100$ SPSs.
-This means that for pulsed microwave SPSs with a repetition rate of $R\sim 0.4$ MHz, we can achieve $30$-photon generation rates of $C_{30}\sim 300$ kHz, and $100$-photon rates of $C_{100}\sim 100$ kHz.
-Remarkably, this is more than seven orders of magnitude more efficient than state-of-the-art boson sampling experiments with up to $N=14$ photons Wang et al.
+Moreover, the use of the cavity as a common synchronizer has a low overhead compared to having and calibrating a large number of $N$ independent control lines and driving sources, which may lead to serious limitations in the wiring and the scalability of the superconducting chip Tamate et al.
+(2022).
+Using state-of-the-art parameters, we show it is feasible to build microwave SPSs with single-photon purity and indistinguishability of $99\%$, as measured via standard Hanbury Brown and Twiss (HBT), and Hong-Ou-Mandel (HOM) experiments.
+Most importantly, we predict an overall single-photon probability or brightness of ${\cal P}_{1}\gtrsim 0.99$ and the possibility to efficiently synchronize up to $N=100$ SPSs with large multi-photon probabilities such as $P_{10}\sim 0.90$, $P_{30}\sim 0.72$, and $P_{100}\sim{\color[rgb]{0,0,0}0.16}$, for $N=10$, $30$, and $100$ SPSs, respectively.
+This means that for pulsed microwave SPSs with a repetition rate of $R\sim 0.4$ MHz, we can achieve $30$-photon generation rates of $C_{30}\sim{\color[rgb]{0,0,0}700}$ kHz, and $100$-photon rates of $C_{100}\sim{\color[rgb]{0,0,0}200}$ kHz.
+Remarkably, this is more than seven orders of magnitude more efficient than state-of-the-art boson sampling experiments with up to $N=14$ single photons Wang et al.
 (2019a).
 The performance of our synchronized multiphoton generation scheme benefits from the high efficiency of the superconducting circuit implementation, as well as from the absence of switches and other optical elements.
-We also show that our predictions are robust to small inhomogeneities and disorder in the SPS parameters.
+We show that our predictions are robust to small inhomogeneities in the SPS parameters, and to the multi-mode structure of the synchronizing cavity.
+Finally, we also consider the parameter conditions to synchronize up to $N=1000$ SPSs, which may be possible in more advanced circuit QED setups with weak and highly symmetric cavity-emitter couplings Huang et al.
+(2021); Hazra et al.
+(2021).
 The paper is organized as follows.
@@
 In Sec.
-V we analyze the effect of inhomogeneities and disorder in the SPS parameters, and finally in Sec.
-VI we summarize our conclusions.
-We complement the discussion by analytical and numerical methods to quantify the photon counting and correlations, which are briefly presented in Appendices A, B, and C.
+V we analyze the effects of inhomogeneities and disorder in the SPS parameters and in Sec.
+VI the effects of the multi-mode nature of the cavity.
+Finally in Sec.
+VII we summarize our conclusions.
+We complement the discussion with analytical and numerical methods to quantify the photon counting and photon correlations, which are presented in Appendices A, B, and C.
 II Setup and synchronization
@@
 Fig.1(a) and (b)].
-We describe this qubit-antenna interaction in the the Born-Markov approximation, introducing the $\gamma_{j}$ at which a qubit deposits photons into its own antenna.
+We describe this qubit-antenna interaction in the Born-Markov approximation, introducing the $\gamma_{j}$ at which a qubit deposits photons into its antenna.
 We also introduce a photon loss rate $\gamma_{\rm loss}^{j}$ characterizing the emission of photons into any other unwanted channel.
 Finally, we consider the cavity decay rate $\kappa$ and introduce white noise dephasing rates $\gamma_{\phi}^{j}$ on each of the qubits.
-The complete dynamics of this open system is described with a master equation,
+The complete dynamics of this open system are described with a master equation,
 $\dot{\rho}(t)={}$ $-i[H(t),\rho]+\kappa{\cal{D}}[a]\rho+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho$
@@
 The common mode and the independent output channels can be integrated and scaled up.
-In this scenario, the main limitation arise from the creation of nearly identical or tuneable emitters such as quantum dots Ellis et al.
-(2018), or alternatively the trapping of many atoms Goban et al.
+In this scenario, the main limitation arises from the creation of nearly identical or tuneable emitters such as quantum dots Ellis et al.
+(2018), or the trapping of many atoms Goban et al.
 (2015) or ions Mehta et al.
@@
 In Secs.
-III and IV we analyze in detail the performance and scalability of this superconducting circuit platform.
+III and IV we analyze in detail the state-of-art parameters, the performance and the scalability of this superconducting circuit platform.
 II.2 Cavity-mediated synchronization and residual correlations
 The key mechanism to achieve an efficient cavity-mediated synchronization of the SPSs is to generate a large coherent state $|\alpha|\gg 1$ in the cavity mode.
-This can be achieved via a very strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning.
+This can be achieved via a strong drive, $|\Omega_{0}|\gg\kappa,|\Delta|$, where $\Delta=\omega_{d}-\omega_{c}$ is the cavity-drive detuning.
 We model the resulting state using a displacement of the Fock operator
@@
 Similarly, once we switch the drive off $(f(t\geq T)=0),$ the cavity displacement vanishes in a time scale $\sim 1/\kappa.$
-As shown in App. A.1, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
+As shown in App. A.2, our control of the cavity displacement translates into a cavity-mediated driving on all coupled qubits, given by the effective Hamiltonian,
 […]
 This is the mechanism that allows us to excite all SPSs simultaneously.
-In order for our cavity mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed.
+For our cavity-mediated control to succeed, the qubit back-action, cavity-induced interactions, and correlations between qubits must be suppressed.
 To do so, we restrict the couplings and detunings to keep the system in the bad-cavity limit
@@
 As shown in App.
-A.2, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
-$\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\cal{D}}\left[\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{-}\right]\tilde{\rho}$
-$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$, (12)
-with the effective Hamiltonian $\tilde{H}(t)$ given by
+A.3, the dynamics of the qubit’s reduced state $\tilde{\rho}(t)={\rm Tr}_{c}\{\rho(t)\}$ can then be modeled by an effective master equation, which in the rotating frame with the drive frequency $\omega_{d}$ reads
+$\dot{\tilde{\rho}}(t)={}$ $-i[\tilde{H}(t),\tilde{\rho}]+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+{\color[rgb]{0,0,0}\sum_{j,l=1}^{N}\gamma^{jl}_{\rm cm}\tilde{{\cal D}}[\sigma_{j}^{-},\sigma_{l}^{-}]\tilde{\rho}}$
+$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\tilde{\rho}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\tilde{\rho}$. (13)
+Here, the effective Hamiltonian $\tilde{H}(t)$ has the form,
 $\tilde{H}(t)={}$ $-\frac{1}{2}\sum_{j=1}^{N}(\delta_{j}-\delta^{j}_{\rm cm})\sigma_{j}^{z}+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}f(t)(\sigma_{j}^{-}+\sigma_{j}^{+})$
 $+\sum_{j>l}J^{jl}_{\rm cm}\left(\sigma_{j}^{+}\sigma_{l}^{-}+\sigma_{l}^{+}\sigma_{j}^{-}\right).$ (14)
-Here, the cavity-mediated driving $\Omega_{\rm cm}^{j}$ reads
+with $\Omega_{\rm cm}^{j}$ the cavity-mediated driving given by
 $\Omega_{\rm cm}^{j}={}$ $|\alpha_{\rm ss}|g_{j},$ (15)
-with the steady state coherent amplitude given in Eq. (9). In addition, the qubits experience a Lamb-shift
+with the steady state coherent amplitude $|\alpha_{\rm ss}|$ given in Eq. (10). In addition, the qubits experience a Lamb-shift
 $\delta^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\Delta}{(\kappa/2)^{2}+\Delta^{2}},$ (16)
@@
 $J^{jl}_{\rm cm}={}$ $\frac{g_{j}g_{l}\Delta}{(\kappa/2)^{2}+\Delta^{2}}.$ (17)
-The cavity also induces a collective decay or “superradiance” on the qubits, described by a collective jump operator $W_{\rm cm}=\sum_{j=1}^{N}\sqrt{\gamma^{j}_{\rm cm}}\sigma^{-}_{j}.$ The cavity-mediated decay rates
-$\gamma^{j}_{\rm cm}={}$ $\frac{(g_{j})^{2}\kappa}{(\kappa/2)^{2}+\Delta^{2}},$ (17)
-combined with the local decay and dephasing rates $\gamma_{j}$, $\gamma_{j}^{\rm loss}$, and $\gamma_{\phi},$ describe the incoherent dynamics of the qubit, as shown in Eq.
+The cavity also induces a collective decay or “superradiance” on the qubits, described by a generalized Lindblad term $\tilde{{\cal D}}[x,y]\rho=x\rho y^{\dagger}-(y^{\dagger}x\rho+\rho y^{\dagger}x)/2$ in Eq.
+(13) and cavity-mediated decay rates
+${\color[rgb]{0,0,0}\gamma^{jl}_{\rm cm}}={}$ $\frac{{\color[rgb]{0,0,0}g_{j}g_{l}}\kappa}{(\kappa/2)^{2}+\Delta^{2}}.$ (18)
+Finally, the effective dynamics of the qubits also include local decay and dephasing rates $\gamma_{j}$, $\gamma_{j}^{\rm loss}$, and $\gamma_{\phi}$, as described by the normal Lindblad terms in Eq.
 (13).
-We will see that, for an optimal performance of the synchronized $N$-photon emitter, the residual cavity-mediated interactions $J_{\rm cm}^{jl}$ and decay $\gamma_{\rm cm}^{j}$ must be much smaller than the cavity-mediated driving $\Omega_{\rm cm}^{j}.$
 II.3 Synchronization dynamics and parameter regime
-Our goal is to realize the synchronized excitation and emission of each qubit, so that they act as nearly independent SPSs.
-The first stage of operation involves exciting each two-level systems from the ground state $|g\rangle_{j}$ to the excited state $|e\rangle_{j}$ via a fast cavity-mediated $\pi$-pulse of duration $T\sim\pi/(2\Omega_{\rm cm}^{j})$.
+Our goal is to realize the synchronized excitation and emission of each qubit so that they act as nearly independent SPSs.
+The first stage of operation involves exciting each two-level system from the ground state $|g\rangle_{j}$ to the excited state $|e\rangle_{j}$ via a fast cavity-mediated $\pi$-pulse of duration $T\sim\pi/(2\Omega_{\rm cm}^{j})$.
 We then expect that all emitters will produce synchronized photons, one on each of the antennas, on a timescale $t\gg 1/\gamma_{j}$.
@@
 and $\delta=\delta_{\rm cm}$, as explained in the following.
-First, to achieve synchronization and indistinguishability of the photon emissions, the system parameters should be as homogeneous as possible, and specially the qubit frequencies $\omega_{q}\approx\omega_{q}^{j}$, couplings $g\approx g_{j}$, and antenna decays $\gamma\approx\gamma_{j}$.
-In the remainder of the paper we will thus assume that all system parameters are homogeneous, except in Sec.
+First, to achieve synchronization and indistinguishability of the photon emissions, the system parameters should be as homogeneous as possible, especially the qubit frequencies $\omega_{q}\approx\omega_{q}^{j}$, couplings $g\approx g_{j}$, and antenna decays $\gamma\approx\gamma_{j}$.
+In the remainder of the paper, we will thus assume that all system parameters are homogeneous, except in Sec.
 V, where we analyze inhomogeneities and disorder.
@@
 Figure 2: Synchronization of two deterministic SPSs via a cavity.
-(a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity with fast residual oscillations due to non-RWA terms [cf.
+(a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim{\color[rgb]{0,0,0}10^{2}}$ photons is quickly created in the cavity with fast oscillations due to the large detuning $\Delta$ and the non-RWA terms [cf.
 red solid and black dashed lines].
-This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf.
+This induces an effective cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf.
 blue dash-dotted line].
 After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$.
-(b) Long time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its own antenna [cf.
+(b) Long-time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its antenna [cf.
 green dashed line].
-Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf.
+Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim{\color[rgb]{0,0,0}10^{-2}}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf.
 yellow dotted line].
@@
 We observe great agreement and no heating effects despite the large cavity driving strength $\Omega_{0}=4\omega_{d}\gg\kappa,|\Delta|$.
-These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row C of Table 1 and discussed in Sec.
+These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row A of Table 1 and discussed in Sec.
 III.1.
-$\frac{\gamma}{2\pi}[{\rm MHz}]$ $\frac{\gamma_{\rm loss}}{2\pi}[{\rm kHz}]$ $\frac{\gamma_{\phi}}{2\pi}[{\rm kHz}]$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\phi}/\gamma$ $\gamma_{\rm loss}/\gamma$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
-A 0.4 0 0 $1.0\cdot 10^{-2}$ 0 0 $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.997 0.994 $4.4\cdot 10^{-3}$ $2.0\cdot 10^{-5}$
-B 0.4 1 1 $1.0\cdot 10^{-2}$ $2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.994 0.989 $4.4\cdot 10^{-3}$ $2.5\cdot 10^{-3}$
-C 1.0 5 10 $2.6\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $5.0\cdot 10^{-3}$ $3.8\cdot 10^{-4}$ $4.7\cdot 10^{-5}$ 0.989 0.979 $1.0\cdot 10^{-2}$ $9.7\cdot 10^{-3}$
-D 3.0 50 100 $7.7\cdot 10^{-2}$ $3.3\cdot 10^{-2}$ $1.7\cdot 10^{-2}$ $1.3\cdot 10^{-4}$ $1.6\cdot 10^{-5}$ 0.968 0.936 $3.0\cdot 10^{-2}$ $3.1\cdot 10^{-2}$
+$g/2\pi$ $\kappa/2\pi$ $\Delta/2\pi$ $|\alpha_{\rm ss}|^{2}$ $\gamma/2\pi$ $\gamma_{\rm loss}/2\pi$ $\gamma_{\phi}/2\pi$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
+$\rm MHz$}] $\rm MHz$}] $\rm MHz$}] $\rm kHz$}] $\rm kHz$}] $\rm kHz$}]
+A 4.0 50 1400 $10^{2}$ 966 5 10 $2.4\cdot 10^{-2}$ $4.2\cdot 10^{-4}$ $1.2\cdot 10^{-2}$ 0.989 0.979 $1.4\cdot 10^{-2}$ $1.4\cdot 10^{-2}$
+B 4.0 50 1400 $10^{2}$ 3000 10 20 $7.5\cdot 10^{-2}$ $1.4\cdot 10^{-4}$ $3.8\cdot 10^{-3}$ 0.982 0.964 $2.8\cdot 10^{-2}$ $7.0\cdot 10^{-3}$
+C 0.2 400 50 $4\cdot 10^{4}$ 400 1 1 $1.0\cdot 10^{-2}$ $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.994 0.989 $4.4\cdot 10^{-3}$ $2.5\cdot 10^{-3}$
+D 0.2 400 50 $4\cdot 10^{4}$ 1000 1 1 $2.6\cdot 10^{-2}$ $3.8\cdot 10^{-4}$ $4.7\cdot 10^{-5}$ 0.991 0.983 $1.3\cdot 10^{-2}$ $7.7\cdot 10^{-3}$
 Table 1: Parameters and figures of merit for a circuit-QED implementation of two synchronized SPSs.
-We study four parameter sets: A, B, C, and D, which show increasing antenna coupling $\gamma$ and decoherence rates $\gamma_{\rm loss}$, $\gamma_{\phi}$.
-The rest of the parameters are common to all sets and read $g/2\pi=0.2{\rm MHz}$, $\kappa/2\pi=400{\rm MHz}$, $\Delta/2\pi=50{\rm MHz}$, and $\Omega_{0}/2\pi=40{\rm GHz}$, which leads to $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$ photons in the cavity on average.
-The detuning of the qubits is chosen to compensate for the cavity-induced Lamb-shift $\delta_{q}=\delta_{\rm cm}$.
-Depending on the decoherence rates of sets A-D, the inequalities in Eq.
-(18) are satisfied differently, as shown in columns 5th to 9th, and this influences the performance of the four figures of merit: single-photon generation probability ${\cal P}_{1}$ [cf.
+We consider four parameter sets: A and B correspond state-of-the-art parameters $g,\kappa,\Delta$ and $|\alpha_{\rm ss}|^{2}$ for different values of antenna couplings $\gamma$ and decoherence rates $\gamma_{\rm loss}$, $\gamma_{\phi}$.
+C and D correspond to more challenging long-term parameters due to the lower coupling $g$ and larger $|\alpha_{\rm ss}|^{2}$.
+All parameters lead to the same effective cavity-mediated drive $\Omega_{\rm cm}/2\pi\approx 40$ MHz.
+The detuning of the qubits is chosen to compensate for any cavity-mediated shift, including the Lamb-shift $\delta_{q}\sim\delta_{\rm cm}$ or residual Bloch-Siegert shifts.
+Depending on the couplings and decoherence rates of sets A-D, the inequalities in Eq.
+(19) are satisfied differently, as shown in columns 8th to 10th, and this influences the performance of figures of merit in the last four columns: single-photon generation probability ${\cal P}_{1}$ [cf.
 Sec.
@@
 Fig.
-2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ [cf.
-red solid line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf.
+2(a) displays the short time dynamics: the creation of a large coherent state in the cavity which induces an oscillatory occupation around the large steady-state value, $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim{\color[rgb]{0,0,0}10^{2}}$ [cf.
+red solid line and dashed black line], while mediating the simultaneous excitation of the qubits up to $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ (when the cavity driving is switched off) [cf.
 blue dash-dotted line].
 Fig.
-2(b) shows the long time dynamics, in which the photons are generated.
+2(b) shows the long-time dynamics, in which the photons are generated.
 Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf.
@@
 yellow dotted line].
-This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq.
-(2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
+This demonstrates that both emitted photons are highly independent of each other $S_{2}\approx 1$ in Eq.
+(3)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim{\color[rgb]{0,0,0}10^{-2}}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
 We performed the numerical simulations in Figs.
@@
 (13)-(18).
-They show an excellent agreement with deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf.
+They show an excellent agreement with small deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf.
 Fig.
-2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim 10^{4}\gg 1$.
+2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim{\color[rgb]{0,0,0}10^{2}}\gg 1$.
 This demonstrates that the strong cavity driving $\Omega_{0}\gg\kappa,|\Delta|$ does not introduce heating in the system, but only creates a large coherent state in the cavity with some fast oscillations around the mean value $|\alpha_{\rm ss}|^{2}\gg 1$ [cf.
@@
 2(a)].
-Since this fast oscillations have no detrimental impact in the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (12) which uses the mean steady value (9) for describing the effect of the cavity field (see App.
+Since these fast oscillations have no detrimental impact on the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (13) which uses the mean steady value (10) for describing the effect of the cavity field (see App.
 A for more details on the approximations involved in the model).
@@
 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits.
-For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz.
+For the cavity mode $a$, we consider a superconducting transmission line resonator with decay rate $\kappa/2\pi={\color[rgb]{0,0,0}50}{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz.
 These superconducting resonators support strong drives and can hold a large number of photons in steady state $|\alpha_{\rm ss}|^{2}\gg 1$ Pietikäinen et al.
 (2017); Jenkins et al.
-(2014); Macrì et al.
+(2014); Macri et al.
 (2020), as it is required in our scheme.
-For this, we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al.
-(2017); Jenkins et al.
-(2014).
+In particular, we consider $|\alpha_{\rm ss}|^{2}\approx 100$ of average photons in steady state, induced by a cavity driving of $\Omega_{0}/2\pi=14{\rm GHz}$ and a detuning of $\Delta/2\pi=1400{\rm MHz}$.
+For standard transmission line resonators such as in Ref.
+Pietikäinen et al.
+(2017), this order of driving strength $\Omega_{0}$ is obtained with a voltage source of power $P_{\rm in}\sim-50$ dBm.
 As two-level emitters, we consider flux-tunable transmon qubits DiCarlo et al.
@@
 (2014); Wang et al.
-(2020) weakly coupled to the resonator mode with strength $g/2\pi=0.2{\rm MHz}$.
-The cavity mediated driving amplitude is $\Omega_{\rm cm}/2\pi=|\alpha_{\rm ss}|g/2\pi\approx 39$MHz.
+(2020) weakly coupled to the resonator mode with strength $g/2\pi={\color[rgb]{0,0,0}4}{\rm MHz}$.
+The cavity mediated driving amplitude then reads $\Omega_{\rm cm}/2\pi=|\alpha_{\rm ss}|g/2\pi\approx 40$MHz.
 This is strong but still smaller than typical transmon anharmonicities $U/2\pi\sim 300-400{\rm MHz}$ Pechal et al.
@@
 (2020), thus preventing leakage to higher excitation states.
-The frequency of the qubits—which will be on the few GHz range—can be fine-tuned by sending a DC current through the antenna that is coupled to each qubit [cf.
+The frequency of the qubits—which will be in the few GHz range—can be fine-tuned by sending a DC current through the antenna that is coupled to each qubit [cf.
 Fig.
 1(b)].
-This is required to bring the qubits in resonance with the external drive, and to compensate the cavity-induced Lamb-shifts $\delta_{q}=\delta_{\rm cm}$ [See Sec.
+This is required to bring the qubits in resonance with the external drive, and to compensate any cavity-induced shifts such as Lamb-shifts $\delta_{q}\sim\delta_{\rm cm}$ or residual Bloch-Siegert shifts Pietikäinen et al.
+(2017) [See Sec.
 V for small deviations of this condition].
@@
 1(b)].
-The coupling rate between the qubit and antenna $\gamma$ can be chosen depending on the estimated decoherence parameters $\gamma_{\rm loss}$ and $\gamma_{\phi}$, so that the photons are efficiently ($\gamma\gg\gamma_{\rm loss}$) and coherently ($\gamma\gg\gamma_{\phi}$) collected.
-In our analysis, we consider four different combinations of the parameter sets ($\gamma,\gamma_{\rm loss}$$\gamma_{\phi}$), shown in Table 1.
-Parameter set D corresponds to the typical relaxation $\gamma_{\rm loss}/2\pi\sim 50$kHz and dephasing rates $\gamma_{\phi}/2\pi\sim 100$kHz found in microwave quantum photonics experiments Pechal et al.
-(2014); Zhou et al.
-(2020); Wang et al.
-(2020).
-For this case we choose a relatively large decay rate of $\gamma/2\pi=3$MHz so that the emission time is faster than the decoherence timescales, but at the same time it is still slower than the effective driving $\Omega_{\rm cm}/2\pi\sim 39$MHz [cf.
-column D in Table 1].
-In addition, parameter set C corresponds to decoherence rates reported in quantum computing experiments Barends et al.
+The coupling rate $\gamma$ between the qubit and antenna can be chosen to optimally fulfill the inequalities in Eq.
+(19) and it also determines the time-scale of photon emission $t\sim 15/\gamma$.
+For the circuit QED parameters discussed above, we consider two values of $\gamma$: Parameter set A with $\gamma/2\pi\approx 0.97$ MHz optimized to maximize the emission probabilities ${\cal P}_{1}\approx 0.99$ and $P_{2}\approx 0.98$, as well as parameter set B with $\gamma/2\pi=3$ MHz optimized for faster emission rate $C_{2}\approx 1.2$ MHz and lower cavity-mediated effects, $J_{\rm cm}/\gamma\sim 10^{-3}$.
+The value of $\gamma$ also limits the maximum unwanted decoherence rates $\gamma_{\rm loss}$ and $\gamma_{\phi}$ that can be tolerated by the device to work properly.
+In particular, we consider $\gamma_{\rm loss},\gamma_{\phi}\sim 2\pi\cdot 10{\rm kHz}\ll\gamma$ as realized in state-of-art experiments Barends et al.
 (2013); Caldwell et al.
@@
 (2019); Andersen et al.
-(2020), which have reached relaxation and dephasing rates as low as $\gamma_{\rm loss}/2\pi\sim 5$kHz and $\gamma_{\phi}/2\pi\sim 10$kHz, respectively.
-For these parameters we can choose a slightly lower coupling to the antenna of $\gamma/2\pi\sim 1$MHz, and thus improve the ratio $\gamma/\Omega_{\rm cm}\sim 2.6\cdot 10^{-2}$ [cf.
-column C in Table 1].
-Finally, set B corresponds to near-future parameters and set A to the ideal case of no decoherence, which we use as a reference to quantify the intrinsic limitations of the synchronization method.
-In the remainder of the paper, we use these four parameter sets A-D to characterize the performance and scaling of the synchronized SPSs for this circuit-QED implementation.
+(2020).
+Table 1 summarizes all the parameter sets considered in this work and their main figures of merit as discussed below.
+Note that we also consider extra parameter sets C and D, which reduce cavity-mediated effects by two orders of magnitude compared to A and B, respectively.
+This is achieved by having a 20-fold lower cavity-emitter coupling rate $g/2\pi=0.2$ MHz compensated by a larger cavity mean photon number $|\alpha_{\rm ss}|^{2}\sim 4\cdot 10^{4}$ so that the effective drive on the qubits is preserved, $\Omega_{\rm cm}=g|\alpha_{\rm ss}|\sim 2\pi\cdot 40$MHz.
+Engineering such a low coupling with high homogeneity on all emitters is challenging with current circuit QED technology because the coupling becomes of the same order as parasitic capacitive couplings typically found in superconducting circuits Wang et al.
+(2020).
+Nevertheless, reducing the inhomogeneity of these parasitic couplings may be possible in the long term via highly symmetric ring resonators Huang et al.
+(2021); Hazra et al.
+(2021).
+Parameters C and D consider this possibility and below we discuss how the scalablity can reach up to $N\sim 1000$ SPSs in this case.
+The second important requirement of parameters C and D is the large mean photon displacement $|\alpha_{\rm ss}|^{2}\sim 4\cdot 10^{4}$, but this is well below the values $10^{5}-10^{9}$ photons reported with current technology Pietikäinen et al.
+(2017); Jenkins et al.
+(2014) and is achieved by having a lower detuning $\Delta/2\pi=50$ MHz, larger cavity drive $\Omega_{0}/2\pi=40$ GHz and a decay of $\kappa/2\pi=400$ MHz.
+We note the strong drive required by our scheme is not harmful to the superconducting device since most of its energy is reflected by the cavity and it mainly induces a coherent displacement with no thermal component (see App.
+A for more details).
+In addition, in Sec.
+VI.3, we show that using a long superconducting resonator $L\sim 66$cm Sundaresan et al.
+(2015); Chang et al.
+(2020) the multi-mode structure of the cavity allows to reduce the required external drive down to $\Omega_{0}/2\pi\sim 2.6$ GHz while having the same effective quantities considered in sets C and D.
+In the remainder of the paper, we use the four parameter sets A-D in Table 1 to characterize the performance and scaling of the synchronized SPSs for this circuit-QED implementation.
 Figure 3: Few-photon statistics ${\cal P}_{n}$ and two-photon dependence $D_{2}$ as a function of various system parameters such as (a) cavity occupation $|\alpha_{\rm ss}|^{2}$, (b) qubit-cavity coupling $g/\kappa$, (c) qubit-antenna decay $\gamma/\Omega_{\rm cm}$, (d) qubit-drive detuning $(\delta-\delta_{\rm cm})/\Omega_{\rm cm}$, (e) dephasing $\gamma_{\phi}/\gamma$, and (f) relaxation loss $\gamma_{\rm loss}/\gamma$.
 The probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) are plotted using the left vertical axis, whereas the two-photon dependence $D_{2}$ (red), corresponds the right vertical axis.
-Solid lines are computed using the effective master equation in Eq.
-(12) and the discrete symbols using the full master equation in Eq.
-(6), which show excellent agreement.
-Parameters are given by set C of Table 1, except for the quantity that is changed on each subplot.
+All quantities are computed with the effective master equation (13) using parameter set A of Table 1, except for the quantity that is changed on each subplot.
 III.2 Efficiency of single-photon generation
@@
 This also applies to each of the probabilities ${\cal P}_{n}^{j}$ in our synchronized setup, if we want it to operate as a collection of $N$ independent emitters.
-To verify this, we have computed the few-photon probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) for a setup with $N=2$ superconducting qubits, varying different parameters around the set C from Table 1.
-The results in Figs.
-3(a)-(f), have been computed using the full model (6) and the effective equation (12), and are shown in lines and markers, respectively.
-The agreement between both methods confirms that the adiabatic elimination of the cavity is an excellent approximation over a broad range of parameters.
+To verify this, in Figs.
+3(a)-(f) we have computed the few-photon probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) for a setup with $N=2$ superconducting qubits, varying different parameters around the set A of Table 1.
 In Fig.
 3(a) we plot the effect of cavity occupation $|\alpha_{\rm ss}|^{2}$ (controlled via the cavity drive $\Omega_{0}$) on the single-photon efficiency of the SPS.
-All errors reduce with the cavity population $|\alpha_{\rm ss}|^{2},$ because a larger cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$ increases the speed at which the qubit is excited and reduces the probability of emitting two consecutive photons.
+All errors reduce or at least stay nearly constant when increasing the cavity population $|\alpha_{\rm ss}|^{2}$, because a larger cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$ increases the speed at which the qubit is excited and reduces the probability of emitting two consecutive photons.
 However, the driving strength is limited by our need to have an off-resonant cavity and by the anharmonicity of the qubits [cf.
@@
 However, a large coupling strength enhances the cavity-mediated dipole-dipole interactions $J_{\rm cm}\sim g^{2}$ and the collective decay $\gamma_{\rm cm}\sim g^{2}$.
-Both result in the transfer of excitations between qubits, increasing the probability of no emission ${\cal P}_{0}$ and of two photon emission ${\cal P}_{2}.$ For the parameters in Table 1 we find an optimal operation point around $g/\kappa=4\cdot 10^{-4}$.
+Both result in the transfer of excitations between qubits, increasing the probability of no emission ${\cal P}_{0}$ and two-photon emission ${\cal P}_{2}.$ For the parameters A in Table 1 we find an optimal operation point around $g/\kappa={\color[rgb]{0,0,0}10^{-1}}$.
 In Fig.
 3(c) we analyze the effect of the coupling to the antenna coupling $\gamma.$ Once more, there is an optimal point that satisfies (19), balancing the imperfections due to unwanted re-excitations ($\Omega_{\rm cm}\gg\gamma$) and cavity-mediated effects ($\gamma\ll J_{\rm cm},\gamma_{\rm cm}$).
-The optimal lays around $\gamma/\Omega_{\rm cm}\sim 2\cdot 10^{-2}$ for the parameter set C in Table 1.
+The minima lay around $\gamma/\Omega_{\rm cm}\sim[10^{-2},10^{-1}]$ for the parameter set A in Table 1.
 Fig.
-3(d) clearly shows an optimal operation of the emitters when they are on resonance with the drive $\delta=\delta_{\rm cm}$.
-At this point, a $\pi$ pulse excites the qubits with high fidelity, and the Lamb-shift is compensated.
+3(d) shows an optimal operation of the emitters when they are on resonance with the drive $\delta\sim\delta_{\rm cm}$.
+At this point, a $\pi$ pulse excites the qubits with high fidelity, and the Lamb shift is compensated, which is the main qubit’s frequency shift in our setup.
 Finally, in Fig.
@@
 (19)].
-Nevertheless, when $\gamma_{\phi},\gamma_{\rm loss}\sim\gamma$ we see a rapid increase of events where no photon is detected ${\cal P}_{0}.$ This is due to emitters becoming effectively off-resonant from the drive when $\gamma_{\phi}\sim\gamma$, or due to a decrease in the the collection efficiency in the antenna, when $\gamma_{\rm loss}\sim\gamma.$
-Figure 4: Multi-photon contamination of the SPS measured via the second-order correlation function.
-(a) Schematic of the Hanbury Brown and Twiss (HBT) experiment using a beam-splitter (BS) and two-photon coincidence measurements.
-(b) HBT correlation function $G^{(2)}_{\rm HBT}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$.
-Parameters correspond to set C of Table 1.
-(c) Enlarged plot of $G^{(2)}_{\rm HBT}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$.
-(d,e) Normalized correlation function $g^{(2)}_{\rm HBT}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
-Other parameters correspond to set C (blue) and set B (red).
-Decoherence has a small detrimental effect on the observed multi-photon contamitation via $g^{(2)}_{\rm HBT}[0]$.
+Nevertheless, when $\gamma_{\phi},\gamma_{\rm loss}\sim\gamma$ we see a rapid increase of events where no photon is detected ${\cal P}_{0}.$ This is due to emitters becoming effectively off-resonant from the drive when $\gamma_{\phi}\sim\gamma$, or due to a decrease in the collection efficiency in the antenna, when $\gamma_{\rm loss}\sim\gamma.$
 III.3 Efficiency of synchronized N-photon generation
@@
 The two-photon generation probability satisfies $P_{2}\approx({\cal P}_{1})^{2}\lesssim 1,$ and deviates from unity because of errors in the single-photon efficiency ${\cal P}_{1}\lesssim 1$.
-To quantify more precisely the independence of the $N$ emitted photons, we define the $N$-photon dependence or demultiplexing error as,
+To quantify more precisely the independence of the $N$ emitted photons, we compute the $N$-photon dependence or demultiplexing error $D_{N}$ from the definition in Eq. (3) as,
 $D_{N}=\frac{P_{N}}{({\cal P}_{1})^{N}}-1,$ (22)
-which describes the deviation from the ideal limit of $N$ perfectly independent and synchronized SPSs. This quantity is also related to the demultiplexing inefficiency in Eq. (2) as $S_{N}=1+D_{N}$.
-We have computed $P_{N}$ and $D_{N}$ in the simplest case of two SPSs, using the same photon counting methods introduced in Appendix B.
+which describes the deviation from the ideal limit of $N$ perfectly independent and synchronized SPSs.
+We have computed $P_{N}$ and $D_{N}$ in the simplest case of two SPSs as an example, using the same photon counting methods introduced in Appendix B.
 The results of these simulations are shown as red curves of Figs.
@@
 red curves and right vertical axis].
-We observe that $D_{2}$ is nearly insensitive to changes in the cavity occupation $|\alpha_{\rm ss}|^{2}$, the dephasing $\gamma_{\phi}/\gamma$, and out-coupling efficiency $\gamma_{\rm loss}/\gamma$, with typical values on order $D_{2}\sim 10^{-7}$ for parameters around set C of Table 1.
-Nevertheless, the variables $g$, $\gamma$, and $\delta$ can drastically change $D_{2}$ in various orders of magnitude.
+We observe that $D_{2}$ is nearly insensitive to changes in the cavity occupation $|\alpha_{\rm ss}|^{2}$, the detuning $\delta$, the dephasing $\gamma_{\phi}/\gamma$, and out-coupling efficiency $\gamma_{\rm loss}/\gamma$, with typical values on order $D_{2}\sim{\color[rgb]{0,0,0}10^{-4}}$ for parameters around set A of Table 1.
+Nevertheless, the variables $g$ and $\gamma$ can drastically change $D_{2}$ in various orders of magnitude.
 In particular, for larger cavity coupling $g$ or smaller antenna decay $\gamma$, $D_{2}$ increases because the cavity-mediated correlations $\sim g^{2}$ become more important in the timescale of the photon emission $\sim 15/\gamma$ [cf.
@@
 (19)].
-Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference.
-(a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements.
-(b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 4 excitation pulses with repetition rate $R=\gamma/15$.
-Parameters correspond to set C of Table 1.
-(c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa=[1,5,10]\cdot 10^{-4}$.
-(d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
-Other parameters correspond to set C (blue) and set B (red).
-The indistinguishability dramatically reduces with dephasing, a behaviour in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al.
-(2003).
 Finally, we note that the exact pulse shape $f(t)$ is not relevant for the synchronization dynamics as long as it approximately induces a $\pi$-pulse on the qubits.
@@
 Therefore, in the remainder of the paper we safely consider the ideal square pulse, but we keep in mind that a realistic experimental pulse shape gives similar results when properly optimized.
+Figure 4: Multi-photon contamination of the SPS measured via the second-order correlation function.
+(a) Schematic of the Hanbury Brown and Twiss (HBT) experiment using a beam-splitter (BS) and two-photon coincidence measurements.
+(b) HBT correlation function $G^{(2)}_{\rm HBT}(\tau)$ as a function of delay $\tau$, in the case of a train of 2 excitation pulses with repetition rate $R=\gamma/15$.
+Parameters correspond to set A of Table 1.
+(c) Enlarged plot of $G^{(2)}_{\rm HBT}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa={\color[rgb]{0,0,0}[2,8,20]\cdot 10^{-2}}$.
+(d,e) Normalized correlation function $g^{(2)}_{\rm HBT}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
+Decoherence has a small detrimental effect on the observed multi-photon contamination via $g^{(2)}_{\rm HBT}[0]$.
 III.4 Single-photon purity
@@
 $G_{\rm HBT}^{(2)}(\tau)\!=\!$ $\int\!\mathrm{d}t\,\langle b^{j{\dagger}}_{\rm out}(t)b^{j{\dagger}}_{\rm out}(t+\tau)b^{j}_{\rm out}(t+\tau)b^{j}_{\rm out}(t)\rangle,$ (23)
-where $\tau$ is the time delay between the two photon detections, and $b^{j}_{\rm out}(t)$ is the annihilation operator for an outgoing photon on the $j$-th antenna Gardiner and Zoller (2004).
+where $\tau$ is the time delay between the two-photon detections, and $b^{j}_{\rm out}(t)$ is the annihilation operator for an outgoing photon on the $j$-th antenna Gardiner and Zoller (2004).
 Appendix C contains details on the input-output theory and the calculation of these correlations.
 For simplicity, our analysis assumes a homogeneous setup and thus we omit the $j$ index in $G_{\rm HBT}^{(2)}(\tau)$.
-In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for train of 4 excitation pulses with a repetition rate
+In Fig. 4(b) we show the behaviour of the correlation $G_{\rm HBT}^{(2)}(\tau)$ in our setup for parameters A and a train of 2 consecutive excitation pulses with a repetition rate
 $R=\gamma/15.$ (24)
-We observe clear peaks at integer multiples of the repetition time $\tau=n/R$, corresponding to the detection of two photons coming from different pulses.
+We observe a clear peak at the repetition time $\tau=1/R$, corresponding to the detection of two photons coming from different pulses.
 Relevant to characterize the few-photon statistics of the SPS is the small peak that appears near zero time delay, $\tau\sim 0.$ A small area in this peak signals a small probability of two- and multi-photon emission Kiraz et al.
@@
 $g_{\rm HBT}^{(2)}[0]=\frac{\int_{0}^{1/(2R)}d\tau G_{\rm HBT}^{(2)}(\tau)}{\left(\int_{0}^{1/(2R)}dt\langle b_{\rm out}^{j{\dagger}}(t)b_{\rm out}^{j}(t)\rangle\right)^{2}}.$ (25)
-Here, the numerator corresponds to the area of the zeroth peak at $\tau\sim 0$ and the normalization is the area of one of the high peaks at $\tau\sim n/R$, and thus $g_{\rm HBT}^{(2)}[0]\sim{\cal P}_{2}/({\cal P}_{1})^{2}$.
+Here, the numerator corresponds to the area of the zeroth peak at $\tau\sim 0$ and the normalization is the area of the high peak at $\tau=1/R$, and thus $g_{\rm HBT}^{(2)}[0]\sim{\cal P}_{2}/({\cal P}_{1})^{2}$.
 In Figs.
 4(d)-(e), we plot the normalized correlation $g_{\rm HBT}^{(2)}[0]$ as a function of $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma$, respectively.
-We observe that $g_{\rm HBT}^{(2)}[0]$ is very insensitive to both types of decoherence, which is also consistent to the behaviour of ${\cal P}_{2}$ in Fig.
+We observe that $g_{\rm HBT}^{(2)}[0]$ is very insensitive to both types of decoherence, which is also consistent with the behavior of ${\cal P}_{2}$ in Fig.
 3(e)-(f) for the same parameters.
-In addition, we see that parameter set B gives a two-photon contamination of $g_{\rm HBT}^{(2)}[0]\sim 2.5\cdot 10^{-3}$, slightly lower than parameter set C [cf.
-Fig.
-4(e)-(f) and Table 1].
-Finally, we note that the single-photon purity $p$ is sometimes defined as $p=1-g_{\rm HBT}^{(2)}[0]$, which for parameter sets B and C of our circuit-QED implementation is higher than $p=0.99$.
+For parameters A, we obtain a two-photon contamination of $g_{\rm HBT}^{(2)}[0]{\color[rgb]{0,0,0}\sim 1.4\cdot 10^{-2}}$.
+We also calculated this quantity for parameters B-D of Table 1, for which we obtained $g_{\rm HBT}^{(2)}[0]\sim 2.8\cdot 10^{-2}$, $g_{\rm HBT}^{(2)}[0]\sim 4.4\cdot 10^{-3}$, and $g_{\rm HBT}^{(2)}[0]\sim 1.3\cdot 10^{-2}$, respectively.
+Finally, we note that the single-photon purity $p$ is typically defined as $p=1-g_{\rm HBT}^{(2)}[0]$, which for parameter sets A, B, C, and D read $p=0.986$, $p=0.972$, $p=0.996$, and $p=0.987$, respectively.
+We see that slower sources have a larger single-photon purity (see A and C compared to B and D).
 III.5 Indistinguishability
-Another important aspect of a high-performance multiphoton demultiplexing scheme is the degree of indistiguishability of the photons emitted by two different sources Senellart et al.
+Figure 5: Indistinguishability of single-photons from two different SPSs quantified by Hong-Ou-Mandel (HOM) interference.
+(a) Schematic of the HOM setup using a beam splitter to interfere two nearly independent photons, followed by two-photon coincidence measurements.
+(b) HOM correlation function $G^{(2)}_{\rm HOM}(\tau)$ as a function of delay $\tau$, in the case of a train of 2 excitation pulses with repetition rate $R=\gamma/15$.
+Parameters correspond to set A of Table 1.
+(c) Enlarged plot of $G^{(2)}_{\rm HOM}(\tau)$ at the zeroth peak, $\tau\sim 0$, for three values of couplings $g/\kappa={\color[rgb]{0,0,0}[2,8,20]\cdot 10^{-2}}$.
+(d,e) Normalized correlation function $g^{(2)}_{\rm HOM}[0]$ at zero delay, as function of dephasing $\gamma_{\phi}/\gamma$ and out-coupling inefficiency $\gamma_{\rm loss}/\gamma$, respectively.
+The indistinguishability dramatically reduces with dephasing, a behavior in good agreement with the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx({\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)/(1+{\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)$ Bylander et al.
+(2003).
+Another important aspect of a high-performance multiphoton demultiplexing scheme is the degree of indistinguishability of the photons emitted by two different sources Senellart et al.
 (2017).
@@
 $G_{\rm HOM}^{(2)}(\tau)\!=\!\int dt\langle b_{\text{BS}}^{1{\dagger}}(t)b_{BS}^{2{\dagger}}(t+\tau)b_{BS}^{2}(t+\tau)b_{BS}^{1}(t)\rangle,$ (26)
-where $\tau$ is the delay time between the two photon detections.
+where $\tau$ is the delay time between the two-photon detections.
 In addition, $b_{BS}^{1}(t)=[b^{1}_{\rm out}(t)+b^{2}_{\rm out}(t)]/\sqrt{2}$, and $b_{BS}^{2}(t)=[b^{1}_{\rm out}(t)-b^{2}_{\rm out}(t)]/\sqrt{2}$, describe the output fields of photons after interfering at the BS [cf.
@@
 This is perfect HOM interference, and any deviation from it (assuming an ideal BS) can be used to quantify the distinguishability of the generated single-photons.
+A standard figure of merit of indistinguishability that accounts for losses and other imperfections is the normalized HOM correlation function $g_{\rm HOM}^{(2)}[0]$ Fischer et al. (2016)
+$g_{\rm HOM}^{(2)}[0]=\frac{\int_{0}^{1/(2R)}d\tau G_{\rm HOM}^{(2)}(\tau)}{\prod_{k=1}^{2}\left(\int_{0}^{1/(2R)}dt\langle b_{BS}^{k{\dagger}}(t)b_{BS}^{k}(t)\rangle\right)}.$ (27)
+In analogy to Eq.
+(25), the numerator of Eq.
+(27) corresponds to the area below the zeroth peak, and the normalization to the area of the high peak at the repetition time $\tau={\color[rgb]{0,0,0}1/R}$.
+The indistinguishability of two SPSs can be simply defined as $I=1-g^{(2)}_{\rm HOM}[0].$
+We have computed $G_{\rm HOM}^{(2)}(\tau)$ as a function of $\tau$, for a setup with two emitters, parameters A of Table 1, and a train of 2 excitation pulses [cf.
+Fig.
+5(b)].
+Here, we observe a clear peak at the repetition time $\color[rgb]{0,0,0}\tau=1/R$} due to the detection of two nearly independent photons coming from two different pulses.
+The correlations $G_{\rm HOM}^{(2)}(\tau)$ are strongly suppressed at the origin, $\tau\sim 0$, showing only a minor zeroth peak which is enlarged in Fig.
+5(c).
+The small area of this zeroth peak manifests the high indistinguishability of the single-photons.
+We also see that this behavior is nearly insensitive to the coupling $g/\kappa$, but may strongly depend on the decoherence parameters.
+Figures 5(d)-(e) display $g^{(2)}_{\rm HOM}[0]$ as a function of the decoherence rates $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma.$ Note that in our implementation with parameters A we reach $I=0.986$, whereas for parameters B-D we obtain $I=0.993$, $I=0.998$, and $I=0.992$, respectively (cf.
+Table 1).
+We also find that the distinguishability $g^{(2)}_{\rm HOM}[0]$ is quite insensitive to $\gamma_{\rm loss}/\gamma$, but it dramatically increases with dephasing, following the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx({\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)/(1+{\color[rgb]{0,0,0}2}\gamma_{\phi}/\gamma)$ Bylander et al.
+(2003).
+This contrasts with the behaviour of $D_{2}$ in Fig.
+3(e), illustrating that the dependence error $D_{2}$ of two SPSs is qualitatively different from the distinguishability $g^{(2)}_{\rm HOM}[0]$ measured by HOM interference.
+Consequently, both figures of merit have to be considered when designing high-performance multiplexed and synchronized SPSs.
+IV Scalability of synchronized multiphoton generation
 Figure 6: Scalability performance of the synchronized multiphoton generation scheme.
@@
 (b) NN-photon generation rate $C_{N}=RP_{N}$ for pulsed excitation with repetition rate $R=\gamma/15$.
-Despite parameter set D shows the lowest $P_{N}$ efficiency, its faster repetition rate $R\sim 1.2$MHz allows for appreciably higher $C_{N}$.
+Despite parameter set B shows the lowest efficiency $P_{N}$, the faster repetition rate $R\approx 1.26$MHz allows for appreciably higher $C_{N}$.
 (c) Demultiplexing error $D_{N}$ up to $N=5$, and quadratic fit $D_{N}=\epsilon N(N-1)$ (cf.
 solid lines and inset).
-The fitted scaling factor depends on the system parameters [cf.
-Eq.
-(28)] and read $\epsilon_{A}=6.7\times 10^{-7}$, $\epsilon_{B}=6.7\times 10^{-7},\epsilon_{C}=1.0\times 10^{-7}$, and $\epsilon_{D}=1.3\times 10^{-8}$ for parameter sets A, B, C, and D, respectively.
-The extrapolation of all these calculations are valid as long as $D_{N}\lesssim 1$.
-A standard figure of merit of indistinguishability that accounts for losses and other imperfections is the normalized HOM correlation function $g_{\rm HOM}^{(2)}[0]$ Fischer et al. (2016)
-$g_{\rm HOM}^{(2)}[0]=\frac{\int_{0}^{1/(2R)}d\tau G_{\rm HOM}^{(2)}(\tau)}{\prod_{k=1}^{2}\left(\int_{0}^{1/(2R)}dt\langle b_{BS}^{k{\dagger}}(t)b_{BS}^{k}(t)\rangle\right)}.$ (26)
-In analogy to Eq.
-(24), the numerator of Eq.
-(26) corresponds to the area below the zeroth peak and the normalization to the area of one of the high peaks at the repetition times $\tau=n/R$.
-The indistinguishability of two SPSs can be simply defined as $I=1-g^{(2)}_{\rm HOM}[0].$
-We have computed $G_{\rm HOM}^{(2)}(\tau)$ as a function of $\tau$, for a setup with two emitters and a train of 4 excitation pulses [cf.
-Fig.
-5(b)].
-Here, we observe clear peaks at the repetition times $\tau=n/R$ due to the detection of two nearly independent photons coming from two different pulses.
-The correlations $G_{\rm HOM}^{(2)}(\tau)$ are strongly suppressed at the origin, $\tau\sim 0$, showing only a minor zeroth peak which is enlarged in Fig.
-5(c).
-The small area of this zeroth peak manifests the high indistinguishability of the single-photons prepared in our setup with parameter set C.
-We also see that this behaviour is nearly insensitive to the coupling $g/\kappa$, but may strongly depend on the decoherence parameters.
-Figures 5(d)-(e) display $g^{(2)}_{\rm HOM}[0]$ as function of the decoherence rates $\gamma_{\phi}/\gamma$ and $\gamma_{\rm loss}/\gamma.$ Note that in our implementation with parameters B and C we reach values above $I=0.99$, as inferred from Fig.
-5(e) and Table 1.
-We also find that the distinguishability $g^{(2)}_{\rm HOM}[0]$ is quite insensitive to $\gamma_{\rm loss}/\gamma$, but it dramatically increases with dephasing, following the analytical prediction $g^{(2)}_{\rm HOM}[0]\approx(\gamma_{\phi}/\gamma)/(1+\gamma_{\phi}/\gamma)$ Bylander et al.
-(2003).
-This contrasts with the behaviour of $D_{2}$ in Fig.
-3(e), illustrating that the dependence error $D_{2}$ of two SPSs is qualitatively different from the distinguishability $g^{(2)}_{\rm HOM}[0]$ measured by HOM interference.
-Consequently, both figures of merit have to be considered when designing high-performance multiplexed and synchronized SPSs.
-IV Scalability of synchronized multiphoton generation
-We are now in position to discuss the most important results of the present work: the scalability performance of the multiphoton emitter.
-To do so, we analyze the $N$-photon generation probability $P_{N}$ and the demultiplexing error $D_{N}$ as a function of $N$ and the various system parameters.
-From these results we demonstrate the feasibility to synchronize $N\sim 100-1000$ emitters with high $N$-photon generation rates $C_{N}=RP_{N}$.
+The fitted scaling factor depends on the system parameters and reads $\epsilon_{A}=1.3\times 10^{-4}$, $\epsilon_{B}=1.2\times 10^{-5}$, $\epsilon_{C}=6.7\times 10^{-7}$, and $\epsilon_{D}=1.0\times 10^{-7}$ for parameter sets A, B, C, and D, respectively.
+The extrapolation of all these calculations is valid as long as $D_{N}\ll 1$.
+We are now in a position to discuss the most important results of the present work: the scalability performance of the multiphoton emitter.
+To do so, we analyze the $N$-photon generation probability $P_{N}$ and the demultiplexing error $D_{N}$ as a function of $N$, for the four parameter sets of Table 1.
 We calculate $P_{N}$ in Eq.
@@
 6(a) displays $P_{N}$ as a function of $N$ computed from an average over $M=6000$ trajectories and considering the four parameter sets of Table 1.
-In the four cases and up to $N=12$, we numerically confirm that the SPSs are nearly perfectly synchronized and independent, satisfying $P_{N}\approx({\cal P}_{1})^{N}$ within the the statistical error $\Delta P_{N}\sim 10^{-3}$ [cf.
+In all these cases and up to $N=12$, we numerically confirm that the SPSs are nearly perfectly synchronized and independent, satisfying $P_{N}\approx({\cal P}_{1})^{N}$ within the statistical error $\Delta P_{N}\sim 10^{-3}$ [cf.
 black lines in Fig.
@@
 The reduction of $N$-photon generation probability with $N$ is thus mainly limited by the imperfections in the single-photon efficiency ${\cal P}_{1}$ and not by the synchronizing and demultiplexing scheme.
-For the decoherence-free parameters A, we predict a $12$-photon efficiency as high as $P_{12}\approx 0.97$, and when including realistic decoherence sources as in parameter set C [cf.
+For the state-of-the-art parameters A and B, we predict a $10$-photon efficiency of $P_{10}\approx 0.90$ and $0.83$, respectively [cf.
 inset in Fig.
-6(a)], this reduces only to $P_{12}\approx 0.88$.
-For a pulsed excitation of the SPSs with a repetition rate $R=\gamma/15$ (as in Secs.
-III.4-III.5), these high efficiencies imply $12$-photon generation rates $C_{12}=RP_{12}$ of $\sim 0.1-1$MHz [cf.
+6(a)].
+The reduction of efficiency for B is mainly due to the larger $\gamma$, but this also allows it to have a much larger $10$-photon emission rate of $C_{10}=1.0$ MHz compared to $C_{10}=0.4$ MHz for A [cf.
 Fig.
-6(b) and inset].
-Since parameter set D has the largest antenna decay $\gamma/2\pi\sim 3$MHz, it shows the largest $N$-photon generation rates $C_{N}$ despite having the lowest efficiencies $P_{N}$ and the largest decoherence rates.
+6(b)].
+The long-term parameters C and D present similar $N$-photon emission rates, but much larger efficiencies of $P_{10}\approx 0.94$ and $0.91$, respectively.
 To quantify more precisely the deviations from the ideal scaling $P_{N}\approx({\cal P}_{1})^{N}$ in Figs.
@@
 (22).
-For $N=2$ we know this error is as small as $D_{2}\sim 10^{-7}$ [cf.
-Figs.
+For $N=2$ we know this error is as small as $D_{2}\sim{\color[rgb]{0,0,0}10^{-4}}$ [cf.
+Fig.
 3] and therefore the QT calculations with an uncertainty $\Delta P_{N}\sim 10^{-3}$ do not provide enough precision.
@@
 Appendix B.1], which does not suffer from any statistical uncertainty.
-In this alternative method we simulate the photon counters at each antenna by an additional two-level system.
+In this alternative method, we simulate the photon counters at each antenna by an additional two-level system.
 This increases the Hilbert space dimension to $2^{2N}$ and thus limits the numerical computations to a maximum of $N\sim 5$ emitters.
@@
 6(c) we show the results of $D_{N}$ as a function $N$ for the four parameter sets of Table 1, and up to $N=5$.
-We confirm that $D_{N}<10^{-5}$ is well below the QT precision, and most importantly, we observe that the data is very well approximated by the quadratic fit
-$D_{N}=\epsilon N(N-1),$ (27)
-where the scaling factor $\epsilon\ll 1$ depends on the system parameters [cf.
+Importantly, we observe that the data is very well approximated by the quadratic fit $D_{N}=\epsilon N(N-1)$ stated in Eq.
+(4), with a scaling factor $\epsilon\ll 1$ that depends on the system parameters.
+For parameter sets A, B, C, and D, we obtain $\epsilon_{A}=1.3\times 10^{-4}$, $\epsilon_{B}=1.2\times 10^{-5}$, $\epsilon_{C}=6.7\times 10^{-7}$, and $\epsilon_{D}=1.0\times 10^{-7}$, respectively [cf.
 inset of Fig.
 6(c)].
-To analyze how the cavity-induced correlations influence $\epsilon$ and thus the scalability of the multiphoton scheme, we vary the system parameters around parameter set B and extract $\epsilon$ from fits to numerical calculations of $D_{N}$ up to $N=5$.
-The results of this analysis are represented by the data points in Fig.
-7(a), which correspond to independent sweeps over each of the system parameters.
-Remarkably, the apparently complicated behaviour of the system is very well approximated by the analytical formula [cf.
-solid lines in Fig.
-7(a)],
-$\epsilon\approx(1.42)\frac{g^{4}}{\gamma^{2}}\frac{\kappa}{\left[(\kappa/2)^{2}+\Delta^{2}\right]^{3/2}},$ (28)
-which manifests the dominant dependence on $g$, $\gamma$, $\kappa$, and $\Delta$, and omits the marginal effects of the effective driving $\Omega_{\rm cm}$, and the decoherence rates $\gamma_{\rm loss}$ and $\gamma_{\phi}$.
-This confirms our intuition that the synchronization and demultiplexing performance of our scheme is limited exclusively by residual cavity-induced correlations, provided $J_{\rm cm},\gamma_{\rm cm}\ll\gamma$ [cf.
+Notice that $\epsilon$ is smaller when the conditions $\gamma_{\rm cm},J_{\rm cm}\ll\gamma$ are better fulfilled overall, confirming that the synchronization performance is limited mainly by residual cavity-induced correlations [cf.
 Eq.
 (19)].
-Although Eq.
-(28) was extracted from a variation of parameters around set B of Table 1, this expression accurately predicts the scaling factors $\epsilon$ obtained from the fits on all parameters sets A-D in Fig.
-6(c) with an error smaller than $8\%$.
-Interestingly, we can exploit Eq.
-(28) to predict new parameter sets whose demultiplexing error grows quickly to $D_{N}\gtrsim 1$ for moderate system sizes $N\lesssim 10$.
-Setups with these characteristics can then be used to study the validity limits of the approximated $D_{N}$ scaling in Eq.
-(27).
+We will now analyze the validity limits of the quadratic scaling $D_{N}=\epsilon N(N-1)$.
+This is hard to do for the optimized parameter A-D because the deviation is expected for $N\gg 10$ due to the low correlation errors $\epsilon$.
+However, we can reduce $\gamma$ and/or increase $g$ to observe the growth of the demultiplexing error with $N$ up to $D_{N}\gtrsim 1$ with moderate system sizes $N\lesssim 10$.
 In Fig.
-7(b) we illustrate this analysis by calculating $D_{N}$ for two parameter sets similar to B, but: (i) with a $10$-fold coupling enhancement $g/2\pi=2$MHz (blue data), and (ii) with a $1/2$ reduction of decay $\gamma/2\pi=0.2$MHz in addition to the same $10$-fold increase in $g$ (red data).
-We perform $D_{N}=\epsilon N(N-1)$ fits to both data sets up to $N=5$, and we clearly observe the on-set of deviations from this scaling only when $D_{N}\gtrsim 1$ [cf.
+7, we illustrate this analysis by calculating $D_{N}$ for two modified parameter sets similar to A but with a 10-fold lower antenna decay $\gamma/2\pi=0.1$ MHz (blue diamonds), and with that lower decay and also larger coupling $\gamma/2\pi,g/2\pi]=[0.1,6]$MHz (red crosses).
+We also include the results for the standard parameter set A as a reference (blue circles).
+We perform quadratic fits $D_{N}=\epsilon N(N-1)$ to the three data sets up to $N=5$, obtaining increasing correlation errors: $\epsilon=1.4\times 10^{-4}$ (green), $\epsilon=8.5\times 10^{-3}$ (blue), and $\epsilon=3.5\times 10^{-2}$ (red).
+Remarkably, we indeed observe the on-set of deviations from the quadratic scaling only when $D_{N}\gtrsim 1$, which occurs for the red data at $N\sim 7$ [cf.
 solid red line in Fig.
 7].
-For $D_{N}\ll 1$, both data sets show excellent agreement with the quadratic scaling (27), and this constitutes the range of validity of this approximation.
-Figure 7: Limits on the scalability of the synchronized multiphoton generation.
-(a) Scaling factor $\epsilon$ as a function of various system parameters.
-The dimensionless variable $x=1$ corresponds to the parameters of set B in Table 1, which are denoted by $\tilde{g}$, $\tilde{\gamma}$, $\tilde{\Delta}$, $\tilde{\kappa}$, $\tilde{\Omega}_{\rm cm}$, $\tilde{\gamma}_{\rm loss}$, $\tilde{\gamma}_{\phi}$, and in this plot we change each of them from $25\%$ to $400\%$.
-We use these results to numerically extract Eq.
-(28).
-(b) Demultiplexing error $D_{N}$ calculated up to $N=10$ with QT averaged over $M=10^{4}$ trajectories.
-Both parameter sets are similar to B, except for $g/2\pi=2$MHz (blue data) and $g/2\pi,\gamma/2\pi]=[2,0.2]$MHz (red data).
+For $D_{N}\ll 1$, all these data sets show excellent agreement with the quadratic scaling $D_{N}=\epsilon N(N-1)$, and this constitutes the range of validity of this approximation.
+Figure 7: Demultiplexing error $D_{N}$ calculated up to $N=9$ with QT averaged over $M=10^{4}$ trajectories.
+Green circles correspond to the parameter set A of Table 1, whereas the other two data sets only differ in having $\gamma/2\pi=0.1$MHz (blue diamonds) and $\gamma/2\pi,g/2\pi]=[0.1,6]$MHz (red crosses).
 Solid lines correspond to fits of $D_{N}=\epsilon N(N-1)$ up to $N=5$, which result in the $\epsilon$ factors shown in the legend.
-For both data sets we observe that the quadratic scaling is a good approximation as long as $D_{N}\ll 1$, confirming the expected limit of validity.
+We observe that the quadratic scaling deviates only for $D_{N}\gtrsim 1$, confirming the validity range $D_{N}\ll 1$.
 […]
@@
 So far, we have analyzed the performance of the scheme assuming a perfectly homogeneous setup.
-In large scale implementations, however, the SPS parameters will unavoidably present some degree of inhomogeneity.
-In this last section we analyze the impact of these imperfections.
+In large-scale implementations, however, the SPS parameters will unavoidably present some degree of inhomogeneity.
+In this section, we analyze the impact of these imperfections.
 We consider inhomogeneity deviations on all qubit parameters, which we generically denote by
@@
 $\Pi(\delta y_{j})=\frac{1}{\sqrt{2\pi(\Delta y)^{2}}}{\rm exp}\left(-\frac{1}{2}\left[\frac{\delta y_{j}}{\Delta y}\right]^{2}\right).$ (29)
-Here, $\Delta y$ denotes the standard deviation associated to the disorder on each qubit parameter $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$.
+Here, $\Delta y$ denotes the standard deviation associated with the disorder on each qubit parameter $y=\omega_{q},g,\gamma,\gamma_{\phi},\gamma_{\rm loss}$.
 To quantify the impact of each of the inhomogeneities $\delta y_{j}$ on the performance of the multiphoton emitter, we compute the average $N$-photon efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{y}$ as
@@
 (8).
-We perform calculations up to $N=4$ slightly inhomogeneous photon sources, for parameters B and C, and we do not observe any detrimental effect up to a disorder strength of $10\%$ of the average values [cf.
+We perform calculations up to $N=4$ slightly inhomogeneous photon sources, for parameters A and C, and we do not observe any detrimental effect up to a disorder strength of $10\%$ of the average values [cf.
 Fig.
@@
 Figure 8: Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{y}$, as a function of disorder in antenna decay $y=\Delta\gamma/\gamma$, qubit dephasing $y=\Delta\gamma_{\phi}/\gamma_{\phi}$, and waveguide loss $y=\Delta\gamma_{\rm loss}/\gamma_{\rm loss}$.
-(a)-(b) Up to a disorder strength of $10\%$, we do not observe any impact of these three disorder sources on the synchronization and demultiplexing dynamics, as calculated up to $N=4$ for parameter sets B and C, respectively.
+(a)-(b) Up to a disorder strength of $10\%$, we do not observe any impact of these three disorder sources on the synchronization and demultiplexing dynamics, as calculated up to $N=4$ for parameter sets A and C, respectively.
 All results are obtained from averaging $M=10^{3}$ realizations for each type of disorder.
-On the other hand, inhomogeneities in the qubit frequencies $\delta\omega^{j}_{q}$ and in the couplings $\delta g_{j}$ are much more harmful for the multiphoton synchronization performance and therefore they need to be controlled more precisely in an experimental implementation of the device.
+On the other hand, inhomogeneities in the qubit frequencies $\delta\omega^{j}_{q}$ and in the couplings $\delta g_{j}$ are much more harmful to the multiphoton synchronization performance and therefore they need to be controlled more precisely in an experimental implementation of the device.
 Inhomogeneous couplings $g_{j}$, in particular, induce different cavity-mediated driving strengths $\Omega^{j}_{\rm cm}=\Omega_{\rm cm}+|\alpha_{\rm ss}|\delta g_{j}$, and therefore different times to realize an exact $\pi$-pulse on each qubit $\sim\pi/(2\Omega_{\rm cm}^{j})$.
@@
 and therefore a low coupling disorder $\delta g_{j}\ll g$ is required for a high performance of the synchronized multiphoton emitter.
-To precisely quantity the deviation from the ideal photon independence condition, we define the average dependence or demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{y}$ over disorder $y$ as
+To precisely quantify the deviation from the ideal photon independence condition, we define the average dependence or demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{y}$ over disorder $y$ as
 $\langle\!\langle D_{N}\rangle\!\rangle_{y}=\frac{\langle\!\langle P_{N}\rangle\!\rangle_{y}}{\left(\langle\!\langle{\cal P}_{1}\rangle\!\rangle_{y}\right)^{N}}-1.$ (33)
@@
 In Figs.
-9(a)-(b) we compute $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=4$ for parameter set B, and disorder strengths $\Delta g/g=0\%$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green).
-To achieve reasonable statistical errors $\Delta D_{N}\sim 10^{-6}$, these calculations require up to $M=3\cdot 10^{5}$ realizations.
-We observe that the $N$-photon efficiency and dependence, $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$, are minimally modified as long as $\delta g_{j}/g\lesssim 1\%$.
-Fortunately, current superconducting circuit technology allows for less than $1\%$ of geometrical errors in device fabrication, which is thus enough to suppress the disorder effects in the design of capacitive qubit-cavity couplings $g_{j}$.
+9(a)-(b) we compute $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=4$ for parameter set A, and disorder strengths $\Delta g/g=0\%$ (blue), $\Delta g/g={\color[rgb]{0,0,0}1}\%$ (red), and $\Delta g/g={\color[rgb]{0,0,0}2.5}\%$ (green).
+To achieve reasonable statistical errors $\Delta D_{N}\lesssim 10^{-{\color[rgb]{0,0,0}5}}$, these calculations require $\color[rgb]{0,0,0}M\sim 1.7\times 10^{5}$} realizations.
+We observe that the $N$-photon efficiency and dependence, $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{g}$, are minimally modified for $\delta g_{j}/g\lesssim{\color[rgb]{0,0,0}2.5\%}$.
+For the average coupling of $g/2\pi=4$ MHz, considered in parameter sets A and B, suppressing these disorder effects thus requires coupling inhomogeneities in the range $\delta g_{j}/2\pi\sim 0.1$ MHz.
+This level of inhomogeneity has already been reported in superconducting circuits Wang et al.
+(2020) and may be further reduced by using symmetric ring cavity configurations Huang et al.
+(2021); Hazra et al.
+(2021).
+For parameter sets C and D, however, reaching the same level of precision requires a coupling inhomogeneity down to $\delta g_{j}/2\pi\sim 5$ kHz which has not been achieved with current superconducting technology, but may be possible in the future by reducing residual stray capacitances in the circuit and also considering symmetric cavity configurations.
 Figure 9: Impact of inhomogeneities in qubit-cavity couplings $\delta g_{j}$ and qubit frequencies $\delta\omega_{j}$ on the performance of the synchronized multiphoton emitter.
-(a) Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ up to $N=4$ for coupling disorder strengths $\Delta g/g=0$ (blue), $\Delta g/g=0.5\%$ (red), and $\Delta g/g=1\%$ (green).
+(a) Average NN-photon generation probabilities $\langle\!\langle P_{N}\rangle\!\rangle_{g}$ up to $N=4$ for coupling disorder strengths $\Delta g/g=0$ (blue), $\Delta g/g={\color[rgb]{0,0,0}1}\%$ (red), and $\Delta g/g={\color[rgb]{0,0,0}2.5}\%$ (green).
 (b) Average demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{g}$ up to $N=3$, for the same disorder values as in (a).
@@
 We consider disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green).
-All calculations are performed using the parameter set B of Table 1 and up to $M=3\cdot 10^{5}$ realizations of each type of disorder.
+All calculations are performed using the parameter set A of Table 1 and up to $M=1.7\times 10^{5}$ realizations of each type of disorder.
 Finally, we analyze the impact of inhomogeneous qubit frequencies $\omega_{q}^{j}=\omega_{q}+\delta\omega^{j}_{q}$, which lead to slightly different qubit detunings with respect to the cavity driving frequency $\omega_{d}$, i.e.
@@
 $P_{e}^{j}\approx 1-\frac{1}{4}\left(\frac{\delta\omega^{j}_{q}}{\Omega_{\rm cm}}\right)^{2}+{\cal O}{\color[rgb]{0,0,0}\left(\frac{\delta\omega^{j}_{q}}{\Omega_{\rm cm}}\right)^{4}.}$ (35)
-Therefore, frequency inhomogeneties should also be controlled $\delta\omega_{q}^{j}\ll\Omega_{\rm cm}$, in order to have a high quality synchronization and demultiplexing.
+Therefore, frequency inhomogeneities should also be controlled $\delta\omega_{q}^{j}\ll\Omega_{\rm cm}$, to have a high-quality synchronization and demultiplexing.
 We discussed in Sec.
@@
 In Figs.
-9(c)-(d), we calculate $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ up to $N=4$ and for frequency disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 0\%$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ (green).
+9(c)-(d), we calculate $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ up to $N=4$ and for frequency disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0\%$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green).
 These results confirm that under these realistic disorder conditions the $N$-photon generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ is minimally altered with respect to the homogeneous prediction [cf.
@@
 This occurs at the expense of reducing the generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ so it is not a good limit for our purpose.
-The synchronized multiphoton generation scheme is thus resilient to moderate disorder in all system paremeters and we expect high $N$-photon generation efficiencies in realistic circuit-QED implementations of the device.
-VI Conclusions and Outlook
-In this work, we propose a scalable design for generating efficiently a large number of synchronized, independent and indistinguishable photons propagating over independent channels.
+We conclude that the synchronized multiphoton generation scheme is resilient to moderate disorder in all system parameters and that the required level of homogeneity is achievable in state-of-the-art circuit-QED implementations with parameter sets A and B.
+VI Multi-mode nature of the cavity
+So far we have assumed a mono-mode model for the cavity that synchronizes all emitters.
+Nevertheless, when increasing the length of the cavity to accommodate more qubits, the free spectral range of the cavity reduces and thus more normal cavity modes can participate in the dynamics Sundaresan et al.
+(2015).
+In this section, we consider these multi-mode effects and show that they can be described with the same effective model in Eqs.
+(13)-(18), but with a redefinition of the cavity-mediated parameters $\Omega_{\rm cm}$, $\delta_{\rm cm}$, $J_{\rm cm}$, and $\gamma_{\rm cm}$ [cf.
+Sec.
+VI.1].
+We analyze multi-mode effects for various cavity lengths and configurations [cf.
+Secs.
+(VI.2)-(VI.4)] and show how to fine-tune the parameters to recover the same synchronization dynamics discussed previously with a mono-mode model.
+Moreover, we find that the participation of more cavity modes can even be beneficial for the multi-photon generation scheme as it substantially reduces the required driving strength on the cavity.
+VI.1 Effective model including the multi-mode contributions
+To account for the multi-mode nature of the cavity, we generalize the Hamiltonian and master equation in Eqs.
+(5)-(7), by considering that the qubits interact with many cavity modes $m$, each of them described by a frequency $\omega_{c}^{m}$, a decay $\kappa_{j}$ and a qubit-cavity coupling $g_{j,m}$ (see App.
+A.1 for more details).
+These quantities scale with mode number $m\in\mathbb{N}$ as Sundaresan et al.
+(2015),
+$\omega_{c}^{m}$ $=m\omega_{c}^{1},$ (36)
+$\kappa_{m}$ $=m\kappa_{1},$ (37)
+$g_{j,m}$ $=\sqrt{m}g_{j,1},$ (38)
+with $\omega_{c}^{1}$, $\kappa_{1}$, and $g_{j,1}$ the frequency, decay, and coupling corresponding to the fundamental mode $m=1$.
+Note that these fundamental parameters in turn depend on the geometry and length $L$ of the cavity, such as $\omega_{c}^{1}=\pi c/L$ for an open transmission line resonator, and thus the free spectral range scales as $\sim 1/L$.
+It is possible to find parameters for which all cavity modes are in the bad-cavity limit, i.e.
+they are off-resonant and weakly coupled to the qubits, $g_{j,m}\ll\kappa_{m},|\Delta_{m}|$, so that their dynamics can be adiabatically eliminated.
+Here, $\Delta_{m}=\omega_{d}-\omega_{c}^{m}$ is the detuning between cavity modes and the external drive $\omega_{d}$, which is taken to be nearly resonant to the emitters $\omega_{d}\approx\omega_{j}^{q}$.
+As shown in Appendices A.2-A.3, the effective dynamics of the qubits under the above conditions is indeed described by the same effective Master equation and Hamiltonian in Eqs.
+(13)-(14), but with modified cavity-mediated quantities given by
+$\Omega_{\rm cm}^{j}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{\Omega_{0}g_{j,m}}{\sqrt{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}},$ (39)
+$\gamma^{jl}_{\rm cm}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}},$ (40)
+$J^{jl}_{\rm cm}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}\Delta_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}},$ (41)
+and $\delta^{j}_{\rm cm}=J_{\rm cm}^{jj}$.
+These expressions are the multi-mode generalization of Eqs.
+(15)-(18).
+They take into account the contribution of all cavity modes, with the largest effect coming from the modes closer to the qubit resonance $\omega_{d}\sim\omega_{j}^{q}$.
+Therefore, the effect of the multi-mode nature of the cavity is to modify the effective parameters of the dynamics.
+However, it is always possible to fine-tune the parameters to obtain the same effective physics discussed previously in the manuscript.
+In the following, we show this on three different multi-mode cavity configurations, namely an open transmission line resonator of moderate length [cf.
+Fig.
+10(a)], a very long transmission line resonator [cf.
+Fig.
+10(b)], and a ring cavity [cf.
+Fig.
+10(c)].
+For simplicity, we perform the analysis for homogeneous emitters in the remainder of the section.
+Figure 10: Multi-mode effects in the cavity-mediated synchronization of multiple single-photon sources.
+We consider three different configurations that lead to the same effective dynamics discussed in this work: (a) an open transmission line resonator of moderate length L=9.17L=9.17mm (solid) and L=8.46L=8.46mm (dashed), (b) a long open transmission line resonator with length L=662L=662mm, and (c) a ring cavity with length L=74.6L=74.6mm.
+For each of these configurations (a)-(c), panels (d)-(f) show the mode frequency $\omega_{c}^{m}$, the coupling/decay ratio $g_{m}/\kappa_{m}$, and the mean steady-state occupation $|\alpha^{\rm ss}_{m}|^{2}$ as a function of the mode index $m$.
+The pink horizontal line in the frequency plots indicates the transition frequency $\omega_{q}/2\pi$ of the single-photon emitters on each case to see which cavity modes are closer to resonance.
+VI.2 Transmission line resonator of moderate size
+In this subsection, we show that multi-mode effects are weak in the synchronization setup for cavities of moderate size and when the qubits couple close but still off-resonant to the fundamental resonator mode $m=1$.
+For this, we consider a transmission line resonator of length L=9.17L=9.17mm [cf.
+Fig.
+10(a)], so that the fundamental frequency reads $\omega_{c}^{1}/2\pi=7.2$ GHz.
+All other parameters correspond to set A of Table 1, namely $\Delta_{1}/2\pi=1400\textrm{ MHz}$, $g_{1}/2\pi=4\textrm{ MHz}$, $\kappa_{1}/2\pi=50\textrm{ MHz}$, and $\Omega_{0}/2\pi=14$ GHz.
+Fig.
+10(d) displays the mode frequencies $\omega_{c}^{m}$, the coupling/decay ratio $g_{m}/\kappa_{m}$, and the average occupation $|\alpha^{\rm ss}_{m}|^{2}$ as a function of mode label $m$.
+The pink horizontal line indicates the resonance frequency of the emitters, $\omega_{q}/2\pi=8.6$ GHz, to compare it with the mode frequencies.
+We see that all modes satisfy the bad-cavity limit with $g_{m}/\kappa_{m}<10^{-1}$ and that the occupation is indeed concentrated in the fundamental mode $|\alpha^{\rm ss}_{1}|^{2}\sim 100$ as it is closest to resonance.
+All other modes have occupations more than two orders of magnitude lower and they only slightly modify the effective quantities.
+Furthermore, if we consider a slightly shorter cavity of length $L=8.46$ mm and fine-tune the system parameters as $\Omega_{0}/2\pi=10.4\textrm{ GHz}$, $\kappa_{1}/2\pi=50\textrm{ MHz}$, and $\Delta_{1}/2\pi=2000\textrm{ MHz}$, we recover the same effective quantities obtained from the mono-mode model in Sec.
+III.1 within a $0.1\%$ error, except for $J_{\rm cm}$ which is even reduced by $28\%$.
+The modifications of the mode structure and couplings are indicated by the dashed lines in Fig.
+10(d), where one can see that steady-state occupations are less concentrated in mode $m=1$.
+The participation of more modes with similar weights allows the system to reduce cavity-mediated couplings as well as the required cavity drive.
+VI.3 Long transmission line resonator
+In this subsection, we show that the multi-mode structure of the cavity can be exploited to substantially reduce the driving strength required for parameters C and D from $\Omega_{0}/2\pi\sim 40$ GHz down to $\Omega_{0}/2\pi\sim 2.64$ GHz, while still inducing the same effective synchronization dynamics on the emitters.
+The idea is to consider a long resonator so that the free spectral range reduces and more modes can contribute to the effective parameters.
+In particular, we consider a long open transmission line resonator of length L=662L=662mm [cf.
+Fig.
+10(b)], which can fabricated by bending the transmission line inside the chip Sundaresan et al.
+(2015); Chang et al.
+(2020).
+The length of this resonator is roughly $70$ times larger than in the previous subsection so the mode $m=50$ of frequency $\omega_{c}^{50}/2\pi=5$ GHz becomes the closest to resonance [cf.
+left panel of Fig.
+10(e)].
+Here, we consider qubits of frequency $\omega_{q}/2\pi\approx 4.98$ GHz, and all other parameters are chosen to induce the same effective dynamics as in the mono-mode model with parameters C, namely $g_{50}/2\pi=0.1\textrm{ MHz}$, $\kappa_{50}/2\pi=10.9\textrm{ MHz}$, $\Delta_{50}/2\pi=17.1\textrm{ MHz}$, and $\Omega_{0}/2\pi=2.64\textrm{ GHz}$.
+Note that the driving strength $\Omega_{0}$ is more than an order of magnitude lower than required in the modo-mode configuration [cf.
+Fig.
+10(a)] This is because of the nearly $10$ off-resonant modes that play a non-negligible role in contributing to the effective parameters [cf.
+right panel of Fig.
+10(e)], and in particular to the cavity-mediated driving $\Omega_{\rm cm}=\sum_{m}g_{m}|\alpha_{\rm ss}^{m}|^{2}\sim 2\pi\cdot 39$ MHz.
+Although we have discussed that having a strong resonator drive on order $\Omega_{0}/2\pi\sim 40$GHz, as required for parameter set C, is not harmful to the superconducting system (as most of this energy is reflected by the cavity), it is instructive to show that the multi-mode structure of the cavity does not introduce detrimental effects and that it may be even beneficial for reducing drive power if needed.
+VI.4 Ring resonator
+In Sec.
+V, we discussed that the multi-photon synchronizing scheme is resilient to disorder in the couplings $g_{j}$ as long as this is controlled with a relative precision on the order or below $\sim 2.5\%$.
+In this subsection, we consider a ring cavity configuration [cf.
+Fig.
+10(c)], which is a very appealing alternative to improve the fabrication of homogeneous couplings $g_{j,m}\approx g_{m}$ due to its intrinsic circular symmetry.
+Prototypes of ring cavities Huang et al.
+(2021); Hazra et al.
+(2021) already demonstrate the generation of a very uniform electromagnetic field amplitude, leading to a coupling inhomogeneity below $1\%$.
+Note that for the same length $L$ of an open cavity, the ring cavity has a fundamental frequency twice as large $\omega_{c}^{1}=2\pi c/L$, due to its periodic boundary conditions.
+In particular, we consider a ring cavity of length $L=74.6\textrm{ mm}$, so that mode $m=3$ with $\omega_{c}^{3}/2\pi=5.31$GHz is the closest to the qubit resonance of $\omega_{q}/2\pi=6$ GHz [cf.
+left panel of Fig.
+10(f)].
+Other parameters are taken to induce an effective dynamics identical to parameters A in the mono-mode model, namely $\kappa_{3}/2\pi=15.6\textrm{ MHz}$, $g_{3}/2\pi=2.32\textrm{ MHz}$, $\Delta_{3}/2\pi=690\textrm{ MHz}$, and $\Omega_{0}/2\pi=3.7\textrm{ GHz}$.
+For these parameters, we see that the two modes $m=3,4$ are appreciably populated [cf.
+right panel of Fig.
+10(f)].
+As for the open cavity, increasing the length $L$ of the ring cavity also allows accommodating more qubits and reducing the driving strength $\Omega_{0}$ on the cavity, but the most important added feature of this symmetric configuration is the potential to reduce inhomogeneities in the fabrication of the multi-photon synchronizing device.
+VII Conclusions and Outlook
+In this work, we propose a scalable design for generating efficiently a large number of synchronized, independent, and indistinguishable photons propagating over independent channels.
 The synchronization is provided by a strongly driven resonator in the bad cavity and weak-coupling limits.
-The resonator allows for a simultaneous and robust control of all emitters, which deposit photon into individual waveguides, at a high rate and with a high collection efficiency.
+The resonator allows for a simultaneous and robust control of all emitters, which deposit photons into individual waveguides, at a high rate and with a high collection efficiency.
 Although our scheme can be implemented in cavity QED and nanophotonic platforms Welte et al.
@@
 (2019); Ellis et al.
-(2018); Goban et al.
-(2015), we have discussed a particularly efficient superconducting circuits implementation, with microwave resonators, superconducting qubits and microwave transmission lines Pechal et al.
+(2018), we have discussed an efficient circuit QED implementation, Pechal et al.
 (2014); Zhou et al.
@@
 (2013); Caldwell et al.
-(2018); McKay et al.
-(2019); Andersen et al.
-(2020); Dassonneville et al.
-(2020a).
-The only intrinsic limitation for the scalability of the synchronized multiphoton device are the cavity-mediated interactions and collective decay, which can create correlations between SPSs and the emitted photons.
-Nevertheless, for state-of-the-art circuit QED parameters, we show that these correlations are strongly suppressed on the timescale of photon emission, and that they only induce a nearly negligible quadratic demultiplexing error $D_{N}\approx\epsilon N(N-1)$ with $\epsilon\sim 10^{-8}-10^{-6}$.
-Remarkably, this allows for the synchronization of up to hundreds of nearly independent single-photons, even in the presence of decoherence and disorder.
+(2018), where the cavity synchronization mechanism is especially useful to reduce overhead with the number of independent controls Tamate et al.
+(2022).
+The only intrinsic limitations for the scalability of the synchronized multiphoton device are the cavity-mediated interactions and collective decay, which can create correlations between SPSs and the emitted photons.
+Nevertheless, we show that these correlations are strongly suppressed on the timescale of photon emission and that they only induce a nearly negligible quadratic demultiplexing error $D_{N}\approx\epsilon N(N-1)$ with $\epsilon\sim{\color[rgb]{0,0,0}10^{-4}-10^{-5}}$ for state-of-the-art circuit-QED parameters.
+Remarkably, this allows for the synchronization of up to hundreds of nearly independent single-photons, even in the presence of decoherence, disorder, and multi-mode cavity effects.
 Given that, as we show, each SPS can achieve single-photon efficiency, purity, and indistinguishability above $99\%$, and the parallel operation in our device enables the efficient creation of large $N$-photon states.
-For instance, we predict a $30$-photon probability of $P_{30}\approx 0.72$ at a rate $C_{30}\approx 270$kHz and a $100$-photon probability of $P_{100}\approx 0.33$ at a rate $C_{100}\approx 130$kHz.
-This is seven orders of magnitude more efficient than the most sophisticated multiplexed SPSs up to $N=14$ Lenzini et al.
+For instance, we predict a $30$-photon probability of $P_{30}\sim 0.72$ at a rate $C_{30}{\color[rgb]{0,0,0}\sim 300}$kHz and a $100$-photon probability of $P_{100}\sim{\color[rgb]{0,0,0}0.16}$ at a rate $C_{100}\sim 200$kHz.
+This is seven orders of magnitude more efficient than the most sophisticated multiplexed SPSs up to $N=14$ single photons Lenzini et al.
 (2017); Wang et al.
@@
 All these promising figures of merit can be further improved when implementing a more refined model for the SPSs such as three-level emitters Fischer et al.
-(2016), but this is independent from the efficient synchronization scheme we propose.
+(2016), but this is independent of the efficient synchronization scheme we propose.
 Scalable and deterministic sources of multi-photon states will be a key resource for realizing large-scale quantum information processing such as quantum optical neural networks Steinbrecher et al.
 (2019) or fault-tolerant photonic quantum computation Takeda and Furusawa (2019).
-In the short-term, the implemenation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al.
+In the short-term, the implementation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al.
 (2019) or quantum metrology Motes et al.
@@
 Many-body methods such as Matrix-product-states Huang et al.
-(2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Muñoz et al.
+(2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Sánchez-Muñoz et al.
 (2018); Dassonneville et al.
-(2020b); Lescanne et al.
+(2020); Lescanne et al.
 (2020); Le Jeannic et al.
-(2021) could be use to characterize them in the lab.
+(2021) could be used to characterize them in the lab.
 Acknowledgements
@@
 Porras, for helpful discussions.
-Work in Madrid was supported by project PGC2018-094792-B-I00 (MCIU/AEI/FEDER, UE), CAM/FEDER project No.
-S2018/TCS-4342 (QUITEMAD-CM), and CSIC Quantum Technology Platform PT-001.
+Work in Madrid was supported by project PGC2018-094792-B-I00 (MCIU/AEI/FEDER, UE), Proyecto Sinergico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM), and CSIC Quantum Technology Platform PT-001.
 T.R.
 further acknowledges funding from the EU Horizon 2020 program under the Marie Skłodowska-Curie grant agreement No.
-798397.
+798397 and the Ramón y Cajal program RYC2021-032473-I, financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR.
 M.L.
-acknowledges support from the China Scholarship Council and National Natural Science Foundation of China under Grant No.
+acknowledges support from the China Scholarship Council and the National Natural Science Foundation of China under Grant No.
 11475021.
 Appendix A Derivation of effective model for cavity synchronization
-In this appendix we outline the derivation of the effective master equation (12) by performing a displacement transformation on the cavity mode [cf.
+In this appendix, we outline the derivation of the effective master equation (13) starting from general dynamics describing $N$ emitters coupled to a multi-mode cavity [cf.
 Sec.
-A.1], and then an adiabatic elimination of the cavity fluctuations [cf.
+A.1].
+In particular, we perform a displacement transformation on all the cavity modes [cf.
 Sec.
-A.2].
-A.1 Coherent displacement of driven cavity mode
+A.2], and then an adiabatic elimination of the cavity fluctuations [cf.
+Sec.
+A.3] as shown in the following.
+A.1 Open quantum dynamics of emitters coupled to a multi-mode cavity
+The master equation and Hamiltonian describing $N$ two-level emitters coupled to a driven multi-mode cavity read,
+$\dot{\rho}(t)={}$ $-i[H(t),\rho]+\sum_{m}\kappa_{m}{\cal{D}}[a_{m}]\rho+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho$
+$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\rho+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}},[\sigma_{j}^{+}\sigma_{j}^{-}]\rho,$ (42)
+$H(t)={}$ $\sum_{m=1}\omega_{c}^{m}a_{m}^{\dagger}a_{m}+2\Omega(t)\sum_{m=1}(a_{m}+a_{m}^{\dagger})\cos(\omega_{d}t)$
+$+{}$ $\frac{1}{2}\sum_{j=1}^{N}\omega_{j}^{q}\sigma_{j}^{z}+\sum_{j=1}^{N}\sum_{m=1}g_{j,m}(a_{m}+a_{m}^{\dagger})(\sigma_{j}^{-}+\sigma_{j}^{+}).$ (43)
+Here, the system parameters are the same as in the mono-mode case given in Eqs.
+(5)-(7), except that the cavity is generalized to have many modes $m$, each one described by ladder operators ($a_{m}$, $a_{m}^{\dagger}$), a frequency $\omega_{c}^{m}$, a decay $\kappa_{j}$ and a coupling to the qubits $g_{j,m}$.
+A.2 Coherent displacement of driven cavity modes
 The quantum Langevin equations Gardiner and Zoller (2004) associated to the master equation (42) read
-$\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}$
-$-2i\Omega(t)\cos(\omega_{d}t),$ (37)
-$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(a+a)^{{\dagger}}\sigma_{j}^{z}$
-$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (38)
-$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})-2ig_{j}(a^{{\dagger}}+a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
-$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (39)
-Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}(t)$ corresponds to the input noise field of the cavity mode, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$.
-A coherent driving on the cavity induces a coherent state component on this mode, and thus it is convenient to displace it using the transformation $a\rightarrow\alpha+\delta a$ as in Eq.
+$\dot{a}_{m}={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]a_{m}-i\sum_{j=1}^{N}g_{j,m}(\sigma_{j}^{-}+\sigma_{j}^{+})$ (44)
+$-2i\Omega_{m}(t)\cos(\omega_{d}t)-\sqrt{\kappa_{m}}a_{\rm in}^{m},$
+$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}$ (45)
+$+i\sum_{m}g_{j,m}(a_{m}+a_{m}^{{\dagger}})\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}$
+$+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\chi(t)\sigma_{j}^{-},$ (46)
+$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma^{j}_{\rm loss})(1+\sigma_{j}^{z})$ (47)
+$-2i\sum_{m}g_{j,m}(a_{m}^{{\dagger}}+a_{m})(\sigma_{j}^{+}-\sigma_{j}^{-})$
+$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$
+Here, $\chi(t)$ is a stochastic white noise dephasing fluctuation satisfying $\langle\!\langle\chi(t)\chi(t^{\prime})\rangle\!\rangle=\delta(t-t^{\prime})$ Ramos and García-Ripoll (2018), $a_{\rm in}^{m}(t)$ corresponds to the input noise field of the cavity mode $m$, $b_{\rm in}^{j}(t)$ the input field of photons in antenna $j$, and $c_{\rm in}^{j}(t)$ the photonic input field of unwanted channels coupled to qubit $j$.
+A coherent driving on the cavity induces a coherent state component on this mode, and thus it is convenient to displace it using the transformation $a_{m}\rightarrow\alpha_{m}+\delta a_{m}$ as in Eq.
 (8).
-Here, $\delta a$ corresponds to the quantum fluctuation of the cavity mode around its classical value $\alpha$.
-By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation reads
-$\delta\dot{a}={}$ $-\left[\kappa/2+i\omega_{c}\right]\delta a-i\sum_{j=1}^{N}g_{j}(\sigma_{j}^{-}+\sigma_{j}^{+})-\sqrt{\kappa}a_{\rm in}(t),$ (40)
+Here, $\delta a_{m}$ corresponds to the quantum fluctuation of the cavity mode around its classical value $\alpha_{m}$.
+By separating the classical and quantum cavity components, the displaced Langevin equation for the fluctuation of cavity mode $m$ reads,
+$\delta\dot{a}_{m}={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]\delta a_{m}-i\sum_{j=1}^{N}g_{j,m}(\sigma_{j}^{-}+\sigma_{j}^{+})$
+$-\sqrt{\kappa_{m}}a_{\rm in}^{m}(t),$ (48)
 and for the qubits,
-$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+ig_{j}(\delta a+\delta a^{{\dagger}})\sigma_{j}^{z}$
-$+2ig_{j}{\rm Re}\{\alpha\}\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}$
-$-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (41)
-$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-2ig_{j}(\delta a^{{\dagger}}+\delta a)(\sigma_{j}^{+}-\sigma_{j}^{-})$
-$-4ig_{j}{\rm Re}\{\alpha\}(\sigma_{j}^{+}-\sigma_{j}^{-})-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
+$\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss})/2+i\omega_{j}^{q}\right]\sigma_{j}^{-}+2i\sum_{m}g_{j,m}{\rm Re}\{\alpha_{m}\}\sigma_{j}^{z}$
+$+i\sum_{m}g_{j,m}(\delta a_{m}+\delta a_{m}^{{\dagger}})\sigma_{j}^{z}+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}$
+$+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$ (49)
+$\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j})(1+\sigma_{j}^{z})-4i\sum_{m}g_{j,m}{\rm Re}\{\alpha_{m}\}(\sigma_{j}^{+}-\sigma_{j}^{-})$
+$-2i\sum_{m}g_{j,m}(\delta a_{m}^{{\dagger}}+\delta a_{m})(\sigma_{j}^{+}-\sigma_{j}^{-})$
+$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}(t)+{\rm h.c.}\right]$
 $-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}(t)+{\rm h.c.}\right].$ (50)
-Here, $\alpha(t)$ is determined via the classical differential equation (8).
-The master equation associated to the above Langevin equations (40)-(42) with the displaced cavity reads
-$\dot{\rho}^{\prime}(t)={}$ $-i[H^{\prime}(t),\rho^{\prime}]+\kappa{\cal{D}}[\delta a]\rho^{\prime}+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}$
+Here, $\alpha_{m}(t)$ is determined via the classical differential equation,
+$\dot{\alpha}_{m}(t)={}$ $-\left[\kappa_{m}/2+i\omega_{c}^{m}\right]\alpha_{m}(t)-2i\Omega_{0}f(t)\cos(\omega_{d}t),$ (51)
+which is a multi-mode generalization of Eq. (10). The master equation associated to the above Langevin equations (48)-(50) with the displaced cavity reads
+$\dot{\rho}^{\prime}(t)={}$ $-i[H^{\prime}(t),\rho^{\prime}]+\sum_{m}\kappa_{m}{\cal{D}}[\delta a_{m}]\rho^{\prime}+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}$
 $+\sum_{j=1}^{N}\gamma^{j}_{\rm loss}{\cal{D}}[\sigma_{j}^{-}]\rho^{\prime}+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}}[\sigma_{j}^{+}\sigma_{j}^{-}]\rho^{\prime},$ (52)
 where the system Hamiltonian after the displacement transformation $H^{\prime}(t)$ is given by,
-$H^{\prime}(t)={}$ $\omega_{c}\delta a^{{\dagger}}\delta a+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
-$+\sum_{j=1}^{N}2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (44)
-Notice that the master equation (43) for the dynamics of $\delta a$ is the same as the master equation (6) for the total field $a$, except for the driving term in displaced Hamiltonian (44), which acts on the qubits rather than on the cavity as indicated in Eq.
+$H^{\prime}(t)={}$ $\sum_{m}\omega_{c}^{m}\delta a_{m}^{{\dagger}}\delta a_{m}+\sum_{j=1}^{N}\frac{\omega_{j}^{q}}{2}\sigma_{j}^{z}$
+$+\sum_{j=1}^{N}\sum_{m}g_{j,m}(\delta a_{m}+\delta a_{m}^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
+$+\sum_{j=1}^{N}\sum_{m}2g_{j,m}{\rm Re}\{\alpha_{m}(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (53)
+Notice that, the displaced master equation (52) has the same shape as Eq.
+(7) except for the inclusion of multi-mode nature and is expressed in terms of the fluctuations $\delta a_{m}$.
+The driving term in displaced Hamiltonian (53) acts on the qubits rather than on the cavity modes and this is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
+Equation (52) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega_{0}|\gg\kappa_{m},|\omega_{c}^{m}-\omega_{d}|$ because most of the cavity photons can be taken into account by the classical field $\alpha_{m}(t)$, and the fluctuations are only weakly populated, $\langle\delta a_{m}^{\dagger}\delta a_{m}\rangle\ll 1$.
+The steady state solution of Eq.
+(51) with $f(t)=1$ is given by $\alpha_{\rm ss,m}=-i\Omega_{0}/(\kappa_{m}/2-i\Delta_{m})$, with $\Delta_{m}=\omega_{d}-\omega_{c}^{m}$ the detuning of the $m$-th mode with respect to the cavity drive.
+This implies that only a few cavity modes with $\omega_{c}^{m}\sim\omega_{d}$ have a significant displacement $\alpha_{m}$ during evolution.
+In addition, we consider that all cavity modes are weakly coupled to the emitters, which are nearly resonant to the cavity drive, i.e.
+$g_{j,m}\ll\omega_{q}^{j}\approx\omega_{d}$, so the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq.
 (53).
-This is the effective cavity-mediated driving on the qubits that allows for the synchronization of many SPSs in our scheme.
-The displaced master equation (43) is very efficient to perform numerical simulations of the dynamics in the case of strong cavity driving $|\Omega(t)|\gg\kappa,|\Delta|$ because most of the cavity photons can be taken into account by the classical field $\alpha(t)$, and the fluctuation $\langle\delta a^{\dagger}\delta a\rangle$ is only weakly populated.
-This largely reduces the Hilbert space dimension needed to numerically simulate the dynamics of the driven cavity mode coupled to all qubits.
-In addition, if we consider the cavity in the weak coupling and bad cavity limit given by Eq.
-(11), the dynamics of the system can be further simplified by applying the rotating wave approximation (RWA) to the Hamiltonian in Eq.
-(44).
-On the one hand, since $g_{j}\ll\omega_{c}+\omega_{q}^{j}$ and the fluctuation is weakly populated, we can approximate the qubit-fluctuation coupling as
-$g_{j}(\delta a+\delta a^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})\approx g_{j}(\delta a^{\dagger}\sigma_{-}+{\rm h.c.}).$ (45)
+On the one hand, we can approximate the qubit-fluctuation coupling as
+$\sum_{m}g_{j,m}$ $(\delta a_{m}+\delta a_{m}^{{\dagger}})(\sigma_{j}^{-}+\sigma_{j}^{+})$
+$\approx\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}(\delta a_{m}^{\dagger}\sigma_{j}^{-}+{\rm h.c.}).$ (54)
 On the other hand, the cavity-mediated driving on the qubits can also be simplified by applying RWA as
-$2g_{j}{\rm Re}\{\alpha(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})\approx g_{j}|\alpha_{\rm ss}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi)}+{\rm h.c.}).$ (46)
-Here, $|\alpha_{\rm ss}|=\Omega_{0}/\sqrt{(\kappa/2)^{2}+\Delta^{2}}$ and $\phi={\rm Arctan}(2\Delta/\kappa)-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile.
-To demonstrate this last approximation (46), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha(0)=0$.
-Then, the solution of the classical equation (8) for $\alpha(t)$ reads
-$\alpha(t)={}$ $|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}(1-e^{-(\kappa/2-i\Delta)t})$
-$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa/2+2i\omega_{d})t}),$ (47)
-where we have also used the inequalities $|\Delta|,\kappa\ll\omega_{d}$ valid for our parameter conditions.
-The first term in (47) reaches the steady state value $\sim|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ in a timescale $\sim 1/\kappa$ and the second term causes the fast but low amplitude oscillation observed in Fig.
+$\sum_{m}$ $g_{j,m}2{\rm Re}\{\alpha_{m}(t)\}(\sigma_{j}^{+}+\sigma_{j}^{-})$ (55)
+$\approx\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}|\alpha_{\rm ss,m}|f(t)(\sigma_{j}^{+}e^{-i(\omega_{d}t-\phi_{m})}+{\rm h.c.}),$
+where $|\alpha_{\rm ss,m}|=\Omega_{0}/\sqrt{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}$ and $\phi_{m}={\rm Arctan}(2\Delta_{m}/\kappa_{m})-\pi/2$ are the steady state amplitude and phase of the cavity displacement, respectively, and $f(t)$ is a step function profile.
+To demonstrate this last approximation (55), we first assume the square pulse is switched on $f(t)=1$ and the cavity is initially empty $\alpha_{m}(0)=0$.
+Then, the solution of the classical equation (9) for $\alpha_{m}(t)$ reads
+$\alpha_{m}(t)={}$ $|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}(1-e^{-(\kappa_{m}/2-i\Delta_{m})t})$
+$-\frac{\Omega_{0}}{2\omega_{d}}e^{i\omega_{d}t}(1-e^{-(\kappa_{m}/2+2i\omega_{d})t}),$ (56)
+where we have also used the inequalities $|\Delta_{m}|,\kappa_{m}\ll\omega_{d}$ valid for our parameter conditions.
+The first term in (56) reaches the steady state value $\sim|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}$ in a timescale $\sim 1/\kappa_{m}$ and the second term causes the fast but low amplitude oscillation observed in Fig.
 2(a).
-Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss}|\gg 1$ as it is the case for our parameters.
-Since in bad cavity limit (11) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa$, we can use the steady state value $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits.
+Even for a large driving strength $\Omega_{0}\gtrsim\omega_{d}$, the second term can be neglected with respect to the first when $|\alpha_{\rm ss,m}|\gg 1$ as it is the case for our parameters.
+Since in bad cavity limit (12) the timescale for the qubit dynamics is much longer than $\sim 1/\kappa_{m}$, we can use the steady state value $\alpha_{m}(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi_{m})}$ to approximate the cavity displacement when the drive is switched on $f(t)=1$ as it would be instantaneous for the qubits.
 Similarly, after the cavity reaches the steady state and it is switched off ($f(t)=0$), the dynamics in Eq.
-(8) predicts an exponential decay of $\alpha(t)$ with rate $\sim\kappa$.
-In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha(t)\approx 0$ for $f(t)=0$.
-In summary, under the above approximations we have $\alpha(t)\approx|\alpha_{\rm ss}|e^{-i(\omega_{d}t-\phi)}f(t)$ so that when replacing it in the last term of Eq.
-(44), and applying RWA provided $g_{j}|\alpha_{\rm ss}|\ll\omega_{d}$, we obtain Eq.
+(9) predicts an exponential decay of $\alpha_{m}(t)$ with rate $\sim\kappa_{m}$.
+In the bad cavity limit we can also approximate this emptying of the cavity as instantaneous for the qubit and thus $\alpha_{m}(t)\approx 0$ for $f(t)=0$.
+In summary, under the above approximations we have $\alpha_{m}(t)\approx|\alpha_{\rm ss,m}|e^{-i(\omega_{d}t-\phi_{m})}f(t)$ so that when replacing it in the last term of Eq.
+(53), and applying RWA provided $g_{j,m}|\alpha_{\rm ss,m}|\ll\omega_{c}^{m},\omega_{q}^{j}\approx\omega_{d}$, we obtain Eq.
 (55).
 Replacing Eqs.
-(45)-(46) in (44), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (43) with the RWA Hamiltonian:
-$H^{\prime}(t)\approx{}$ $-\Delta\delta a^{{\dagger}}\delta a-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}+\sum_{j=1}^{N}g_{j}(\delta a^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a)$
+(54)-(55) in (53), and going to a rotating frame with respect to the cavity drive frequency $\omega_{d}$, the dynamics of the system in the displaced picture is finally given by the master equation (52) with the RWA Hamiltonian,
+$H^{\prime}(t)\approx{}$ $-\sum_{\omega_{c}^{m}\sim\omega_{d}}\Delta_{m}\delta a_{m}^{{\dagger}}\delta a_{m}-\frac{1}{2}\sum_{j=1}^{N}\delta_{j}\sigma_{j}^{z}$
+$+\sum_{j=1}^{N}\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}(\delta a_{m}^{{\dagger}}\sigma_{j}^{-}+\sigma_{j}^{+}\delta a_{m})$
 $+\sum_{j=1}^{N}\Omega_{\rm cm}^{j}(\sigma_{j}^{+}+\sigma_{j}^{-}).$ (57)
-Here, we have defined $\Omega_{\rm cm}^{j}=g_{j}|\alpha_{\rm ss}|$ as in Eq.
-(14) and absorbed the constant phase $\phi$ in the definition of the qubit and fluctuation operators.
+Here, $\Omega_{\rm cm}^{j}=\sum_{\omega_{c}^{m}\sim\omega_{d}}g_{j,m}|\alpha_{\rm ss,m}|$ is the multi-mode generalization of the cavity-mediated driving in Eq.
+(15) and we have also absorbed the constant phase $\phi_{m}$ in the definition of the qubit and fluctuation operators.
 The qubit detunings read $\delta_{j}=\omega_{d}-\omega_{q}^{j}$ as in the main text.
-A.2 Adiabatic elimination of cavity fluctuations
-As discussed in the previous subsection, in the bad cavity limit $g_{j},|\delta_{j}|\ll\kappa$, the cavity reaches very quickly a coherent steady state, with its quantum fluctuations $\delta a$ close to the vacuum state $|0\rangle$.
-In this situation it is convenient to adiabatically eliminate the cavity fluctuation and obtain an effective dynamics for the degrees of freedom of the qubits only.
+A.3 Adiabatic elimination of cavity fluctuations
+As discussed in the previous subsection, in the bad cavity limit $g_{j,m},|\delta_{j}|\ll\kappa_{m}$, the cavity modes reach very quickly a coherent steady state, with its quantum fluctuations $\delta a_{m}$ close to the vacuum state $|0\rangle$.
+In this situation, it is convenient to adiabatically eliminate the cavity fluctuation and obtain effective dynamics for the degrees of freedom of the qubits only.
 To do so, we formally integrate Eq.
-(40) and apply Markov approximation provided the qubits’ dynamics evolves slowly on the time-scale $\sim 1/\kappa$.
+(48) and apply Markov approximation provided the qubits’ dynamics evolve slowly on the time-scale $\sim 1/\kappa_{m}$.
 As a result, we get
-$\delta a(t)\approx{}$ $-i\sum_{j}\frac{g_{j}}{(\kappa/2-i\Delta)}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa}}{(\kappa/2-i\Delta)}a_{\rm in}(t)$
-${\color[rgb]{0,0,0}-i\sum_{j}\frac{g_{j}}{\kappa/2+i(2\omega_{d}-\Delta)}\sigma_{j}^{+}(t).}$ (49)
-Replacing the above expressions in the original quantum Langevin equations (38)-(39), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations
+$\delta a_{m}(t)\approx{}$ $-i\sum_{j}\frac{g_{j,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{-}(t)-\frac{\sqrt{\kappa_{m}}a_{\rm in}^{m}(t)}{(\kappa_{m}/2-i\Delta_{m})}$
+$-i\sum_{j}\frac{g_{j,m}}{\kappa_{m}/2+i(2\omega_{d}-\Delta_{m})}\sigma_{j}^{+}(t).$ (58)
+Replacing the above expressions in the original quantum Langevin equations (45)-(47), applying RWA, and going to a rotating frame with frequency $\omega_{d}$, we obtain the effective dynamics of the qubits, given by the Langevin equations,
 $\dot{\sigma}_{j}^{-}={}$ $-\left[(\gamma_{j}+\gamma^{j}_{\rm loss}+\gamma_{\rm cm}^{j})/2-i(\delta_{j}-\delta^{j}_{\rm cm})\right]\sigma_{j}^{-}$ (59)
-$+i\Omega_{\rm cm}^{j}f(t)\sigma_{j}^{z}+\sum_{l\neq j}\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{z}\sigma_{j}^{-}+\sqrt{\gamma^{j}_{\rm cm}}\sigma_{j}^{z}a_{\rm in}$
-$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-},$
+$+i\Omega_{\rm cm}^{j}f(t)\sigma_{j}^{z}+\sum_{l\neq j}\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{z}\sigma_{l}^{-}$
+$+\sqrt{\gamma_{j}}\sigma_{j}^{z}b_{\rm in}^{j}+\sqrt{\gamma^{j}_{\rm loss}}\sigma_{j}^{z}c_{\rm in}^{j}-i\sqrt{2\gamma_{\phi}^{j}}\Theta(t)\sigma_{j}^{-}$
+$+\sum_{\omega_{c}^{m}\sim\omega_{d}}\sqrt{\frac{(g_{j,m})^{2}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}}\sigma_{j}^{z}a_{\rm in}^{m},$
 $\dot{\sigma}_{j}^{z}={}$ $-(\gamma_{j}+\gamma_{\rm loss}^{j}+\gamma_{\rm cm}^{j})(1+\sigma_{j}^{z})-2i\Omega_{\rm cm}^{j}f(t)(\sigma_{j}^{+}-\sigma_{j}^{-})$
-$-2\sum_{l\neq j}\left[\frac{g_{j}g_{l}}{(\kappa/2-i\Delta)}\sigma_{j}^{+}\sigma_{l}^{-}+{\rm h.c.}\right]$
-$-2\sqrt{\gamma^{j}_{\rm cm}}\left[\sigma_{j}^{+}a_{\rm in}+{\rm h.c.}\right]-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]$
-$-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (51)
-Here, the effective cavity-mediated decay $\gamma_{\rm cm}^{j}$, detuning $\delta_{\rm cm}^{j}$, hopping $J_{\rm cm}^{jl}$, and driving strength $\Omega_{\rm cm}^{j}$ are defined in Eqs. (14)-(17).
-Finally, we can find a master equation which leads to the same effective Langenvin equations in Eqs.
-(50)-(51).
-This effective master equation, which incorporates the adiabatic elimination of the cavity mode, is given in Eqs.
-(12)-(13).
+$-2\sum_{l\neq j}\sum_{\omega_{c}^{m}\sim\omega_{d}}\left[\frac{g_{j,m}g_{l,m}}{(\kappa_{m}/2-i\Delta_{m})}\sigma_{j}^{+}\sigma_{l}^{-}+{\rm h.c.}\right]$
+$-2\sum_{\omega_{c}^{m}\sim\omega_{d}}\sqrt{\frac{(g_{j,m})^{2}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}}\left[\sigma_{j}^{+}a_{\rm in}^{m}+{\rm h.c.}\right]$
+$-2\sqrt{\gamma_{j}}\left[\sigma_{j}^{+}b_{\rm in}^{j}+{\rm h.c.}\right]-2\sqrt{\gamma^{j}_{\rm loss}}\left[\sigma_{j}^{+}c_{\rm in}^{j}+{\rm h.c.}\right].$ (60)
+Finally, we can find a master equation whose effective dynamics for the qubits are equivalent to the above quantum Langevin equations (59)-(60).
+This is the case for the effective master equation (13) with effective Hamiltonian (14), provided the cavity-mediated decay $\gamma_{\rm cm}^{jl}$, detuning, hopping $J_{\rm cm}^{jl}$, driving $\Omega_{\rm cm}^{j}$, and detuning $\delta_{\rm cm}^{j}$ are given as in Eqs.
+(39)-(41), with $\omega_{q}^{j}\approx\omega_{d}$.
+These expressions reduce to Eqs.
+(15)-(18) in the case of a mono-mode cavity.
 Appendix B Photon counting and calculation of photon generation probabilities
 Photon detection and counting lies at the heart of Quantum Optics and thus plenty of methods have been developed over time Gardiner and Zoller (2004); Plenio and Knight (1998).
-In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs.
+In this Appendix, we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs.
 In Sec.
@@
 B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach.
-This is useful when the Hilbert space of the system becomes very large, but at the expense of losing precision in the computation of the averages.
+This is useful when the Hilbert space of the system becomes very large but at the expense of losing precision in the computation of the averages.
 B.1 Photon counting in master equation formalism
-In the master equation dynamics the information about the emission of photons into the bath is omitted, and therefore one typically resorts to the input-output formalism (cf.
+In the master equation dynamics, the information about the emission of photons into the bath is omitted, and therefore one typically resorts to the input-output formalism (cf.
 Sec.
 C and Ref.
-Gardiner and Zoller (2004)) in order to relate measurable photonic quantities to multi-time system correlations.
+Gardiner and Zoller (2004)) to relate measurable photonic quantities to multi-time system correlations.
 Nevertheless, obtaining the photon statistics from these system correlations involves the computation of multi-dimensional integrals over time, which can be very computationally costly and inefficient when scaling up the number of emitters or photons to probe.
@@
 For any Markovian master equation for a system state $\tilde{\rho}(t)$, one can find an extended master equation that incorporates the few-photon counting statistics in the dynamics.
-The general recipe is actually very simple and thus we explain it directly on the synchronized multiphoton device described by Eq.
+The general recipe is very simple and thus we explain it directly on the synchronized multiphoton device described by Eq.
 (13).
@@
 where $\tilde{\rho}(t)$ is the state of the system governed by the original master equation without photon counting (13).
-In practical calculations we therefore solve for the extended state $\rho_{\rm ext}(t)$ in Eq.
+In practical calculations, we therefore solve for the extended state $\rho_{\rm ext}(t)$ in Eq.
 (63), and we then obtain the few-photon statistics of the emitted photons by taking simple expectation values on $\rho_{\rm ext}(t)$.
@@
 $P_{N}=\Pi_{1\dots 1}^{1\dots N}={\rm Tr}\{\Lambda_{1}^{1}\dots\Lambda_{1}^{N}\rho_{\rm ext}\}.$ (71)
-In the general case, however, it is important to ensure that the occupation of the last state of the counters is negligible, so that the photon statistics is not affected by the finite size of the counters.
+In the general case, however, it is important to ensure that the occupation of the last state of the counters is negligible so that the photon statistics are not affected by the finite size of the counters.
 B.2 Photon counting in quantum trajectories formalism
 The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al.
-(1992); Mølmer et al.
+(1992); Molmer et al.
 (1993); Dum et al.
@@
 Nevertheless, the advantage of QT formalism is that we evolve pure states instead of density matrices and that it avoids extending the Hilbert space dimension to include counters as in the extended master equation method.
-This two key aspects allows us to dramatically reduce the Hilbert space for the simulations and thus to treat much larger systems composed of many more SPSs [cf.
+These two key aspects allow us to dramatically reduce the Hilbert space for the simulations and thus to treat much larger systems composed of many more SPSs [cf.
 the scalability calculations in Fig.
@@
 III.4-III.5, we discuss how to quantify multi-photon contamination and photon indistinguishability in the emission of SPSs via second-order photon correlation functions.
-These photon correlations are measured via coincidence counts either in the Hanbury Brown and Twiss (HBT) or the Hong-Ou-Mandel (HOM) configurations, and can be expressed as
+These photon correlations are measured via coincidence counts either in the Hanbury Brown and Twiss (HBT) or the Hong-Ou-Mandel (HOM) configurations and can be expressed as
 $G_{\rm HBT}^{(2),j}(\tau)={}$ $\int_{0}^{\infty}\!\!\!\!dt\langle b^{j{\dagger}}_{\rm out}(t)b^{j{\dagger}}_{\rm out}(t+\tau)b^{j}_{\rm out}(t+\tau)b^{j}_{\rm out}(t)\rangle,$ (82)